### v2

# On a generalisation of Cameron’s base size conjecture

Abstract.
Let $G\leqslant{\rm Sym}(\Omega)$ be a finite transitive permutation group with point stabiliser $H$. A base for $G$ is a subset of $\Omega$ whose pointwise stabiliser is trivial, and the minimal cardinality of a base is called the base size of $G$, denoted by $b(G,\Omega)$. Equivalently, $b(G,\Omega)$ is the minimal positive integer $k$ such that $G$ has a regular orbit on the Cartesian product $\Omega^{k}$. A well-known conjecture of Cameron from the 1990s asserts that if $G$ is an almost simple primitive group and $H$ is a so-called non-standard subgroup, then $b(G,\Omega)\leqslant 7$, with equality if and only if $G$ is the Mathieu group ${\rm M}_{24}$ in its natural action of degree $24$. This conjecture was settled in a series of papers by Burness et al. (2007-11).
In this paper, we complete the proof of a natural generalisation of Cameron’s conjecture. Our main result states that if $G$ is an almost simple group and $H_{1},\ldots,H_{k}$ are any core-free non-standard maximal subgroups of $G$ with $k\geqslant 7$, then $G$ has a regular orbit on $G/H_{1}\times\cdots\times G/H_{k}$, noting that Cameron’s original conjecture corresponds to the special case where the $H_{i}$ are pairwise conjugate subgroups. In addition, we show that the same conclusion holds with $k\geqslant 6$, unless $G={\rm M}_{24}$ and each $H_{i}$ is isomorphic to ${\rm M}_{23}$. For example, this means that if $G$ is a simple exceptional group of Lie type and $H_{1},\ldots,H_{6}$ are proper subgroups of $G$, then there exist elements $g_{i}\in G$ such that $\bigcap_{i}H_{i}^{g_{i}}=1$. By applying recent work in a joint paper with Burness, we may assume $G$ is a group of Lie type and our proof uses probabilistic methods based on fixed point ratio estimates.
1. Introduction
Let $G\leqslant\operatorname{Sym}(\Omega)$ be a transitive permutation group on a finite set $\Omega$. Recall that a base for $G$ is a subset of $\Omega$ whose pointwise stabiliser is trivial. We write $b(G,\Omega)$ for the base size of $G$, which is the minimal cardinality of a base for $G$. This classical invariant has been extensively studied since the early days of group theory in the nineteenth century, finding a wide range of applications to other areas of mathematics such as graph theory and model theory, for example. Bases are also fundamental in the computational study of finite groups and the implementation of efficient algorithms. There is an extensive literature on bases for permutation groups and we refer the reader to the survey articles [7, 62, 67] and [20, Section 5] for further details.
Determining the base size of a permutation group is a fundamental problem in group theory and it has attracted significant interest in recent years. In particular, there has been particular focus in the study of base sizes of primitive groups, motivated by some highly influential conjectures by Babai, Cameron, Kantor, and Pyber stated in the 1990s (Recall that a transitive permutation group $G\leqslant{\rm Sym}(\Omega)$ is primitive if a point stabiliser is a maximal subgroup of $G$). All those conjectures have now been resolved.
Recall that a finite group $G$ is almost simple if there exists a nonabelian simple group $G_{0}$ such that $G_{0}\trianglelefteqslant G\leqslant{\rm Aut}(G_{0})$. We call $G_{0}$ the socle of $G$. We say that a subgroup $H$ of a finite group $G$ is core-free if $\bigcap_{g\in G}H^{g}=1$. Note that if $G$ is almost simple, then $H\leqslant G$ is core-free if it does not contain $G_{0}$. Recall the following definition:
Definition.
Let $G$ be a finite almost simple group with socle $G_{0}$ and let $H$ be a core-free maximal subgroup of $G$. We say that $H$ is standard if one of the following holds:
  1. (i)
    $G_{0}=A_{n}$ and the natural action of $H$ on $\{1,\ldots,n\}$ is not primitive;
  2. (ii)
    $G$ is a classical group and $H$ is a subspace subgroup of $G$.
We say that $H$ is non-standard otherwise.
We will define subspace subgroups of classical groups precisely in Section 3 (see Definition 3.4), but roughly speaking, if $G$ is a classical group, then $H$ is standard if it is reducible. That is, $H\cap G_{0}$ acts reducibly on the natural module. Note that if there exists an isomorphism of $G$ to some other group that maps $H$ to a standard subgroup, then we will also say that $H$ is standard in this case. For instance, we note that ${\rm PSL}_{4}(2)\cong A_{8}$ and the subgroup of type ${\rm GL}_{2}(4)$ gets mapped to the intransitive subgroup $(S_{3}\times S_{5})\cap A_{8}$ by an isomorphism. We will therefore say that the subgroup of type ${\rm GL}_{2}(4)$ is standard in ${\rm PSL}_{4}(2)$. Those cases are handled in Table 3.
If $G\leqslant{\rm Sym}(\Omega)$ is a primitive group with point stabiliser $H$, then we will say that the action of $G$ on $G/H$ is non-standard if $H$ is non-standard. It is natural to distinguish between standard and non-standard actions when studying bases. More specifically, we observe that there is a dichotomy between standard and non-standard actions in terms of base sizes. In particular, standard actions tend to admit ‘large bases’. For example, the base sizes of alternating and symmetric groups in standard actions have been determined due to work by several authors (see [21, 35, 45, 46, 68, 69]) and we observe that in the majority of cases, $b(G,\Omega)$ grows with $n$, and consequently $b(G,\Omega)$ can get arbitrarily large as $n\rightarrow\infty$. As we now explain, almost simple groups in non-standard actions exhibit a completely different behaviour with respect to base sizes.
It was conjectured by Cameron and Kantor in [29] that if $G$ is a primitive almost simple permutation group in a non-standard action, then $b(G,\Omega)$ is bounded above by an absolute constant independent of $G$. This was proved by Liebeck and Shalev in [55]. In response, Cameron conjectured in [28, p.122] that the best possible value of this absolute constant is $7$. This conjecture was proved in a series of papers by Burness et al. [18, 23, 25, 26]. In particular, we have the following result:
Theorem.
Let $G\leqslant{\rm Sym}(\Omega)$ be an almost simple primitive permutation group with non-standard point stabiliser $H$. Then $b(G,\Omega)\leqslant 7$ with equality if and only if $(G,H)=({\rm M}_{24},{\rm M}_{23})$.
The aim of this paper is to generalise Theorem Theorem in the setting we describe below:
Let $G$ be a finite group and let $H_{1},\ldots,H_{k}$ be a collection of core-free subgroups of $G$, allowing repetitions. Consider the natural action of $G$ on the Cartesian product
$X=G/H_{1}\times\cdots\times G/H_{k}$
and observe that $G$ has a regular orbit on $X$ if and only if
$\bigcap_{i=1}^{k}H_{i}^{g_{i}}=1$
for some elements $g_{i}\in G$. Following [2], we say that a $k$-tuple $(H_{1},\ldots,H_{k})$ of core-free subgroups of $G$ is regular if $G$ has a regular orbit on $X$, and non-regular otherwise. We will always be interested in the regularity of subgroup tuples all of whose components are core-free, and so, given a subgroup tuple, we will always assume that its components are core-free. We will also say that $\tau$ is conjugate if the $H_{i}$ are pairwise conjugate. Finally, we call the minimal $r$ such that all $r$-tuples of core-free subgroups of $G$ are regular the regularity number of $G$, which we denote by $R(G)$.
Recall that a transitive group $G\leqslant{\rm Sym}(\Omega)$ has a base of size $k$ if and only if $G$ has a regular orbit on $\Omega^{k}$. Therefore, a conjugate tuple $(H_{1},\ldots,H_{k})$ is regular if and only if $G$ admits a base of size $k$ in its natural action on $G/H_{1}$. This implies that $k\geqslant b(G,H_{1})$.
If $G$ is almost simple and all the components of a tuple $\tau=(H_{1},\ldots,H_{k})$ are non-standard, we will say that $\tau$ is non-standard. Hence, in this language, Theorem Theorem asserts that every conjugate non-standard $7$-tuple of $G$ is regular.
The regularity of subgroup tuples was introduced and studied in [2] for almost simple groups with alternating and sporadic socle, where Burness and the author stated generalisations of some of the most influential base size conjectures in the regularity setting, including Cameron’s conjecture. In particular, if $G$ is a finite almost simple group and $R_{\rm ns}(G)$ denotes the minimal integer $r$ such that all non-standard tuples of $G$ are regular, then they conjecture that $R_{\rm ns}(G)\leqslant 7$, with equality if and only if $G={\rm M}_{24}$. Note that this conjecture is a vast generalisation of Theorem Theorem. As noted above, Theorem Theorem implies that all conjugate non-standard $7$-tuples of a finite almost simple group $G$ are regular and it is claimed that the conjugacy assumption can be removed.
The above conjecture was proved in [2] for almost simple groups with alternating and sporadic socle, which reduces the problem to the almost simple groups of Lie type. The main aim of this paper is to complete the proof of this conjecture. We now state our main theorem:
Theorem 1.
Let $G$ be a finite almost simple group. Then $R_{\rm ns}(G)\leqslant 7$ with equality if and only if $G={\rm M}_{24}$.
It is shown in [2], that up to conjugacy, the only non-regular non-standard $6$-tuple when $G={\rm M}_{24}$ is $(H,H,H,H,H,H)$, where $H={\rm M}_{23}$. Moreover, note that if $G$ is sporadic or exceptional, then all of its subgroups are non-standard, so Theorem 1 asserts that $R(G)\leqslant 7$ in this case.
As we will see in Section 3, we can prove a stronger statement than Theorem 1 for classical groups. In particular, if $G$ is an almost simple classical group, then we will show that every non-standard $4$-tuple is regular, unless $G={\rm U}_{6}(2).2$, in which case $R_{\rm ns}(G)=5$ and the only non-regular non-standard $4$-tuple up to conjugacy is $(H,H,H,H)$ with $H={\rm U}_{4}(3).2^{2}$. Combining this with the results in [2], we obtain a more refined version of Theorem 1, which we state below:
Theorem 2.
Let $G$ be a finite almost simple group with socle $G_{0}$:
  • (i)
    If $G_{0}=A_{n}$, then $R_{\rm ns}(G)\leqslant 6$ with equality if and only if $G=A_{8}$. Moreover, if $n\geqslant 13$, then $R_{\rm ns}(G)=2$;
  • (ii)
    If $G_{0}$ is sporadic, then $R_{\rm ns}(G)=R(G)\leqslant 7$ with equality if and only if $G={\rm M}_{24}$. Moreover, the exact value of $R(G)$ is recorded in [2, Table 3];
  • (iii)
    If $G_{0}$ is a classical group, then $R_{\rm ns}(G)\leqslant 5$ with equality if and only if $G={\rm U}_{6}(2).2$;
  • (iv)
    If $G_{0}$ is an exceptional group, then $R_{\rm ns}(G)=R(G)\leqslant 6$.
We now briefly discuss the methods we use to prove Theorem 1 for groups of Lie type. These are similar to those used in the proof of Cameron’s conjecture. In particular, our main tool will be a generalised version of the powerful probabilistic method based on fixed point ratio estimates that was first introduced in [63] by Liebeck and Shalev for the study of bases.
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a tuple of core-free subgroups of $G$. We will write $X_{\tau}=G/H_{1}\times\cdots\times G/H_{k}$ and $\mathbb{P}(G,\tau)$ for the probability that a randomly chosen point in $X_{\tau}$ is in a regular orbit. Our general approach will be to show that
$Q(G,\tau)=1-\mathbb{P}(G,\tau)<1.$
Note that generally, it is impossible to compute $Q(G,\tau)$ for a finite group $G$. However, in most cases it is sufficient to work with the upper bound below:
$Q(G,\tau)\leqslant\widehat{Q}(G,\tau):=\sum_{i=1}^{t}|x_{i}^{G}|\cdot\left(\prod_{j=1}^{k}{\rm fpr}(x_{i},G/H_{j})\right)$
where $x_{1},\ldots,x_{t}$ is a set of representatives of the prime order conjugacy classes in $G$ and ${\rm fpr}(x,G/H)$ denotes the fixed point ratio of $x$ in the natural action of $G$ on the cosets of $H$. That is, the proportion of fixed points of $x$ in this action. We refer the reader to Section 2.1 for more information.
If $G$ is an almost simple classical group with socle $G_{0}$ with natural module $V$ and $\tau$ is a non-standard tuple of $G$, then our main tool when estimating $\widehat{Q}(G,\tau)$ the following ‘zeta type function’
$\eta_{G}(t)=\sum_{C\in\mathscr{C}}|C|^{-t}$
where $\mathscr{C}$ denotes the set of prime order $G_{0}$-conjugacy classes of $G$. More specifically, Burness obtains upper bounds on ${\rm fpr}(x,G/H)$ for a prime order element $x$ and a non-standard subgroup $H$ in terms of the size of the $G$-class of $x$ in [14, 15, 16, 17]. Using these allows us to bound $\widehat{Q}(G,\tau)$ by $\eta_{G}(t)$ for some $t\in(0,1)$, and our problem reduces to showing that $\eta_{G}(t)<1$. In some cases, we will need stronger estimates on ${\rm fpr}(x,G/H)$ than the ones obtained [14, 15, 16, 17], and we obtain those in Section 4. Finally, when ${\rm dim}V\leqslant 5$, using $\eta_{G}(t)$ will no longer be possible, so we take a more direct approach to estimate $\widehat{Q}(G,\tau)$.
If $G$ is exceptional, then we obtain a character theoretic approach instead. In particular, if $H$ is a maximal subgroup of $G$ and $\Omega=G/H$, then we have
${\rm fpr}(x,G/H)=\frac{\chi(x)}{|\Omega|},$
where $\chi(x)$ is the corresponding permutation character. Noting this, we can employ powerful tools developed by Lawther, Liebeck and Seitz in [51] to estimate ${\rm fpr}(x,G/H)$. In this way, for every prime order element $x\in G$, we are able to compute
$f(x,q):=\max\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{M}_{G}\}$
where $\mathcal{M}_{G}$ denotes the set of maximal subgroups of $G$. This allows us to obtain a uniform upper bound on $\widehat{Q}(G,\tau)$ that applies to any $6$-tuple $\tau$ of maximal subgroups of $G$.
The main challenge when the conjugacy assumption on our subgroup tuples is removed is that we need to adopt a uniform approach that is applicable to any non-standard tuple. In particular, our arguments no longer depend on a particular action of our group $G$, which means that we are always bound to worst case fixed point ratio estimates over all maximal subgroups of $G$ for all prime order elements of $G$.
Existing worst case bounds in the literature for a given element $x\in G$ might be good enough for a particular action of $G$, but might not be sufficient in this general setting. This means that we will often need to obtain stronger fixed point ratio estimates for our purposes. We state those in Theorem 4.1 for the classical groups. For exceptional groups, we record worst case fixed point ratio estimates over non-parabolic actions in Propositions 8.1, 8.2, 8.3, and 8.4. Our main source for obtaining those is inspection of [25] and [19], and in some instances we need to improve what currently is available. We give detailed information in Sections 2, 4 and 8.
Additionally, for classical groups there are certain instances where existing information on $\eta_{G}(t)$ is no longer sufficient on our purposes and we will need stronger results. We obtain those in Section 5.
The heavily probabilistic approach we adopt to prove Theorem 1 allows us to also obtain an asymptotic result. In particular, if $G$ is an almost simple group of Lie type, $\mathcal{P}$ denotes the set of non-standard subgroups of $G$ and
$\mathbb{P}(G,c)=\min\{\mathbb{P}(G,\tau)\,:\,\tau\in\mathcal{P}^{c}\},$
then we have the following theorem:
Theorem 3.
If $G$ is a finite almost simple group of Lie type, then $\mathbb{P}(G,6)\rightarrow 1$ as $|G|\rightarrow\infty$.
This paper is organised as follows: In Section 2 we record some preliminary results. In Section 3, we proceed by recording some preliminaries that only apply to classical groups. We first discuss subgroup structure and conjugacy classes of classical groups, and then proceed to give a precise definition of a subspace subgroup of a classical group. The rest of the section is devoted to probabilistic and computational methods, which are the key tools that we will be using.
Section 4 is devoted to improving existing fixed point ratio estimates for low-dimensional classical groups, in Section 5 we prove some results relating to the ‘zeta-type’ function introduced in Definition 3.10, and Section 6, the final section on classical groups is devoted to the proof of Theorem 3.1, a refinement of Theorem 1 for classical groups.
In Section 7 we record some preliminaries that only apply to exceptional groups. In particular, the section is divided into two subsections, where we briefly discuss subgroup structure and conjugacy classes respectively. In Section 8 we record fixed point ratio estimates for non-parabolic actions of exceptional groups, and then in Section 9 we proceed to discuss fixed point ratios for parabolic actions. Finally, Sections 10.1, 10.2, and 10.3 are devoted to the proof of Theorem 7.1, a refinement of Theorem 1 for exceptional groups.
Notation. Our notation is fairly standard. We write ${\rm Inndiag}(G_{0})$ for the group of inner-diagonal automorphisms of a simple group of Lie type $G_{0}$. Moreover, we will also often write $\pi(n)$ to denote the set of prime divisors of a positive integer $n$ and $\pi_{a}(n)$ to denote the set of prime divisors of $n$ that are no less than $a$. In a similar fashion, we will denote the set of prime numbers at most $s$ by $\pi_{s}$ and the set of odd prime numbers at most $s$ by $\pi^{\prime}_{s}$. For a group $G$, we will denote the number of elements of order $s$ in $G$ by $i_{s}(G)$. For classical groups, we write ${\rm PSL}^{+}_{n}(q)={\rm PSL}_{n}(q)={\rm L}_{n}(q)$ and ${\rm PSL}^{-}_{n}(q)={\rm PSU}_{n}(q)={\rm U}_{n}(q)$, We use the same notation and terminology as [22] for field, graph, and graph-field automorphisms. Finally, we follow the definition of [47] for the $\mathscr{C}_{i}$ and $\mathscr{S}$ classes of maximal subgroups of classical groups. We will also be using the standard notation $E_{6}^{\epsilon}(q)$ for the finite exceptional groups of type $E_{6}$, where $E_{6}^{+}(q)=E_{6}(q)$ and $E_{6}^{-}(q)={}^{2}E_{6}(q)$. We will introduce more technical terminology, when and where appropriate in the paper.
Acknowledgements. This paper is part of the author’s doctoral thesis under the supervision of Professor Tim Burness. The author thanks Professor Burness for his encouragement, feedback, and invaluable advice. She also thanks him for introducing her to this interesting problem. The author also thanks Frank Lübeck for sharing GAP-readable data on Foulkes functions of exceptional groups. Finally, she thanks the Heilbronn Institute for Mathematical Research for funding her PhD at the university of Bristol.
2. Preliminaries
In this section, we record some preliminary lemmas that we will repeatedly use to prove Theorem 1. We first discuss a small number theoretic lemma that we will repeatedly use to derive bounds on conjugacy class sizes, we then proceed to discuss probabilistic methods, and finally we briefly comment on computational methods.
Lemma 2.1.
The following are true:
  • If $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ be two sets of integers, all at least $2$, then
    $\frac{\prod_{i=1}^{m}(q^{a_{i}}-1)}{\prod_{i=1}^{n}(q^{b_{i}}-1)}>2q^{\sum a_{i}-\sum b_{i}}$
  • If $a,b\in\mathbb{Z}^{+}$ with $a<b$, then
    $\frac{(q^{a}+1)}{(q^{b}+1)}<q^{a-b}$
This lemma is [51, 1.2 (i) and (iii)]
2.1. Probabilistic methods
Our main tool for proving our main result is the probabilistic method. Recall that if $G\leqslant{\rm Sym}(\Omega)$ is a transitive permutation group with point stabiliser $H$, then the fixed point ratio of $x\in G$ is defined as
${\rm fpr}(x,G/H)=\frac{|C_{\Omega}(x)|}{|\Omega|},$
where $C_{\Omega}(x)=\{\omega\in\Omega\,:\,\omega^{x}=\omega\}$ is the set of fixed points of $x$ and $x^{G}$ denotes the conjugacy class of $x$ in $G$. We will use a generalisation of the powerful probabilistic approach based on fixed point ratio estimates that was first introduced in [63] by Liebeck and Shalev for the study of bases, which is applicable in the more general regularity setting. We now state and prove some preliminary lemmas about fixed point ratios for finite almost simple permutation groups that will help us confirm the regularity of tuples.
We first state two lemmas that appeared in [2], which are natural generalisations of equivalent lemmas used for the study of base sizes of permutation groups, which can be found in [18]. Throughout the paper, if $n$ is a natural number, we will often write $[n]$ to denote the set $\{1,\ldots,n\}$.
Lemma 2.2.
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a core-free tuple of subgroups of $G$. Then $\tau$ is regular if
$\widehat{Q}(G,\tau):=\sum_{i=1}^{t}|x_{i}^{G}|\cdot\left(\prod_{j=1}^{k}{\rm fpr}(x_{i},G/H_{j})\right)<1,$
where $x_{1},\ldots,x_{t}$ is a set of representatives of the conjugacy classes in $G$ of elements of prime order.
Lemma 2.3.
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a core-free tuple of subgroups of $G$. Suppose $x_{1},\ldots,x_{m}$ are elements in $G$ such that $|x_{i}^{G}|\geqslant B$ for all $i\in[m]$ and $\sum_{i=1}^{m}|x_{i}^{G}\cap H_{j}|\leqslant A_{j}$ for all $j\in[k]$. Then
$\sum_{i=1}^{m}|x_{i}^{G}|\cdot\left(\prod_{j=1}^{k}{\rm fpr}(x_{i},G/H_{j})\right)\leqslant B^{1-k}\cdot\prod_{j=1}^{k}A_{j}$
Our aim will be to derive fixed point ratio bounds that apply to all almost simple groups with a fixed socle $G_{0}$. We now prove a small lemma that will allow us to deal with all such cases simultaneously.
Lemma 2.4.
Let $G$ be an almost simple group with socle $G_{0}$, let $H$ be a core-free maximal subgroup of $G$, and let $x\in G$ be an element of prime order $r$. If $G_{1}=\langle G_{0},x\rangle$ and $H_{1}=H\cap G_{1}$, then ${\rm fpr}(x,G/H)={\rm fpr}(x,G_{1}/H_{1})$.
Proof.
First note that the maximality of $H$ implies that $G=G_{0}H$. In particular, $G=G_{1}H$ and therefore $G/H=G_{1}H/H$. This implies that $G_{1}$ also acts transitively on the cosets of $H$.
We will now prove that $G/H=\Omega$ and $G_{1}/H_{1}=\Omega_{1}$ are isomorphic transitive $G_{1}$-sets. Define $\phi:\Omega\rightarrow\Omega_{1}$ by $\phi(Hg)=H_{1}g$ for $g\in G_{1}$. We claim that $\phi$ is a bijection. Suppose that $H_{1}y=H_{1}z$ for some $y,z\in G_{1}$. Then $yz^{-1}\in H_{1}$, and so $yz^{-1}\in H$, since $H_{1}\leqslant H$. Therefore, $\phi$ is injective. Moreover, $|G:H|=|G_{1}:H_{1}|$, and so $\phi$ is bijective.
Now note that if $y,z\in G_{1}$, then
$\phi(Hyz)=H_{1}yz=\phi(Hy)z$
so the actions of $G_{1}$ on $G/H$ and $G_{1}/H_{1}$ are indeed equivalent. This implies that $|C_{\Omega}(x)|=|C_{\Omega_{1}}(x)|$, and therefore
$\mbox{{\rm fpr}}(x,G/H)=\frac{|C_{\Omega}(x)|}{|\Omega|}=\frac{|C_{\Omega_{1}}(x)|}{|\Omega_{1}|}={\rm fpr}(x,G_{1}/H_{1}),$
as required. ∎
Remark 2.5.
Note that if $G_{0}\leqslant\langle G_{0},x\rangle\leqslant{\rm Inndiag}(G_{0})=G_{I}$, then Lemma 2.4 implies that
${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{I}}\cap H|}{|x^{G_{0}}|}$
We will repeatedly use this observation when computing bounds for ${\rm fpr}(x,G/H)$.
2.2. Computational methods
Here we briefly discuss how we use GAP [43] and Magma [10] in our proofs. Note that we only discuss some of the methods here. In particular, for exceptional groups, we heavily use the polynomial functionality in GAP and Magma to compute fixed point ratios, but we discuss this in detail later in Section 9.
Let $G$ be a finite almost simple group of Lie type over $\mathbb{F}_{q}$ with socle $G_{0}$. There are two natural instances where we use Magma in our proofs. The first is in Section 4 where $G$ is classical and we aim to improve the existing fixed point ratio bound given in Theorem 3.9. There are some cases where we can only derive rather crude estimates, which prove the result when $q$ is sufficiently large, but do not suffice when $q$ is relatively small. In all of those cases, the dimension of the natural module for $G_{0}$ is relatively low (in particular no more than $8$), so we construct $G$ using the Magma function AutomorphismGroupSimpleGroup() and call MaximalSubgroups(G) to obtain a set of representatives of the $G$-classes of maximal subgroups of $G$. We then filter out the non-standard subgroups of $G$ and one can verify the necessary fixed point ratio bounds for all of those using the following functions:
fpr := function( G, H, g )
  cl:=Classes(H);
  A:=[cl[i][2] : i in [1..#cl] | IsConjugate(G,cl[i][3],g)];
  Append(~A,0);
  return &+A*#Centraliser(G,g)/#G;
end function;
MaxFPR := function(G,H)
  cl:=Classes(G);
  a:=[i : i in [1..#cl] | IsPrime(cl[i][1])];
  b:=[i : i in a | fpr(G,H,cl[i][3]) gt 0];
  c,d:=Maximum([-Log(fpr(G,H,cl[i][3]))/Log(cl[i][2]): i in b]);
  return c;
end function;
The function fpr takes as input a permutation group $G$, a subgroup $H$ of $G$ and an element $g\in G$ and computes ${\rm fpr}(g,G/H)$. The function MaxFpr takes as input a group $G$ a subgroup $H$ of $G$ and and returns the largest $t$ so that ${\rm fpr}(x,G/H)<|x^{G}|^{-t}$ for all $x\in G$ of prime order. The case where the computation using the code above is the most time consuming is when $G_{0}={\rm PSp}_{8}(3)$, in which case we can get output in less than an hour.
The second instance where we will be using Magma is in some cases where we want to check confirm the regularity of a non-standard tuple $\tau$ of $G$. As mentioned above, our standard method is to bound $\widehat{Q}(G,\tau)$ above by 11, but again, there are instances where the bounds we obtain are not good enough for small values of $q$, so in these cases we use the function RegTuplesPlus presented in [1, p.4] instead, to verify the regularity all $R_{\rm ns}(G)$-tuples of $G$. The function takes as input a finite group $G$, an ordered tuple $M=[H_{1},\ldots,H_{t}]$ of subgroups of $G$ and two positive integers $l$ and $N$ with $l\geqslant 2$. The output is the minimal integer $k\geqslant l$ such that all $k$-tuples of $G$ are regular, together with all the non-regular (k1)(k-1)-tuples of $G$. We refer the reader to [1] for the code and more information.
To be able to use RegTuplesPlus, both $|G|$ needs to be relatively small. In particular, we only make use of this function only if $G_{0}\in\{{}^{2}F_{4}(2)^{\prime},{}^{3}D_{4}(2)\}$ for exceptional groups and when $G$ is at most $8$-dimensional for classical groups. In the classical case, the maximal value of $q$ we use RegTuplesPlus for, highly depends on the dimension of the natural module and the Lie type of the group we are considering, but it is always relatively small. For example, if $G_{0}={\rm PSp}_{8}(q)$, then we only use Magma for the case $q=2$, whilst if $G_{0}=L_{2}(q)$, then we use Magma for $q\geqslant 31$.
Finally, if $G_{0}=F_{4}(2)$ or $E_{6}^{\epsilon}(2)$, then we will use the Character Table Library in GAP [13] to compute fixed point ratios for certain non-parabolic maximal subgroups.
3. Classical group preliminaries
Throughout this section, as well as Sections 4, 5 and 6, we will take $G$ to be a finite almost simple classical group over a finite field $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$ and some positive integer $f$. We will write $G_{0}$ for the socle of $G$ and $V$ will denote the natural module for $G_{0}$. Moreover, let $G_{I}$ denote the group ${\rm Inndiag}(G_{0})$ of inner-diagonal automorphisms of $G_{0}$.
Our goal here is to prove Theorem 1 for classical groups. In particular, we will prove the following stronger statement:
Theorem 3.1.
Let $G$ be a finite almost simple classical group. Then either $R_{\rm ns}(G)\leqslant 5$ with equality if and only if $G={\rm U}_{6}(2).2$. Moreover, one of the following holds:
  • (i)
    $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$;
  • (ii)
    $G_{0}={\rm Sp}_{6}(q)$ with $q$ even or $G_{0}=\Omega_{7}(q)$ with $q$ odd and $\mathbb{P}(G,6)\rightarrow 1$ as $|G|\rightarrow\infty$.
Remark 3.2.
Let us record some comments on the statement of Theorem 3.1.
  • (a)
    We will prove that $(H,H,H,H)$ is the only non-regular $4$-tuple of ${\rm U}_{6}(2).2$ up to conjugacy, where $H={\rm U}_{4}(3).2^{2}$.
  • (b)
    Theorem 3.1 is a direct regularity analogue of [18, Theorem 1]. In particular, Burness shows that $b(G,H)\leqslant 5$ for every almost simple primitive classical group with non-standard point stabiliser $H$, with equality if and only if $(G,H)=({\rm U}_{6}(2).2,{\rm U}_{4}(3).2^{2})$.
  • (c)
    Our methods will be similar to the ones used in [18]. However, recall from Section 1 that removing the conjugacy assumption on our tuples requires us to work with uniform worst case fixed point ratio bounds. This means that whilst we will heavily rely on the main theorem of [14, 15, 16, 17] to obtain sufficient fixed point ratio estimates in the majority of the cases, there will be some cases, which we outline in Section 4, where we will require stronger bounds. We obtain those bounds in Theorem 4.1.
  • (d)
    As we will explain in Section 3.3,we will be heavily using a function $\eta_{G}(t)$, which is a variant of a ’zeta type function’ introduced by Liebeck and Shalev in [61] in the proof of Theorem 3.1. We will heavily rely on results in [18] for this function, but as is the case with fixed point ratios, those results will not always be applicable in this more general setting, so we will need to refine those. We do this in Section 5.
We now discuss some preliminary results around classical groups. In particular, we will briefly discuss Aschbacher’s theorem that describes the subgroup structure of classical groups and record some information on conjugacy classes of prime order elements. We will then proceed by giving a formal definition of a subspace subgroup, and we will finally provide some more information on probabilistic and computational methods that only apply to the classical groups.
3.1. Subgroup structure
The subgroup structure of finite classical groups is described by Aschbacher in [4], where he defines nine separate collections and proves that every maximal subgroup of a finite classical group is contained in one of those collections. Note that the situation is slightly different when ${\rm P\Omega}^{+}_{8}(q)$ or ${\rm PSp}_{4}(2^{f})$ and we discuss this below. The first eight collections, denoted by $\mathscr{C}_{1},\ldots,\mathscr{C}_{8}$ are often referred to as the geometric collections, and the final collection, which we denote by $\mathscr{S}$ is known as the non-geometric collection. The members of the geometric collections can be described in terms of the underlying geometry of $V$. For example, the members of $\mathscr{C}_{2}$ are stabilisers of appropriate direct sum decompositions, and similarly, the members of $\mathscr{C}_{4}$ and $\mathscr{C}_{7}$ are stabilisers of appropriate tensor product decompositions. The $\mathscr{S}$-collection consists of subgroups that correspond to some absolutely irreducible representation on $V$ of a covering group of $H\cap G_{0}$ (see [47, p.3] for a precise definition). Note that the groups in $\mathcal{S}$ are almost simple. Throughout this paper, we will adopt the definition of [47] for the collections of maximal subgroups, which differs slightly from the definition in the original paper by Aschbacher [4]. We give a brief description of each of the collections in Table 1.
Let $H$ be a maximal subgroup of $G$ such that $H\not\in\mathscr{S}$ and set $H_{0}=G_{0}\cap H$. If $H_{0}$ is maximal, then $H_{0}\in\mathscr{C}$, where $\mathscr{C}=\mathscr{C}_{1}\cup\cdots\cup\mathscr{C}_{8}$. If $G_{0}={\rm PSp}_{4}(2^{f})$, or ${\rm P\Omega}_{8}^{+}(q)$, then the existence of exceptional automorphisms gives rise to an extra collection of so-called novelty subgroups. Following the notation in [22], we will denote this class by $\mathscr{N}$. We give a description of all those subgroups in Table 2, although the only cases of groups in $\mathcal{N}$ that will be considered in this paper are the ones where $G_{0}={\rm PSp}_{4}(q)$ for even $q$ and $H$ is of type ${\rm O}^{\epsilon}_{2}(q)\wr S_{2}$ or ${\rm O}_{2}^{-}(q).2$. We note that Aschbacher’s original paper [4] only provides partial information for the case where $G_{0}={\rm P\Omega}^{+}_{8}(q)$ and $G$ contains triality graph automorphisms. In this case, the maximal subgroups were determined in [49] by Kleidman.
We will repeatedly refer to the type of a subgroup, following the terminology in [47]. If $H\in\mathscr{S}$, then the type of $H$ will be the socle of the almost simple group $H$. On the other hand, if $H\in\mathscr{C}$, then the type of $H$ will give an approximate group-theoretic description of $H$ and will often indicate the underlying geometry associated to it. For example, if $G_{0}={\rm PSL}_{n}(q)$ and $H$ is of type ${\rm GL}_{k}(q)\wr S_{t}$, then $H$ is a stabiliser of a direct sum decomposition $V=V_{1}\oplus\cdots\oplus V_{t}$, where ${\rm dim}V_{i}=k$ for all $i$.
Our standard reference for the existence, structure and maximality of subgroups in $\mathscr{C}$ will be [47], where Kleidman and Liebeck provide a complete description of the structure of geometric structures and determine maximality (up to conjugacy) for $n\geqslant 13$. For classical groups with dimension at most $12$ our standard reference becomes [12], where Bray, Holt and Roney-Dougal determine all the maximal subgroups (up to conjugacy) of the low-dimensional classical groups.
In view of the above discussion, we now provide a version of Aschbacher’s theorem on subgroup structure of classical groups:
Theorem 3.3.
Let $G$ be an almost simple classical group with socle $G_{0}$, and let $H$ be a core-free maximal subgroup of $G$. Then $H\in\mathscr{C}\cup\mathscr{S}\cup\mathscr{N}$.
[…]
Table 1. The Aschbacher collections.
[…]
Table 2. The $\mathscr{N}$-collection.
In this paper we are interested in non-standard subgroup tuples of $G$. In the classical setting, a core-free subgroup of $G$ is non-standard if and only if it is not a subspace subgroup. We begin by giving a precise definition of a subspace subgroup.
Definition 3.4.
A core-free maximal subgroup $M$ of $G$ not containing $G_{0}$ is a subspace subgroup if one of the following holds:
  1. (i)
    $M\in\mathscr{C}_{1}$;
  2. (ii)
    $(G_{0},p)=({\rm Sp}_{2m}(q)^{\prime},2)$ and $M\in\mathscr{C}_{8}$;
  3. (iii)
    $(G,M)$ is displayed in Table 3.
In Table 3 we list all the cases not covered by (i) and (ii) in Definition 3.4. Those are all instances where there exists some isomorphism $\phi:G\rightarrow K$ where ${\rm soc}(K)=T\cong G_{0}$ and one of the following holds:
  • (a)
    $\phi(M)=L$ is a subspace subgroup of Type (i) or (ii) from Definition 3.4;
  • (b)
    $T=A_{n}$ and $L$ does not act primitively on $\{1,\ldots,n\}$.
In the first column of Table 3 we list $G_{0}$, in the second column the type of $M$, in the third column we list the pairs with entries $T$ and the type of $L$, and finally, in the fourth column we mention the conditions under which such an isomorphism occurs, if any. We claim that Table 3 indeed covers all relevant cases. First suppose that $\phi$ is an automorphism. Noting that inner, diagonal, and field automorphisms do not permute the Aschbacher families, we deduce that $\phi$ must be a graph automorphism. Moreover, we must have $G_{0}\in\{{\rm P\Omega}_{8}^{+}(q),{\rm PSp}_{4}(q)^{\prime}\}$, as graph automorphisms of linear and unitary groups again preserve Aschbacher families. In those two cases, we verify that $\phi(M)$ is as claimed in Table 3.4 by inspecting [12, Tables 8.14, 8.50]. If $\phi$ is not an automorphism, then the possible isomorphisms of classical groups are listed in [47, 2.9.1], and by inspecting the Tables in [12, Chapter 8] and comparing orders of maximal subgroups, we verify that Table 3 is complete.
[…]
Table 3. The standard subgroups of classical groups arising in part (iii) of Definition 3.4.
3.2. Conjugacy classes
In this section we will give a brief description of the conjugacy classes of prime order elements in classical groups, and we will establish the notation and terminology that we will be using throughout. We refer the reader to [22, Chapter 3] for a more detailed description of those classes and to [22, Chapter 2] for a more detailed analysis of the basic properties of the classical groups, as well as their diagonal, field, graph, and graph-field automorphisms. We note that due to a theorem of Steinberg [73, Theorem 30], every element in an almost simple classical group is a product of an inner, a diagonal, a field, and a graph automorphism.
First let $x\in G_{I}$ be an element of prime order $r$. We will say that $x$ is unipotent if $r=p$ and that $x$ is semisimple if $r\neq p$. The description of $x$ will differ depending on whether it is unipotent or semisimple, so we divide the analysis into two separate sections.
Semisimple elements
Here we will assume that $x$ is semisimple. We start by stating a fundamental theorem ([44, 4.2.2.(j)]) that we will repeatedly use throughout this paper:
Proposition 3.5.
If $T$ is a finite simple group of Lie type and let $x\in T$ be a semisimple element. Then $x^{T}=x^{{\rm Inndiag}(T)}$.
Note that this theorem applies to all finite simple groups, and we will in fact make use of it when dealing with exceptional groups.
The analysis slightly differs when $x$ is an involution, so we will first discuss the structure of semisimple elements of odd order and then proceed to briefly comment on involutions. More information on all this can be found in [22, Chapter 3].
First suppose that $x$ has odd order $r$ and let $i\geqslant 1$ be minimal such that $r$ divides $q^{i}-1$. As in [22], we will write
[…]
First assume that $c\geqslant 2$. Then [22, 3.1.3] implies that $x$ lifts to a unique element $\hat{x}$ of order $r$ in the corresponding matrix group. That is, $x=\hat{x}Z$, where $Z$ is the centre of this matrix group. We will now focus on the linear case, that is $G_{0}={\rm L}_{n}(q)$, to describe the structure of prime order classes, but one can get similar descriptions with minor modifications for the unitary, symplectic, and orthogonal groups.
Now note that $\hat{x}$ is diagonalisable in $\mathbb{F}_{q^{i}}$, but not any proper subfield, so Maschke’s Theorem implies that $\hat{x}$ fixes a direct sum decomposition
[…]
of the natural ${\rm GL}_{n}(q)$-module $V$, where
$C_{V}(\hat{x})=\{v\in V\,:\,v\hat{x}=v\}$
is the $1$-eigenspace of $\hat{x}$ and $U_{j}$ is an $j$-dimensional subspace of $V$, on which $\hat{x}$ acts irreducibly.
Now let $\mathscr{S}_{r}$ denote the set of non-trivial $r$-th roots of unity in $\mathbb{F}_{q^{i}}$, and observe that the standard Frobenius automorphism $\sigma:\mathbb{F}_{q^{i}}\rightarrow\mathbb{F}_{q^{i}}$ defined by $\lambda\mapsto\lambda^{q}$ induces a permutation on $\mathscr{S}_{r}$. In particular, all $\sigma$-orbits have length $i$, so there are precisely t=(r1)/it=(r-1)/i distinct $\sigma$-orbits. We will write $\Lambda_{1},\ldots,\Lambda_{t}$ to denote those $\sigma$-orbits, where
$\Lambda_{j}=\{\lambda_{j},\lambda_{j}^{q},\ldots,\lambda_{j}^{q^{i-1}}\}$
for some $\lambda_{j}\in\mathscr{S}_{r}$. We now note that the set of eigenvalues in $U_{j}\otimes\mathbb{F}_{q^{i}}$ coincides with some $\sigma$-orbit, so we may abuse notation and write
$\hat{x}=[\Lambda_{1}^{a_{1}},\Lambda_{2}^{a_{2}},\ldots,\Lambda_{t}^{a_{t}},I_{e}],$
[…]
$C_{{\rm GL}_{n}(q)}(\hat{x})={\rm GL}_{e}(q)\times\prod_{j=1}^{t}{\rm GL}_{a_{j}}(q^{i})$
With minor modifications we can get similar descriptions for odd order semisimple elements in all classical groups. For example if $G_{0}={\rm PSp}_{n}(q)$, then the main difference is that when $i$ is odd, then the $\sigma$-orbits come in inverse pairs $(\Lambda,\Lambda^{-1})$, and so every element of order $r$ is of the form $(\Lambda_{1},\Lambda_{1}^{-1})^{a_{1}},\ldots,(\Lambda_{t},\Lambda_{t}^{-1})^{a_{t}},I_{e}]$. And again, two elements of order $r$ are ${\rm Sp}_{n}(q)$-conjugate if and only if they have the same eigenvalues in $\mathbb{F}_{q^{i}}$, so we get a similar description for odd order semisimple classes. We also get a similar description of the centraliser structure of $\hat{x}$. In particular, if $\widehat{G}$ denotes the lift of $G_{I}$ in ${\rm GL}_{n}(q)$ (or ${\rm GL}_{n}(q^{2})$ if $G$ is unitary), then [22, Remark 3.3.4, Remark 3.4.4, Remark 3.5.6] give
[…]
Finally, if $c=1$, then we still write $x=\hat{x}Z$, but in this case the lift in $\hat{x}$ is not unique, so additional conjugacy classes arise. A precise analysis of conjugacy in this case is given in [22, 3.2.2, 3.3.3] and the structure of $C_{G_{I}}(x)$ is recorded in [22, Tables B.3 and B.4].
We now briefly comment on the case where $G_{0}={\rm PSL}_{n}(q)$ and $H$ is a $\mathscr{C}_{3}$-subgroup of $G$, as we will often come across this situation in this paper. Set $L={\rm GL}_{n/k}(q^{k})$, let $W$ be the natural $L$-module over $\mathbb{F}_{p^{k}}$ and consider the natural embedding of $L$ in ${\rm GL}_{n}(q)$ with respect to a fixed $\mathbb{F}_{q^{k}}$-basis $\{w_{1},\ldots,w_{n/k}\}$ for $W$. We refer the reader to [22, Construction 2.1.4] for more information. Now set $i_{0}$ to be the smallest positive integer such that $r$ divides $q^{ki_{0}}-1$, and note that
$i_{0}=\begin{cases}i/k\text{ if }k\text{ divides }i\\ i\text{ otherwise}\end{cases}$
If $\Gamma_{1},\ldots,\Gamma_{s}$ denote the distinct $\sigma_{0}$-orbits on $\mathscr{S}_{r}$, where $\sigma_{0}$ is the permutation defined by $\lambda\mapsto\lambda^{q^{k}}$ for all $\lambda\in\mathscr{S}_{r}$, then either each $\sigma$-orbit is a union of $k$ distinct $\sigma_{0}$-orbits, or $i=i_{0}$ and $\sigma$ and $\sigma_{0}$-orbits coincide. In particular, [22, 5.3.2] gives the following description of the embedding of $\hat{x}\in L$ inside ${\rm GL}_{n}(q)$:
Lemma 3.6.
If $\hat{x}=[\Gamma_{1},\ldots\Gamma_{s},I_{e}]\in L$, then, up to conjugacy,
$\hat{x}=[\Gamma_{1}^{a_{1}},(\Gamma_{1}^{q})^{a_{1}},\ldots,(\Gamma_{1}^{q^{k-1}})^{a_{1}},\ldots,\Gamma_{s}^{a_{s}},(\Gamma_{s}^{q})^{a_{s}},\ldots,(\Gamma_{s}^{q^{k-1}})^{a_{s}},I_{ke}]$
on $V$. In particular, if $i=i_{0}$, then $\hat{x}$ is conjugate to $\Gamma_{1}^{ka_{1}},\ldots,\Gamma_{s}^{ka_{s}},I_{ke}]$.
We finish this section by very briefly commenting on the semisimple involutions in ${\rm PGL}_{n}(q)$ and we refer the reader to [22, Sections 3.2.2, 3.3.2, 3.4.2, 3.5.2] for a detailed analysis of semisimple involutions in all classical almost simple groups of Lie type. So now let $G_{0}={\rm PGL}_{n}(q)$ and take $x$ to be a semisimple involution. For every $i$ such that $1\leqslant i\leqslant n/2$ we get a unique class of involutions with representative $t_{i}$, where $t_{i}=\hat{t}Z$ and $\hat{t}$ is an involution in ${\rm GL}_{n}(q)$ whose (1)(-1)-eigenspace has dimension $i$. If $x\in t_{i}^{G}$, then we will say that it is a $t_{i}$-involution. Finally, if $n$ is even, then we get an additional class of involutions with representative $t^{\prime}_{n/2}$. This class of involutions arises from the quadratic field extension $\mathbb{F}_{q^{2}}/\mathbb{F}_{q}$ and we direct the reader to [22, 3.2.2.2] for more details. The precise structure of $C_{G_{I}}(x)$ is recorded in [22, Table B.1].
Unipotent elements
We now shift our attention to unipotent elements in finite classical groups. Let $x\in G$ be an element of order $p$. Then [22, 3.1.3] implies that $x$ again lifts to a unique element $\hat{x}$ of order $p$ in the corresponding matrix group, so that $x=\hat{x}Z$, where $Z$ again denotes the centre of this matrix group. We may now identify $\hat{x}$ with its Jordan decomposition on $V$ and write
$\hat{x}=[J_{p}^{a_{p}},\ldots,J_{1}^{a_{1}}],$
where each $J_{i}$ is a unipotent Jordan block of size $i$ and $a_{i}$ denotes the multiplicity of $J_{i}$ in $\hat{x}$. Note that if $G$ is symplectic or orthogonal, then there are restrictions on the multiplicities of odd and even Jordan blocks respectively. In particular, combining [22, 3.4.1, 3.5.1] we get:
Lemma 3.7.
Let $\hat{x}\in{\rm Sp}_{n}(q)$ (respectively ${\rm O}^{\epsilon}_{n}(q)$) be an element of odd order $p$ with Jordan decomposition $J_{p}^{a_{p}},\ldots,J_{1}^{a_{1}}]$ on $V$. Then $a_{i}$ is even for all odd (respectively even) $i$.
Now [22, 3.1.14] implies that two elements in ${\rm GL}_{n}(q)$ are conjugate if and only if they have the same Jordan decomposition on $V$, and the uniqueness of the lift $\hat{x}$ implies that we again have a complete description of unipotent classes in ${\rm PGL}_{n}(q)$. There are similar descriptions in the unitary, symplectic, and orthogonal cases and we direct the reader to [22, Sections 3.2.3, 3.3.3, 3.4.3, 3.4.4, 3.5.3, 3.5.4] for more information.
The main difference in the symplectic and orthogonal case worth highlighting, is when $p=2$, in which case all unipotent elements have Jordan decomposition $J_{2}^{s},J_{1}^{n-2s}]$ and there are at least two $G_{0}$-classes of elements with this decomposition when $s$ is even.
Finally, we will often be interested in the number and size of $G_{0}$-classes of unipotent elements, so we will often appeal to lemmas in [22] that determine in how many $G_{0}$-classes a given $G_{I}$-class splits. For example, if $G_{0}={\rm PGL}_{n}^{\epsilon}(q)$, then [22, 3.2.7, 3.3.10] tell us that the $G_{I}$-class of an element $x\in G_{I}$ of order $p$ with Jordan decomposition $J_{p}^{a_{p}},\ldots,J_{1}^{a_{1}}]$ splits into $(\nu,q-\epsilon)$ distinct $G_{0}$-classes, where $\nu=\gcd\{j\,:\,a_{j}>0\}$. Similar results are given in [22, 3.4.12, 3.5.14] for symplectic and orthogonal groups.
Finally, if $x$ is a field, graph, or graph-field automorphism, then detailed information can be found in [22, Chapter 3] and the structure of $C_{G_{I}}(x)$ is recorded in the relevant tables of [22, Appendix B]. We will again be particularly interested in the number of $G_{0}$-classes each $G_{0}$-class splits into and we will often refer to the relevant lemmas in [22]. For example, in the linear case, we will repeatedly refer to [22, 3.2.9, 3.2.14, 3.2.15].
We record [16, 3.50] that will be useful to us when dealing with field, graph and graph-field automorphisms.
Lemma 3.8.
Suppose that $G_{0}\neq{\rm P\Omega}^{+}_{8}(q)$, let $H$ be a subgroup of $G$ and suppose that $x\in H\setminus{\rm PGL}(V)$ has prime order. Then $x^{G}\cap H\subseteq\widetilde{H}x$, where $\widetilde{H}=H\cap G_{I}$.
3.3. Probabilistic methods
The proof of Theorem 3.1 will heavily rely on fixed point ratio bounds established in a series of papers by Burness [14, 15, 16, 17]. We now state the main result of those papers, which is Theorem 1 in [14].
Theorem 3.9.
Let $G$ be a finite almost simple classical group over $\mathbb{F}_{q}$ and let $H$ be a non-standard subgroup of $G$. Then
$\mbox{{\rm fpr}}(x,G/H)<|x^{G}|^{-\frac{1}{2}+\frac{1}{n}+\iota(G,H)}$
for all $x\in G$ of prime order, where $\iota(G,H)\geqslant 0$ is a known constant.
Note that in most cases $\iota(G,H)=0$. The precise value of $\iota(G,H)$ when $\iota(G,H)>0$ is listed in [14, Table 1]. Note that in most cases the term $n$ appearing in Theorem 3.9 is the dimension of the natural module, unless $G_{0}={\rm Sp}_{4}(2)^{\prime}$ or ${\rm L}_{3}(2)$, in which case $n=2$ (since ${\rm Sp}_{4}(2)^{\prime}\cong{\rm L}_{2}(2)$ and ${\rm L}_{3}(2)\cong{\rm L}_{2}(7)$). We will refer to $n$ as the dimension of $G$.
Following [61], we now define an ‘zeta-type function’ associated to a finite group $G$, which will be a key tool in the proof of our main theorem for high-dimensional classical groups.
Definition 3.10.
Let $G$ be a finite almost simple group with socle $G_{0}$, and let $\mathscr{C}$ be the set of $G_{0}$-conjugacy classes of prime order elements in $G$. For $t\in\mathbb{R}$, we define:
$\eta_{G}(t)=\sum_{C\in\mathscr{C}}|C|^{-t}$
Evidently, there exists a real number $T_{G}\in(0,1)$ such that $\eta_{G}(T_{G})=1$.
Results on the asymptotic behaviour of $\eta_{G}(t)$ can be extracted from [61, 1.10]. Note that the function considered in [61] is rather different to $\eta_{G}(t)$. In particular, Liebeck and Shalev sum over all $G$-conjugacy classes, whilst we are summing over all prime order $G_{0}$-classes. However, as we explain below, it is still possible to extract asymptotic results for our purposes:
Lemma 3.11.
We have $\sum_{i=1}^{r}|x_{i}^{G}|^{-t}\leqslant\eta_{G}(t)$ for any almost simple group $G$ with socle $G_{0}$ and $t\in(0,1)$, where $x_{1},\ldots,x_{r}$ is a set of representatives of $G$-conjugacy classes of prime order elements in $G$.
Proof.
Let $x_{1},\ldots,x_{s}\in G$ be a set of representatives of the prime order $G_{0}$-classes, and let $c_{i}$ be the number of $G_{0}$-classes in which $x_{i}^{G}$ splits into. Then
$\sum_{i=1}^{r}|x_{i}^{G}|^{-t}\leqslant\sum_{i=1}^{r}c_{i}\left(\frac{|x_{i}^{G}|}{c_{i}}\right)^{-t}=\sum_{j=1}^{s}|x_{j}^{G_{0}}|^{-t}=\eta_{G}(t)$
as required. ∎
Remark 3.12.
Note that the proof of Lemma 3.11 does not rely at all on the assumption that the elements we are summing over have prime order. In fact, by carefully inspecting [61, 1.10], we note that Liebeck and Shalev use the exact same observation in their proof and work in terms of $G_{0}$-classes. This allows us to extract the following asymptotic result:
Proposition 3.13.
Let $G=G(q)$ be an almost simple group with socle $G_{0}$ over $\mathbb{F}_{q}$. Then the following hold:
  • (i)
    $\eta_{G}(t)\rightarrow 0$ as $q\rightarrow\infty$ for any $t>T_{G}$.
  • (ii)
    For a fixed real number $t>0$, there exists some integer $r(t)$, such that for all almost simple groups $G=G(q)$ with rank $r\geqslant r(t)$ we have
    $\eta_{G}(t)\rightarrow 0\text{ as }|G|\rightarrow\infty.$
Proof.
We will prove (i). The argument for (ii) is identical. If $\mathscr{S}$ denotes the set of $G_{0}$-conjugacy classes in $G$, then inspecting [61, 1.10] gives
$\zeta_{G}(t):=\sum_{C\in\mathscr{S}}|C|^{-t}\rightarrow 1\,\,\text{ as }\,\,\rightarrow\infty$
and noting that $\eta_{G}(t)<\zeta_{G}(t)-1$, the result follows. ∎
We now prove a variant of [18, 2.2], which will allow us to use $\eta_{G}$ to establish bounds on $R_{\rm ns}(G)$ for a finite classical almost simple group $G$.
Proposition 3.14.
Let $G$ be a finite almost simple classical group and let $\tau=(H_{1},\ldots,H_{k})$ be a core-free tuple. Suppose that the following hold:
  • (i)
    ${\rm fpr}(x,G/H_{i})<|x^{G_{0}}|^{-m}$ for some some $m>0$ for all $x\in G$ of prime order;
  • (ii)
    $T_{G}<km-1$.
Then $\tau$ is regular.
Proof.
In view of Lemma 2.2, it suffices to show that $\widehat{Q}(G,\tau)<1$. If $x_{1},\ldots,x_{k}$ are representatives of the $G_{0}$-conjugacy classes of prime order elements in $G$, then Lemma 2.2 implies that
$\widehat{Q}(G,\tau)\leqslant\sum_{i=1}^{k}|x_{i}^{G_{0}}|^{-km+1}=\eta_{G}(km-1).$
Finally, noting that $\eta_{G}(t)<1$ for all $t>T_{G}$, the claim follows. ∎
We now record some bounds on $T_{G}$ established in [18, Proposition 2.2, Remark 2.3] that will be key to proving the main result for almost simple classical groups of large dimension.
Proposition 3.15.
Let $G$ be a finite almost simple classical group of rank $n$ with socle $G_{0}$. If $n\geqslant 6$, then $T_{G}<1/3$. Moreover, if $G_{0}={\rm PSL}_{10}^{\epsilon}(q)$, then $T_{G}<1/5$ and if $G_{0}={\rm P\Omega}_{n}^{\epsilon}(q)$ and $n\geqslant 12$, then $T_{G}<4/15$.
Remark 3.16.
Burness again works with a slightly different function than $\eta_{G}(t)$. In particular, he sums over $G$-classes of prime order elements instead of $G_{0}$-classes. However, by closely inspecting his proofs, we note that he also makes use of the observation in Lemma 3.11, and so the bounds in [18, Proposition 2.2, Remark 2.3] still apply in our setting.
We now record a result on fixed point ratios of almost simple classical groups, which is an immediate consequence of [55, Theorem 1]. Note that the original result concerns all the almost simple classical groups, but we will only need the result for groups with socle ${\rm PSL}_{2}(q)$.
Proposition 3.17.
Let $G$ be an almost simple group with socle ${\rm PSL}_{2}(q)$, where $q=p^{f}$, $H$ be a core-free subgroup of $G$, and $x\in G$ an element of prime order $r$. Then
${\rm fpr}(x,G/H)<\frac{4}{3q}$
or ${\rm fpr}(x,G/H)<\alpha$ and $(H,x,\alpha)$ is recorded in Table 4.
[…]
Table 4. The exceptional cases in Proposition 3.17.
In view of Lemma 2.4, Lemma 3.11 and Remark 2.5, we will often need to refer to results in the literature on lower bounds on $|x^{G}_{0}|$. In particular, if $G_{0}$ is in the following collection:
$\mathcal{A}=\{{\rm PSL}^{\epsilon}_{3}(q),{\rm PSL}_{4}^{\epsilon}(q),{\rm PSL}_{5}^{\epsilon}(q),{\rm PSp}_{4}(q)\}$
then we will repeatedly appeal to results in [15] on conjugacy class bounds. In the following lemma we record the bounds that are of interest to us.
Lemma 3.18.
Suppose that $G_{0}\in\mathcal{A}$ and $x\in G$ is an element of prime order $r$. Then $|x^{G_{0}}|>\alpha$, where $\alpha$ is recorded in Table 5 and $\beta=\frac{q}{q+1}$.
Proof.
Sufficient upper bounds on $|x^{G_{I}}|$ (or $|x^{G_{0}}|$ if $x$ is a field, graph, or graph-field automorphism) are established in [16, Sections 3.3-3.5]. Combining those with [22, 3.2.8, 3.2.9, 3.2.14, 3.2.15, 3.3.10, 3.3.12, 3.3.17, 3.4.12, 3.4.15, 3.4.16, 3.5.20], which describe when prime order $G_{I}$-classes split into $G_{0}$-classes, we verify the entries in Table 5. The argument is similar in most cases, so we only treat a few cases that are slightly more involved here.
First assume that $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ and $x\in G_{I}$. Note that [22, 3.2.8, 3.3.10] imply that $|x^{G_{0}}|=|x^{G_{I}}|$, unless $x$ is unipotent with Jordan decomposition $J_{3}]$, in which case $|x^{G_{0}}|\geqslant\frac{1}{3}|x^{G_{I}}|$. Therefore, if $x$ is unipotent, then it follows from [16, 3.22] that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$. Moreover, if $s$ denotes the codimension of the largest eigenspace of the lift $\hat{x}$ of $x$ in ${\rm GL}(V)$, then we note that $s=2$ unless $x$ has Jordan form $J_{2},J_{1}]$, and so with this exception we obtain the better bound
$|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{6}.$
Next suppose that $x$ is semisimple. If $C_{\bar{G_{I}}}(x)$ is connected, where $\bar{G_{I}}$ denotes the ambient algebraic group of $G$, then the bounds in Table 5 are verified via [16, 3.36]. On the other hand, if $C_{\bar{G_{I}}}(x)$ is non-connected, then [15, 3.35] implies that $r$ divides $3$, and so $r=3$, and moreover
$|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6},$
as claimed.
Now assume that $x$ is a field or graph-field automorphism. If $r=2$, then
$|x^{G_{0}}|\geqslant\frac{|{\rm PSL}_{3}^{\epsilon}(q)|}{|{\rm PGU}_{3}(q^{1/2})|}>\frac{1}{2(3,q-\epsilon)}\left(\frac{q}{q+1}\right)q^{(3^{2}-1)/2}>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{4}.$
Similarly, if $r\geqslant 3$, then we obtain $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{16/3}$.
Now assume that $G_{0}={\rm PSp}_{4}(q)$ and $x\in G_{I}$. Appealing to [16, 3.21] we find that $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$ if $x$ is unipotent, and in particular $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{6}$ if $x$ does not have Jordan decomposition $J_{2},J_{1}^{2}]$ on $V$.
Similarly, if $x$ is semisimple and $s$ denotes the codimension of the largest eigenspace of the lift $\hat{x}$ of $x$ on ${\rm Sp}_{4}(q)$, then we note that $s>1$, and so [16, 3.36-3.37] immediately gives us
$|x^{G_{0}}|=|x^{G_{I}}|>\frac{1}{4}q^{4}$
and unless $x$ is a $t_{1}$ or $t_{1}^{\prime}$ involution, then we obtain the better bound $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$ via [16, 3.30, 3.37].
We obtain all other bounds in Table 5 in a similar fashion. ∎
[…]
Table 5. The bounds from Lemma 3.18.
4. Fixed point ratios for classical groups
In this section we improve the fixed point ratio bounds obtained in [14, 15, 16, 17] for some low-dimensional classical groups. The goal here is to obtain sharper bounds that will allow us to use the $\eta$ function machinery discussed in Definition 3.10 and Propositions 3.13 and 3.14 in some low-dimensional cases, where a more direct approach was taken in [18] to derive results on base sizes.
Recall that $G$ is a finite almost simple classical group with socle $G_{0}$ over $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$. Moreover, $V$ denotes the natural module of $G_{0}$ and we will write $H$ for a non-standard subgroup of $G$. We will improve the fixed point ratio bounds in Theorem 3.9 only for certain families of classical groups. The following list contains the possible socle types of the groups we consider and we will treat each possibility in turn:
$\mathcal{B}=\{{\rm PSp}_{8}(q),{\rm PSL}^{\epsilon}_{8}(q),{\rm P\Omega}_{10}^{\epsilon}(q),{\rm PSL}_{6}^{\epsilon}(q),{\rm PSp}_{6}(q),{\rm\Omega}_{7}(q),{\rm PSL}_{5}^{\epsilon}(q),{\rm PSL}_{4}^{\epsilon}(q),{\rm PSL}_{3}^{\epsilon}(q)\}$
We will prove the following theorem:
Theorem 4.1.
If $G_{0}\in\mathcal{B}$, then ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-\beta}$, where $\beta$ is recorded in Table 6.
Remark 4.2.
We note that our bounds have a slightly weaker form than the ones given by Burness in [14, 15, 16, 17]. In particular, he bounds ${\rm fpr}(x,G/H)$ in terms of the $G$-conjugacy class of $x$, whilst we obtain an estimate in terms of the $G_{0}$-class of $x$. However, note that our different definition of $\eta_{G}$ allows us to work with this more relaxed bound.
We will divide the proof of Theorem 4.1 in a series of propositions, according to the socle type of $G$, where we will consider the non-standard subgroups of $G$, for which the claim does not follow immediately from Theorem 3.9. We briefly outline our general approach:
  • (i)
    In the majority of cases we first verify that the claim holds using Magma for small values of $q$, where our general approach is not applicable, using the methods described in Section 2.2.
  • (ii)
    We next consider elements $x\in G_{I}$. To derive an upper bound on ${\rm fpr}(x,G/H)$ we appeal to Lemma 2.4 and Remark 2.5, which allow us to work with the estimate
    ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{I}}\cap H|}{|x^{G_{0}}|}$
    so our first step is to derive an upper bound on $|x^{G_{I}}\cap H|$ and a lower bound on $|x^{G_{0}}|$, which we either do directly, or we inspect the proofs of [14, 15, 16, 17], and we use the bounds derived there when those are sufficient.
  • (ii)
    We then note that $\frac{|x^{G_{I}}\cap H|}{|x^{G_{0}}|}<|x^{G_{0}}|^{-\beta}$ if and only if $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<1-\beta$, so we work with this bound and verify that the claim in Theorem 4.1 holds.
  • (iii)
    We then consider graph, field and graph-field automorphisms. In this case, the estimates in the proofs of [14, 15, 16, 17] are sufficient in most cases, and when they are not, then we work in a similar fashion as Burness to derive tighter bounds.
[…]
Table 6. The value of $\beta$ in Theorem 4.1.
Proposition 4.3.
Suppose that $G_{0}={\rm PSp}_{8}(q)$, $q\geqslant 3$ and $x\in G$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-1/3}$.
Proof.
If $H$ is not of type ${\rm Sp}_{4}(q^{2})$ or ${\rm Sp}_{4}(q)\wr S_{2}$, then the claim follows by Theorem 3.9, so it remains to consider those two cases. The claim can easily be checked using Magma for $q=3$ using the function CheckFpr() provided in Section 2.2, so for the remainder of the proof we may assume that $q\geqslant 4$.
We first consider the case where $x$ is semisimple. First suppose that $H$ is of type ${\rm Sp}_{4}(q^{2})$. Let $i$ be minimal such that $r$ divides $q^{i}-1$ and $i_{0}$ be minimal such that $r$ divides $q^{2i_{0}}-1$, and note that $i_{0}=i/2$ if $i$ is even and $i_{0}=i$ otherwise. Note that we may assume that $x\in H$, since otherwise ${\rm fpr}(x,G/H)=0$. In this case $r$ must divide $|H|$, and so we have $i\in\{1,2,4,8\}$. We need to consider each possibility of $i$ in turn. The calculations for each value of $i$ are straightforward and similar, and so we will only consider the cases where $i=4$ and $i=1$ here and leave the rest for the reader to check.
First suppose that $i=4$. In this case $i_{0}=2$ and $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, ${\rm GU}_{1}(q^{2})^{2}$, or ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$. If $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, then we compute
$|x^{G_{I}}\cap H|=2\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm GU}_{2}(q^{2})|}+\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm GU}_{1}(q^{2})|}=2q^{6}(q^{4}+1)(q^{2}-1)+q^{8}(q^{2}-1)^{2}(q^{4}+1)$
where $G_{I}={\rm Inndiag}(G_{0})$ and
$|x^{G_{0}}|=\frac{|{\rm Sp}_{8}(q)|}{|{\rm GU}_{2}(q^{2})|}=q^{14}(q^{2}-1)^{2}(q^{4}+1)(q^{6}-1),$
and so the claim follows.
If $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})^{2}$, then we have
$|x^{G_{I}}\cap H|\leqslant 4\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm GU}_{1}(q^{2})|^{2}}=4q^{8}(q^{2}-1)^{2}(q^{4}+1)$
and
$|x^{G_{0}}|=\frac{|{\rm Sp}_{8}(q)|}{|{\rm GU}_{1}(q^{2})|^{2}}=q^{16}(q^{2}-1)^{3}(q^{4}+1)(q^{6}-1),$
and so ${\rm fpr}(x,G/H)<|x^{G}|^{-1/3}$ here as well.
Finally, if $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$, then we compute
$|x^{G_{I}}\cap H|\leqslant 2\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm GU}_{1}(q^{2})||{\rm Sp}_{2}(q^{2})|}=2q^{6}(q^{4}+1)(q^{2}-1)$
and
$|x^{G_{0}}|=\frac{|{\rm Sp}_{8}(q)|}{|{\rm GU_{1}}(q^{2})||{\rm Sp}_{4}(q)|}=q^{12}(q^{2}-1)(q^{6}-1)(q^{4}+1)$
and we observe that the desired bound is satisfied.
We finally consider the case $i=i_{0}=1$. Note that we have assumed that $x\in H$ and under this assumption we claim that $C_{G}(x)$ must be of type ${\rm GL}_{4}(q)$, ${\rm GL}_{2}(q)\times{\rm Sp}_{4}(q)$, or ${\rm GL}_{2}(q)^{2}$. Let $\hat{x}$ denote the lift of $x$ in ${\rm Sp}_{4}(q^{2})$. Recall from Section 3.2 that two elements of order $r$ are ${\rm Sp}_{4}(q^{2})$-conjugate if and only if they have the same eigenvalues in $\mathbb{F}_{q^{2i_{0}}}=\mathbb{F}_{q^{2}}$. Moreover, we have seen again in Section 3.2 that since $i_{0}=1$ the eigenvalues of $\hat{x}$ come in inverse pairs $(\lambda,\lambda^{-1})$. Thus, if $\hat{x}$ denotes the lift of $x$ in ${\rm Sp}_{4}(q^{2})$, then we may write $\hat{x}=[\lambda,\lambda^{-1},\mu,\mu^{-1}]$ with $\lambda,\mu\in\mathbb{F}_{q^{2}}$ not necessarily distinct.
As explained in Section 3.2, the equality $i=i_{0}$ implies that $\sigma$ and $\sigma_{0}$-orbits coincide, and in particular, Lemma 3.6 implies that $\hat{x}$ embeds as $(\lambda,\lambda^{-1})^{2},(\mu,\mu^{-1})^{2}]$ in $G$, so all eigenvalues occur with even multiplicity. It follows that $C_{G}(x)$ cannot be of type ${\rm GL}_{1}(q)^{4}$ or ${\rm GL}_{3}(q)\times{\rm Sp}_{2}(q)$, and the claim follows.
We continue by considering each possible centraliser type for $x$ in turn. All cases are entirely similar so we only consider the case where $C_{G}(x)$ is of type ${\rm GL}_{2}(q)\times{\rm Sp}_{4}(q)$ and leave the remaining cases to the reader.
In this case we compute
$|x^{G_{I}}\cap H|=\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm GL}_{1}(q^{2})||{\rm Sp}_{2}(q^{2})|}=q^{6}(q^{2}+1)^{2}(q+1)(q^{4}+1)$
and
$|x^{G_{0}}|=\frac{|{\rm Sp}_{8}(q)|}{|{\rm GL}_{2}(q)||{\rm Sp}_{4}(q)|}=q^{11}(q^{4}+q^{2}+1)(q^{4}+1)(q^{2}+1)(q+1),$
and so the claim holds.
Now suppose that $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$ and again let $i$ be minimal such that $r|q^{i}-1$. We may again assume that $x\in H$, and in particular that $r$ divides $|H|$, since otherwise we have ${\rm fpr}(x,G/H)=0$. It follows that $i\in\{1,2,4\}$. The computations are straightforward here as well, and so, to avoid unnecessary repetition, we only consider the case $i=2$ and leave the remaining cases to the reader.
If $i=2$, then there are $11$ possibilities for $C_{G}(x)$ up to isomorphism. More precisely, $C_{G}(x)$ is one of the following types:
${\rm GU}_{4}(q),{\rm GU}_{3}(q)\times{\rm GU}_{1}(q),{\rm GU}_{3}(q)\times{\rm Sp}_{2}(q),{\rm GU}_{2}(q)^{2},{\rm GU_{2}}(q)\times{\rm GU}_{1}(q)^{2},$
${\rm GU}_{2}(q)\times{\rm GU}_{1}(q)\times{\rm Sp}_{2}(q),{\rm GU}_{2}(q)\times{\rm Sp}_{4}(q),{\rm GU}_{1}(q)^{4},{\rm GU}_{1}(q)^{3}\times{\rm Sp}_{2}(q),$
${\rm GU}_{1}(q)^{2}\times{\rm Sp}_{4}(q),{\rm GU}_{1}(q)\times{\rm Sp}_{6}(q).$
All cases can be handled in a similar way. For example, let us consider the case where $C_{G}(x)$ is of type ${\rm GU}_{2}(q)^{2}$. As described in Section 3.2, if $\hat{x}$ denotes the lift of $x$ in ${\rm Sp}_{8}(q)$, then $\hat{x}$ is ${\rm Sp}_{8}(q)$-conjugate to an element of the form $\Lambda_{1}^{2},\Lambda_{2}^{2}]$ for two distinct $\sigma$-orbits $\Lambda_{1},\Lambda_{2}$.
Let $\widehat{H}={\rm Sp}_{4}(q)\wr S_{2}$ denote the lift of $H$ in ${\rm GSp}_{8}(q)$ and $V_{1}\oplus V_{2}$ denote the decomposition of $V$ into $4$-dimensional $\mathbb{F}_{q}$-spaces $V_{1},V_{2}$ preserved by $\hat{H}$. Note that since $r$ is odd, we must have $\hat{x}\in\widehat{B}={\rm Sp}_{4}(q)^{2}$. Now there are two $\widehat{H}-$classes of elements that fuse to the ${\rm GSp}_{8}(q)$-class of $\hat{x}$, namely one with representative $\hat{y},\hat{y}$], where $\hat{y}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1},\Lambda_{2}]$ and the one with representative $\hat{z},\hat{w}$], where $\hat{z}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1}^{2}]$ and $\hat{w}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{2}^{2}]$. It now remains to compute the sizes of the $\widehat{H}-$classes of $\hat{y},\hat{y}$] and $\hat{z},\hat{w}$] to obtain a bound on $|x^{G_{I}}\cap H|$. It is easy to see that the $\widehat{H}-$class of $\hat{y},\hat{y}$] consists of all elements of the form $\hat{a},\hat{b}$], with $\hat{a},\hat{b}$ $\widehat{H}-$conjugate to $\hat{y}$, and so
$|[\hat{y},\hat{y}]^{\widehat{H}}|=\frac{|{\rm Sp}_{4}(q)|^{2}}{|{\rm GU}_{1}(q)|^{4}}.$
On the other hand, if $\hat{a},\hat{b}$] is $\widehat{H}-$conjugate to $\hat{z},\hat{w}$], then since $\widehat{H}$ can permute $V_{1}$ and $V_{2}$ either $\hat{a}\in\hat{z}^{\widehat{H}}$ and $\hat{b}\in\hat{w}^{\widehat{H}}$, or $\hat{a}\in\hat{w}^{\widehat{H}}$ and $\hat{b}\in\hat{z}^{\widehat{H}}$, and so
$|[\hat{z},\hat{w}]^{\widehat{H}}|=2\left(\frac{|{\rm Sp}_{4}(q)|}{|{\rm GU}_{2}(q)|}\right)^{2}.$
It follows that
$|x^{G_{I}}\cap H|=2\left(\frac{|{\rm Sp}_{4}(q)|}{|{\rm GU}_{2}(q)|}\right)^{2}+\frac{|{\rm Sp}_{4}(q)|^{2}}{|{\rm GU}_{1}(q)|^{4}}=2q^{6}(q-1)^{2}(q^{2}+1)^{2}+q^{8}(q-1)^{4}(q^{2}+1)^{2}$
and moreover,
$|x^{G_{0}}|=\frac{|{\rm Sp}_{8}(q)|}{|{\rm GU}_{2}(q)|^{2}}=q^{14}(q-1)^{2}(q^{2}+1)^{2}(q^{4}+q^{2}+1)(q^{4}+1),$
so ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-1/3}$.
Now suppose that $x$ is a unipotent element of odd prime order. If $H$ of type ${\rm Sp}_{4}(q^{2})$, then the possible Jordan forms on $V$ for an element $x\in H$ are $J_{4}^{2}],[J_{2}^{4}]$, and $J_{2}^{2},J_{1}^{4}]$. We can easily compute $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using the information in [22, Table B.7] in every case and deduce that the desired bound holds. For example, if $x$ has Jordan form $J_{2}^{2},J_{1}^{4}]$, then we have
$|x^{G_{I}}\cap H|\leqslant\frac{2|{\rm Sp}_{4}(q^{2})|}{q^{6}|{\rm Sp}_{2}(q^{2})|}=2(q^{8}-1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{8}(q)|}{q^{11}|{\rm O}^{-}_{2}(q)||{\rm Sp}_{4}(q)|}=\frac{1}{2}q(q-1)(q^{2}+1)(q^{4}+1)(q^{6}-1)$
and those bounds are sufficient.
If $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$, then the possibilities and bounds for $|x^{G}\cap H|$ and $|x^{G}|$ are listed in the proof of [18, 3.2], and one can check that those bounds are sufficient, so we omit the details here.
We now consider the case where $x$ is an involution. If $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$, then one can easily check that the desired bound holds by inspecting the proof of [16, 2.8, 2.9], so we now assume that $H$ is of type ${\rm Sp}_{4}(q^{2})$.
If $x$ is not semisimple with $C_{G}(x)$ of type ${\rm Sp}_{4}(q)^{2}$, then the bounds found in the proof of [16, 3.2] are sufficient, so we may now assume that $q$ is odd and $C_{G}(x)$ is of type ${\rm Sp}_{4}(q)^{2}$. By [22, Section 3.4.2.2] we have $x\in G_{0}$. Now [47, 4.3.10] tells us that
$H\cap G_{0}={\rm PSp}_{4}(q^{2}).\langle\phi\rangle=B.\langle\phi\rangle,$
where $\phi$ is a field automorphism of order 2, and so we compute
$|x^{G_{I}}\cap B|\leqslant\frac{|{\rm PSp}_{4}(q^{2})|}{|{\rm Sp}_{2}(q^{2})|^{2}}\,\,\,\,\,\text{and}\,\,\,\,\,|x^{G_{0}}|\geqslant\frac{|{\rm PSp}_{8}(q)|}{2|{\rm Sp}_{4}(q)|^{2}}.$
Moreover, (68) in [16] gives us
$|x^{G_{I}}\cap(H\setminus B)|\leqslant\frac{|{\rm Sp}_{4}(q^{2})|}{|{\rm Sp}_{4}(q)|}$
and so the claim holds.
Finally, if $x$ is a field automorphism of order $r$, then the bounds presented at the start of [16, Section 3.2] and [16, 2.3] are sufficient. ∎
Proposition 4.4.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{8}(q)$ and $x\in G$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-1/3}$.
Proof.
We only need to consider the case where $H$ is of type ${\rm Sp}_{8}(q)$, since otherwise the claim follows straight from Theorem 3.9. In this remaining case, the claim can be verified using Magma for $q\leqslant 2$, and for $q\geqslant 3$ one can argue in a similar fashion to the proof of Proposition 4.3, so we leave the details for the reader to check.
If $x$ is a field, graph, or graph-field automorphism, then the bounds in [15, 8.1] suffice, so it is left to check the claim for semisimple and unipotent elements. In those cases we follow the same arguments as in [15, 8.1]. We start by noting that $H\cap{\rm PGL}(V)\leqslant{\rm PGSp}_{8}(q)$, and so $|x^{G_{I}}\cap H|\leqslant|x^{{\rm PGSp}_{8}(q)}|$. Now the conjugacy classes in ${\rm PGSp}_{8}(q)$ and ${\rm PGL}_{8}^{\epsilon}(q)$ are discussed in Section 3.2 briefly and in [22, Chapter 7] in more detail, and the centraliser structure for each class can be read off from [22, Tables B.3 and B.7], so one can check that the claim holds. ∎
Proposition 4.5.
Suppose that $G_{0}={\rm P\Omega}^{\epsilon}_{10}(q)$ and $x\in G$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-1/3}$.
Proof.
In this case we may assume that $H$ is of type ${\rm GL}_{5}^{\epsilon}(q)$, as the claim follows by Theorem 3.9 in all other cases. The claim can be verified using Magma for $q\in\{2,3\}$, and for $q\geqslant 4$ the bounds in the proof of [16, 3.3] are sufficient. ∎
Proposition 4.6.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{6}(q)$ and $x\in G$ is an element of prime order $r$. Then either ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-1/3}$, or one of the following holds:
  1. (i)
    $\epsilon=+$, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q)^{2}$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G}|^{-0.32}$
  2. (ii)
    $\epsilon=$-, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q^{2})$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G}|^{-0.32}$.
Proof.
Here we may assume that $H$ is of type ${\rm Sp}_{6}(q)$ since the claim follows directly by Theorem 3.9 in all other cases. First assume $x$ is semisimple of odd order and let $i$ be minimal such that $r|q^{i}-1$. Again, we may assume that $x\in H$ and in particular that $r$ divides $|H|$, since otherwise we have ${\rm fpr}(x,G/H)=0$. Thus, we have $i\in\{1,2,3,4,6\}$. The embedding here is straightforward so we will only give details for the cases where $i\in\{1,4\}$ and leave the remaining cases for the reader.
First suppose $i=4$. In this case $C_{G}(x)$ is of type ${\rm GL}_{1}(q^{4})\times{\rm GL}^{\epsilon}_{2}(q)$ and we have
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GU}_{1}(q^{2})||{\rm Sp}_{2}(q)|}=q^{8}(q^{2}-1)(q^{6}-1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{6}(q)|}{|{\rm GL}_{1}(q^{4})||{\rm GU}_{2}(q)|}=q^{14}(q-1)^{2}(q^{3}-1)(q^{5}-1)(q^{4}+q^{2}+1).$
The claim follows.
Now assume $i=1$. If $\epsilon=+$, then there are six possibilities for $C_{G}(x)$ for an element $x\in H$. More precisely, $C_{G}(x)$ can be of type ${\rm GL}_{3}(q)^{2}$, ${\rm GL}_{2}(q)^{2}\times{\rm GL}_{1}(q)^{2}$, ${\rm GL}_{2}(q)^{3}$, ${\rm GL}_{2}(q)\times{\rm GL}_{1}(q)^{4}$, ${\rm GL}_{1}(q)^{6}$, or ${\rm GL}_{4}(q)\times{\rm GL}_{1}(q)^{2}$. All cases can be handled in a similar way. For example, if $C_{G}(x)$ is of type ${\rm GL}_{3}(q)^{2}$, then we find
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GL}_{3}(q)|}=q^{6}(q+1)(q^{2}+1)(q^{3}+1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{6}(q)|}{|{\rm GL}_{3}(q)|^{2}}=q^{9}(q^{4}+q^{3}+q^{2}+q+1)(q^{2}+1)(q^{3}+1)$
and one can check that $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<0.68$.
Similarly, if $C_{G}(x)$ is of type ${\rm GL}_{2}(q)^{2}\times{\rm GL}_{1}(q)^{2}$, then we have
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GL}_{2}(q)||{\rm GL}_{1}(q)|}=q^{8}(q+1)^{2}(q^{2}-1)(q^{4}+q^{2}+1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{6}(q)|}{|{\rm GL}_{2}(q)|^{2}|{\rm GL}_{1}(q)|}=q^{13}(q+1)(q^{2}+q+1)(q^{2}+1)(q^{4}+q^{3}+q^{2}+q+1)(q^{4}+q^{2}+1)$
and so the claim is true in this case as well. We leave the remaining cases for the reader to check.
Now suppose that $\epsilon=$-. In this case the possible centraliser types for a prime order element $x\in H$ are ${\rm GL}_{3}(q^{2})$, ${\rm GL}_{2}(q^{2})\times{\rm GL}_{1}(q^{2})$, ${\rm GL}_{2}(q^{2})\times{\rm GU}_{1}(q)$, ${\rm GL}_{1}(q^{2})^{2}\times{\rm GU}_{2}(q)$, ${\rm GL}_{1}(q^{2})^{3}$, and ${\rm GL}_{1}(q^{2})\times{\rm GU}_{4}(q)$, and again we can verify that the desired bound holds by treating each possibility in turn. For example, if $C_{G}(x)$ is of type ${\rm GL}_{3}(q^{2})$, then
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GL}_{3}(q)|}=q^{6}(q+1)(q^{2}+1)(q^{3}+1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GU}_{6}(q)|}{|{\rm GL}_{3}(q^{2})|}=q^{9}(q+1)(q^{3}+1)(q^{5}+1)$
and one can check that $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<0.68$.
Now suppose that $x$ is unipotent. By appealing to Section 3.2, we find that the Jordan decompositions corresponding to an element $x\in H$ are
$J_{6}],[J_{4},J_{2}],[J_{4},J_{1}^{2}],[J_{3}^{2}],[J_{2}^{3}],[J_{2}^{2},J_{1}^{2}],[J_{2},J_{1}^{4}],$
so we check each possibility in turn to confirm that the required bound holds. For example, if $x$ is of type $J_{2},J_{1}^{4}]$, then we get
$|x^{G_{I}}\cap H|\leqslant\frac{2|{\rm Sp}_{6}(q)|}{q^{5}|{\rm Sp}_{4}(q)|}\leqslant 2(q^{6}-1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GU}_{6}(q)|}{q^{9}|{\rm GU}_{4}(q)||{\rm GU}_{1}(q)|}=(q^{4}-q^{3}+q^{2}-q+1)(q^{6}-1)$
and hence ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-1/3}$. The other cases can be dealt with in a similar fashion by inspecting  [22, Tables B.3 and B.7], so we omit the details here.
Finally, if $x$ is a graph field or graph field automorphism or a semisimple involution, then the bounds in the proof of [15, 8.1] suffice. ∎
Proposition 4.7.
Suppose that $G_{0}={\rm PSp}_{6}(q)$ $q$ is even and $x\in G$ is an element of prime order $r$. Then either $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-0.4}$, or $H$ is of type $G_{2}(q)$.
Proof.
The claim can be easily verified using Magma for $q\in\{2,4\}$, and so we may assume that $q\geqslant 8$ for the remainder of the proof. By inspecting [12, Tables 8.28 and 8.29], we find that if $H$ is not of type $G_{2}(q)$, then $H_{0}=H\cap G_{0}\in\{{\rm Sp}_{2}(q)\wr S_{3},{\rm Sp}_{2}(q^{3}){:}3,{\rm Sp}_{6}(q_{0})\}$, where $k$ is a prime and $q=q_{0}^{k}$.
First suppose that $H_{0}={\rm Sp}_{2}(q)\wr S_{3}=B.S_{3}$. If $x$ is a field automorphism of order $r$, then the bounds in [16, Section 2.3] give
$|x^{G}\cap H|<48q^{9(1-1/r)}\,\,\,\,\,\text{and}\,\,\,\,\,|x^{G}|>\frac{1}{4}q^{21(1-1/r)}$
and so the claim holds.
Now suppose that $x\in{\rm PGL}(V)$. There are two cases to consider depending whether or not $x^{G}\cap(H\setminus B)\neq\emptyset$. If $x^{G}\cap(H\setminus B)\neq\emptyset$, then either $r=2$ or $r=3$. If $r=2$, then by inspecting the proof of [16, 2.8], we deduce that $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<0.543$, so we may now assume that $r=3$. In this case we may assume $x=b\pi=(x_{1},x_{2},x_{3})\pi$, where $b\in B$ and $\pi\in S_{3}$ is a $3$-cycle and it follows from the proof of [63, 4.5] that $x$ is $B$-conjugate to $\pi$. Now note that if $y=(A,B,C)\in C_{B}(x)$, then since $\pi$ permutes $A,B$, and $C$, we must have $A=B=C$, and so
$|x^{G_{I}}\cap H|=\frac{6|{\rm Sp}_{2}(q)|^{3}}{3|{\rm Sp}_{2}(q)|}=2q^{2}(q^{2}-1)^{2}.$
We now note that with respect to a symplectic basis $\{e_{1},e_{2},e_{3},f_{1},f_{2},f_{3}\}$, $(I_{2},I_{2},I_{2})\pi$ embeds in ${\rm Sp}_{6}(q)$ as the block-diagonal matrix $[A,A]$, where
$A=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$
and since $r=3$ and $q$ is even, $x$ is semisimple. Moreover, since $q$ is even, $3$ must either divide q1q-1 or $q+1$, and so if $i$ is minimal such that $3$ divides $q^{i}-1$, we must have $i=1$ or $2$. Now note that $x$ has eigenvalues $1^{2},(\omega)^{2},(\omega^{2})^{2}$, where $\omega\in\mathbb{F}_{q^{i}}$ is a primitive cube root of unity, and so $C_{G}(x)$ is of type ${\rm GL}_{2}^{\epsilon}(q)\times{\rm Sp}_{2}(q)$, where $\epsilon=(-1)^{i+1}$. Therefore,
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm Sp}_{2}(q)||{\rm GU}_{2}(q)|}=q^{7}(q-1)(q^{2}+1)(q^{4}+q^{2}+1)$
and the claim holds.
The case $x^{G}\cap H\subseteq B$ is straightforward. First suppose that $x$ is semisimple and $i$ is minimal such that $r$ divides $q^{i}-1$. We may again assume that $x\in H$, in which case $r$ divides $|H|$, and we must have $i\in\{1,2\}$. If $i=2$, then $C_{G}(x)$ must be of type
${\rm GU}_{3}(q),{\rm GU}_{2}(q)\times{\rm Sp}_{2}(q),{\rm GU}_{2}(q)\times{\rm GU}_{1}(q),{\rm GU}_{1}(q)\times{\rm Sp}_{4}(q),{\rm GU}_{1}(q)^{2}\times{\rm Sp}_{2}(q),\text{ or }{\rm GU}_{1}(q)^{3}.$
We can now compute bounds for $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using [22, Table B.7]. For example, if $C_{G}(x)$ is of type ${\rm GU}_{2}(q)\times{\rm GU}_{1}(q)$, then as discussed in Section 3.2, we may assume that the lift $\hat{x}$ of $x$ in ${\rm Sp}_{6}(q)$ is $(\Lambda_{1})^{2},\Lambda_{2}]$ for some $\sigma$-orbits $\Lambda_{1},\Lambda_{2}$. Now we find that $\hat{y}=[\hat{y_{1}},\hat{y_{2}},\hat{y_{3}}]\in\hat{x}^{{\rm Sp}_{2}(q)^{3}}$ if and only if $\hat{y}_{i}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{2}]$ for some $i$ and $\hat{y}_{j}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{1}$ for $j\in\{1,2,3\}\setminus\{i\}$. Therefore,
$|x^{G_{I}}\cap H|\leqslant\frac{3|{\rm Sp}_{2}(q)^{3}|}{|{\rm GU}_{1}(q)|^{3}}=3q^{3}(q-1)^{3}$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GU}_{2}(q)||{\rm GU}_{1}(q)|}=q^{8}(q-1)^{2}(q^{2}+1)(q^{4}+q^{2}+1)$
and the claim follows. We can compute in the same way in all other cases. If $i=1$, then the argument is almost identical, so we omit the details to avoid unnecessary repetition.
Now suppose that $x\in H$ is unipotent. Since $x^{G}\cap H\subseteq B$, we deduce that $x$ must be a $b_{1},c_{2}$ or $b_{3}$ involution by consulting [22, 3.4.14]. We consider each possibility in turn by inspecting [22, Table B.7] and we obtain the bounds in the following Table 7, which allow us to verify that the desired claim holds.
[…]
Table 7. Fixed point ratio bounds for unipotent involutions arising in Proposition 4.7.
Now suppose that $H_{0}={\rm Sp}_{2}(q^{3}).3=B.3$. If $x$ is a field automorphism of order $r$, then the proof of [16, 3.2] gives
$|x^{G}\cap H|\leqslant 2q^{9(1-1/r)}\,\,\,\,\,\text{and}\,\,\,\,\,|x^{G}|>\frac{1}{4}q^{21(1-1/r)}$
and the claim follows.
Now suppose that $x\in{\rm PGL}(V)$. If $x^{G}\cap(H\setminus B)\neq\emptyset$, then $x$ acts on ${\rm soc}(H_{0})={\rm Sp}_{2}(q^{3})$ as a field automorphism of order $3$, and so we get
$|x^{G}\cap H|\leqslant 2\frac{|{\rm Sp}_{2}(q^{3})|}{|{\rm Sp}_{2}(q)|}=2q^{2}(q^{4}+q^{2}+1).$
Moreover, with respect to the symplectic basis $\{e_{1},e_{1}^{q},e_{1}^{q^{2}},f_{1},f_{1}^{q},f_{1}^{q^{2}}\}$, $x$ embeds in ${\rm Sp}_{6}(q)$ as $[A,A]$, where $A$ is the matrix defined above. By arguing as in the case $H_{0}={\rm Sp}_{2}(q)\wr S_{3}$, we get
$|x^{G}|\geqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm Sp}_{2}(q)||{\rm GU}_{2}(q)|}=q^{7}(q-1)(q^{2}+1)(q^{4}+q^{2}+1).$
Therefore, the required bound is satisfied and it now remains to deal with the case where $x^{G}\cap H\subseteq B$.
First suppose that $x$ is semisimple (so $r$ is odd), and let $i$ be minimal such that $r|q^{i}-1$ and $i_{0}$ be minimal such that $r|q^{3i_{0}}-1$. Note that $i_{0}=i/3$ if $3$ divides $i$, and $i=i_{0}$ otherwise. We may again assume that $x\in H$, so $r$ divides $|H|$, and in particular $i\in\{1,2,3,6\}$. All cases are similar, so we only consider the cases where $i=6$ and $i=2$ here.
First suppose that $i=6$, so $i_{0}=2$. In this case, $x$ is $G_{I}$-conjugate to $\Lambda]Z$ for some $\sigma$-orbit $\Lambda$, which as discussed in [22, Section 5.3.1] is a union of $3$ distinct $\sigma_{0}$-orbits, where $\sigma_{0}:\alpha\mapsto\alpha^{q^{3}}$. Therefore, $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{3})$ we compute
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{2}(q^{3})|}{|{\rm GU}_{1}(q^{3})|}=q^{3}(q^{3}-1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GU}_{1}(q^{3})|}=q^{9}(q^{2}-1)(q^{3}-1)(q^{4}-1)$
and so the claim holds.
Similarly, if $i=i_{0}=2$, then $C_{G}(x)$ is of type ${\rm GU}_{1}(q)^{3}$, and so
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{2}(q^{3})|}{|{\rm GU}_{1}(q^{3})|}=q^{3}(q^{3}-1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{6}(q)|}{|{\rm GU}_{1}(q)|^{3}}=q^{9}(q-1)^{3}(q^{2}+1)(q^{4}+q^{2}+1)$
and we deduce that ${\rm fpr}(x,G/H)<|x^{G}|^{-0.4}$ as desired.
Finally, if $x\in H$ is unipotent, then it embeds in $G$ as a $b_{3}$ involution and we compute
$|x^{G_{I}}\cap H|=q^{6}-1$
and
$|x^{G_{0}}|\geqslant q^{2}(q^{4}-1)(q^{6}-1)$
Therefore, the claim is true for all elements $x\in H$.
We are now left with the case where $H$ is of type ${\rm Sp}_{6}(q_{0})$, and $q=q_{0}^{k}$ for some prime $k$. If $x\in{\rm PGL}(V)$, then we obtain the required bound by inspecting [22, Table B.7] in a similar fashion to the previous two cases, so we leave the details to the reader.
Now assume that $x$ is a field automorphism of order $r$. If $r\neq k$, then [44, 4.9.1] implies that
$|x^{G}\cap H|\leqslant|x^{{\rm Sp}_{6}(q_{0})}|<2q_{0}^{21(1-1/r)}$
and [15, 3.48] gives
$|x^{G}|>\frac{1}{4}q_{0}^{21(1-1/r)}$
and it can be checked that those bounds are sufficient.
Suppose now that $r=k$. If $r>2$, then one can check that the trivial bound
$|x^{G}\cap H|\leqslant|{\rm Sp}_{6}(q_{0})|$
and
$|x^{G}|>\frac{|{\rm Sp}_{6}(q)|}{|{\rm Sp}_{6}(q_{0})|}$ (1)
suffice. Otherwise, if $r=k=2$, then [42, 4.3] gives us
$|x^{G}\cap H|\leqslant i_{2}({\rm Sp}_{6}(q_{0}))+1$ $\leqslant 1+\frac{|{\rm Sp}_{6}(q_{0})|}{q_{0}^{7}|{\rm Sp}_{2}(q_{0})|}+\frac{|{\rm Sp}_{6}(q_{0})|}{q_{0}^{8}|{\rm Sp}_{2}(q_{0})|}+\frac{|{\rm Sp}_{6}(q_{0})|}{q_{0}^{5}|{\rm Sp}_{4}(q_{0})|}+\frac{|{\rm Sp}_{6}(q_{0})|}{q_{0}^{6}|{\rm Sp}_{2}(q_{0})|}$
$=(q_{0}^{2}+q_{0}+1)(q_{0}^{4}-1)(q_{0}^{6}-1)+q_{0}^{6},$
where $i_{2}({\rm Sp}_{6}(q_{0}))$ denotes the number of involutions in ${\rm Sp}_{6}(q_{0})$. The result follows via (1), so the proof is complete. ∎
Proposition 4.8.
Suppose that $G_{0}={\Omega}_{7}(q)$, $q\geqslant 5$ and $x\in G$ is an element of prime order $r$. Then either $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-0.4}$, or $G$ is of type $G_{2}(q)$.
Proof.
The claim can be verified using Magma for $q=5$, so we may assume that $q\geqslant 7$ for the remainder of the proof. By inspecting [12, Tables 8.39-8.40] we see that if $H$ is not of type $G_{2}(q)$, then $H$ must be of type ${\rm O}_{1}(q)\wr S_{7},{\rm O}_{7}(q_{0})$ with $q=q_{0}^{k}$ for some prime $k$, or ${\rm Sp}_{6}(2)$. We treat each possibility in turn.
If $H$ is of type ${\rm Sp}_{6}(2)$, then it follows directly from [17, 2.15] that ${\rm fpr}(x)<|x^{G}|^{-1/2}$, and so the claim holds.
If $H$ is of type ${\rm O}_{1}(q)\wr S_{7}$, then the bounds in [16, 2.10] suffice.
Finally, if $H$ is of type ${\rm O}_{7}(q)$, then we obtain the result by inspecting the proof of [15, 5.3] and [22, Table B.12], in a similar fashion to the proof of Proposition 4.7. We omit the details. ∎
Proposition 4.9.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$ and $x\in G_{I}$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-9/20}$.
Proof.
The claim can easily be verified using Magma when $q\leqslant 5$, so for the remainder we will assume that $q\geqslant 7$. By inspecting [12, Tables 8.18-8.21], we deduce that the possible types of $H$ are the ones listed in Table 8. Moreover, it follows from Lemma 3.18 that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{8}$.
[…]
Table 8. Non-standard subgroups of almost simple groups with socle ${\rm L}_{5}^{\epsilon}(q)\,(q\geqslant 7)$.
First note that if $H$ is of type $5^{2}.{\rm Sp}_{2}(5)$, then since $q\geqslant 7$ and $q\equiv 1\allowbreak\mkern 4.0mu({\operator@font mod}\,\,5)$, the required bound holds, since $|x^{G}\cap H|\leqslant|H|$.
Similarly, if $H$ is of type ${\rm L}_{2}(11)$ or ${\rm U}_{4}(2)$, then we can construct ${\rm Aut}(H)$ using the Magma command AutomorphismGroupSimpleGroup(), and we observe that
$|x^{G}\cap H|\leqslant\max{\{i_{s}({\rm Aut}(H))\,:\,s\in\pi(|{\rm Aut}(H)|)\}}\leqslant 5184,$
where we recall that $i_{s}({\rm Aut}(H))$ denotes the number of elements of prime order $s$ in ${\rm Aut}(H)$. This is sufficient unless $q=7$ and $\epsilon=+$ and $H$ is of type $U_{4}(2)$. In this case, we observe that if $r\neq 5$, then $i_{r}({\rm Aut}(H))\leqslant 891$, which is sufficient.
On the other hand, if $r=5$, then $x$ is semisimple and the minimal $i$ such that $r|q^{i}-1$ is $4$. This implies that $C_{G}(x)$ is not of type ${\rm GL}_{4}^{\epsilon}(q)\times{\rm GL}_{1}^{\epsilon}(q)$, so $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{12}$, and the claim is again satisfied.
Now assume that $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{5}$. Then $H$ preserves a decomposition $V_{1}\oplus\cdots\oplus V_{5}$ of the natural module $V$. Recall from Lemma 3.18 that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$, unless $x$ is semisimple and $C_{G}(x)$ is of type ${\rm GL}_{s}(q)\times{\rm GL}_{n-s}(q)$ for $s\in\{3,4\}$, or $x$ is unipotent with Jordan form $J_{2}^{s},J_{1}^{5-2s}]$ for $s\in\{1,2\}$, and so if $x$ is not in those classes of prime order elements, the trivial bound
$|x^{G}\cap H|\leqslant|H|<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}\right)^{11/20}<|x^{G_{0}}|^{11/20}$ (2)
holds and we are done. We may now assume that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$. Recall that [47, 4.2.9] implies that ${\rm PGL}(V)\cap H\leqslant(q-\epsilon)^{4}.S_{5}=B.S_{5}$.
First assume that $x$ is unipotent. Then $x^{G_{I}}\cap B=\emptyset$, since $r$ does not divide $|B|$, so we may write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that [15, 3.11] implies that $x$ lifts to a unique element $\hat{x}=(\hat{x}_{1},\ldots,\hat{x}_{5})\pi\in\hat{B}\pi$ of order $r=p$, where $\widehat{B}={\rm GL}^{\epsilon}_{1}(q)^{5}$. Moreover, it is shown in the proof of [63, 4.5] that $\hat{x}$ is $\widehat{B}-$conjugate to $(1,\ldots,1,\hat{x}_{hr+1},\ldots,\hat{x}_{5})\pi$, and as demonstrated in the proof of [16, 2.6], $\hat{x}$ has Jordan form $J_{p}^{h+b_{p}},J_{p-1}^{b_{p-1}},\ldots,J_{1}^{b_{1}}]$, where $J_{p}^{b_{p}},\ldots,J_{1}^{b_{1}}]$ is the Jordan form of the restriction of $x$ on $V_{hr+1}\oplus\cdots\oplus V_{5}$.
From this we deduce that if $\hat{x}$ has Jordan form $J_{2},J_{1}^{3}]$, then $\pi$ must be a $2$-cycle, and so we compute
$|x^{G_{I}}\cap H|=\binom{5}{2}(q-\epsilon)$
and since $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{8}$, we clearly have ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-9/20}$, as claimed.
Similarly, if $\hat{x}$ has Jordan form $J_{2}^{2},J_{1}]$, then $\pi$ must be a product of two transpositions, and so
$|x^{G_{I}}\cap H|=\frac{1}{2}\binom{5}{2}\binom{3}{2}(q-1)^{3}$
which is again sufficient.
Now assume that $x$ is semisimple. Recall that under our assumption $C_{G}(x)$ is of type ${\rm GL}_{4}(q)\times{\rm GL}_{1}(q)$ or ${\rm GL}^{\epsilon}_{3}(q)\times{\rm GL}^{\epsilon}_{2}(q)$. First assume that $C_{G}(x)$ is of type ${\rm GL}_{4}(q)\times{\rm GL}_{1}(q)$, so $\hat{x}$ has eigenvalues $(\lambda)^{4},\mu\in\mathbb{F}_{q}$. If $x^{G_{I}}\cap H\subseteq B$, then $|x^{G}\cap H|$ is bounded above by the number of ways of arranging the eigenvalues of $\hat{x}$ on the main diagonal, so we get $|x^{G_{I}}\cap H|\leqslant 5$, and the bound quickly follows, so we may now assume that $x^{G_{I}}\cap(H\setminus B)\neq\emptyset$. We may gain write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that $\pi$ embeds in ${\rm GL}(V)$ as a permutation matrix, and so we quickly deduce that if $\hat{x}=[\lambda I_{4},\mu]$, then $\pi$ must be a transposition, and consequently, $x$ must embed in $G$ as a $t_{1}$ involution. Therefore, we get
$|x^{G_{I}}\cap H|\leqslant|x^{G_{I}}\cap B|+|x^{G_{I}}\cap(H\setminus B)|=5+10(q-\epsilon)$
and so the claim is true.
If $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{3}(q)\times{\rm GL}^{\epsilon}_{2}(q)$, then using the same arguments we find that
$|x^{G_{I}}\cap B|\leqslant\binom{5}{2}\,\,\,\text{and}\,\,\,|x^{G_{I}}\cap(H\setminus B)|=\frac{1}{2}\binom{5}{2}\binom{3}{2}(q-1)^{3}$
and since $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{8}$ the result follows.
If $H$ is of type ${\rm GL}_{1}^{\epsilon}(q^{5}).5$, then it follows directly from [22, 5.3.2] that there are no elements in $H$ with $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)q^{12}$, and so we have
$|x^{G}\cap H|\leqslant|H|<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)q^{12}\right)^{11/20}<|x^{G_{0}}|^{11/20}$
for all $x\in H$, as required.
Finally, if $H$ is of type ${\rm O}_{5}(q)$, ${\rm GL}_{5}^{\epsilon^{\prime}}(q_{0})$, then it is straightforward to verify the required bound using the relevant tables in [22, Appendix B], as in the proof of Proposition 4.7. ∎
Proposition 4.10.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{4}(q)$ and $x\in G_{I}$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-9/20}$.
Proof.
If $q\leqslant 11$, then we verify that the bound holds using Magma, so we may assume for the remainder of the proof that $q\geqslant 13$. By inspecting [12, Tables 8.8-8.11] we deduce that the possible types of $H$ are the ones listed in Table 9. Moreover, Lemma 3.18 implies that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$ for any prime order element $x\in G_{I}$.
[…]
Table 9. Non-standard subgroups of almost simple groups with socle ${\rm L}_{4}^{\epsilon}(q)\,(q\geqslant 8)$.
If $H\in\mathscr{S}$, then using the Magma command AutomorphismGroupSimpleGroup(), we can construct ${\rm Aut}(H)$ and obtain bounds for $m=\max\{i_{s}({\rm Aut}(H))\,:\,s\in\pi(|{\rm Aut}(H)|)\}$. For $q>13$ we find
$|x^{G}\cap H|\leqslant m\leqslant 5184<|x^{G_{0}}|^{11/20}$ (3)
and this bound is sufficient, so we may now assume that $q=13$. If $H$ is not of type ${\rm U}_{4}(2)$, or $r\neq 5$, then again by computing directly using Magma, we find that $m\leqslant 891$, and so the claim holds.
Now let us assume that $q=13$, $r=5$ and $H$ is of type ${\rm U}_{4}(2)$. Then $\epsilon=+$ and we note that $r$ does not divide q1q-1, so Lemma 3.18 implies that that $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{8}$. The claim is then verified via (3).
Now assume $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{4}$. If $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{10}$, then the trivial bound $|x^{G}\cap H|\leqslant|H|$ is sufficient. We may now assume that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{10}$. Lemma 3.18 implies that if this holds, then either $x$ is unipotent with Jordan form $J_{2}^{s},J_{1}^{4-2s}]$ for $s\in\{1,2\}$, or $x$ is semisimple and $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{s}(q)\times{\rm GL}^{\epsilon}_{4-s}(q)$ for $s\in\{2,3\}$ or ${\rm GL}_{2}(q^{2})$. In those cases we can argue in the same way as in the proof of Proposition 4.9 and obtain the required bound, so we leave the details to the reader.
If $H$ is of type $2^{4}.{\rm Sp}_{4}(2)$, then the claim follows directly from the bounds established in the proof of [15, 6.5].
Finally, if $H$ is of type ${\rm GL}_{4}(q_{0})$, ${\rm GU}_{4}(q_{0})$, or ${\rm O}^{\epsilon}_{4}(q)$ then we can compute directly using the relevant tables in [22, Appendix B]. ∎
Proposition 4.11.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{4}(q)$ and $x\in G$ is a field, graph, or graph-field automorphism of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-3/10}$.
Proof.
If $q\leqslant 11$, then we can verify the claim using Magma, so we may assume that $q\geqslant 13$ for the remainder of the proof. First suppose that $r$ is odd. Note that we must have $q\neq p$, and so $G$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{4}$, ${\rm GL}_{4}^{\epsilon^{\prime}}(q_{0})$ or ${\rm O}^{\epsilon^{\prime\prime}}_{4}(q)$. In those cases the bounds in the proofs of the bounds in the proofs of [16, 2.7][15, 8.2] and [15, 5.1] respectively are sufficient.
Now suppose that $r=2$. If $H\in\mathscr{S}$, then again $q=p$, and so $x$ must be an involutory graph automorphism. The same argument we used in Proposition 4.10 shows that $|x^{G}\cap H|\leqslant 891$. Moreover, Lemma 3.18 implies that $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$ and those bounds are sufficient unless $q=13$. Now suppose that $q=13$. If $\epsilon=+$, then we find
$|x^{G_{0}}|=\frac{1}{2}\frac{|{\rm PGL}_{4}(13)|}{|{\rm PGSp}_{4}(13)|}=371124$
and one can check that the claim holds. On the other hand, if $\epsilon=$-, then $H$ is not of type ${\rm U}_{4}(2)$, so we get
$|x^{G}\cap H|\leqslant i_{2}({\rm P\Gamma L}_{2}(7))\leqslant 241$
and this bound is sufficient.
If $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{4}$, then the bounds in [16, 2.7] are sufficient if $x$ is a field or graph automorphism. If $x$ is a graph-field automorphism then Lemma 3.18 gives $|x^{G_{0}}|>\frac{1}{8}\left(\frac{q}{q+1}\right)q^{15/2}$ and this together with the bound $|x^{G}\cap H|<|H|$ is sufficient.
If $H$ is of type $2^{4}.{\rm Sp}_{4}(2)$, then $q=p$, and so again, $x$ must be an involutory graph automorphism. As demonstrated in the proof of [15, 6.6], we have $|x^{G}\cap H|\leqslant 340$, and one can check that this bound suffices.
Finally, if $H$ is of type ${\rm O}_{4}^{\epsilon^{\prime\prime}}(q)$ or ${\rm GL}_{4}(q_{0})$, then the claim can be verified via the bounds in the proofs of [15, 8.2] and [15, 5.1] respectively. ∎
Proposition 4.12.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{3}(q)$ and $x\in G_{I}$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-9/20}$.
Proof.
The claim can easily be verified using Magma if $q\leqslant 19$, and so we may assume for the remainder of the proof that $q\geqslant 23$. By inspecting [12, Tables 8.3-8.6] we deduce that the possible types of $H$ are the ones listed in Table 10.
[…]
Table 10. Non-standard subgroups of almost simple groups with socle ${\rm L}_{3}^{\epsilon}(q)\,(q\geqslant 16)$.
If $H\in\mathscr{S}$, then as in the proofs of Propositions 4.9, and 4.10, we use the Magma command AutomorphismGroupSimpleGroup() to construct ${\rm Aut}(H)$ and compute $m=\max\{i_{s}(H)\,:\,s\,\text{prime}\}$. If $H$ is of type ${\rm L}_{2}(7)$, then
$|x^{G_{I}}\cap H|\leqslant m=56$
and this bound is sufficient for all $x\in H$, as $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$ by Lemma 3.18 for all $x\in G_{I}$. On the other hand, if $x\in H$ is of type $A_{6}$, then $m=144$, and since $q\geqslant 23$, we deduce that
$|x^{G_{I}}\cap H|\leqslant m=144<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}\right)^{11/20}<|x^{G_{0}}|^{11/20}.$
Now suppose that $H$ is of type $3^{2}.{\rm Sp}_{2}(3)$. Since $q\geqslant 23$, the trivial bound
$|x^{G_{I}}\cap H|\leqslant|H|<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}\right)^{11/20}<|x^{G_{0}}|^{11/20}$
is sufficient.
Next let us assume that $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{3}$. If $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6}$, then we obtain the result via the trivial bound $|x^{G_{I}}\cap H|\leqslant|H\cap G_{I}|$, so we may assume that $|x^{G_{0}}|\leqslant\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6}$. In this case, Lemma 3.18 implies that if $x$ is unipotent, then it has Jordan form $J_{2},J_{1}]$ and if $x$ is semisimple, then $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{2}(q)\times{\rm GL}^{\epsilon}_{1}(q)$. If $x$ is unipotent, then we argue in exactly the same way as in the proof of Proposition 4.9 and we find $|x^{G_{I}}\cap H|\leqslant 3(q+1)$. Moreover, we note that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$, so the claim holds. If $x$ is semisimple, then the argument is also identical with the one in the proof of Proposition 4.9, and we get $|x^{G}\cap H|\leqslant 3(q+1)+3$, which is again sufficient.
If $H$ is of type ${\rm GL}_{1}^{\epsilon}(q^{3})$, then it follows immediately from [22, 5.3.2] that $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6}$ for every $x\in H$, and so the trivial bound $|x^{G_{I}}\cap H|\leqslant|H\cap G_{I}|$ is sufficient.
Finally, if $H$ is of type ${\rm GL}^{\epsilon^{\prime}}_{3}(q_{0})$ or ${\rm O}_{3}(q)$, then we obtain the required result by computing exact fixed point ratios using the relevant tables in [22, Appendix B]. We will only consider the case where $\epsilon=+$ and $H$ is of type ${\rm GU}_{3}(q^{1/2})$ here to avoid unnecessary repetition, as all cases are very similar.
First suppose that $x$ is unipotent. Then $x$ either has Jordan form $J_{3}]$ or $J_{2},J_{1}]$ on $V$. If $x$ has Jordan form $J_{3}]$, then we find
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm GU}_{3}(q_{0})|}{q_{0}^{2}|{\rm GU}_{1}(q_{0})|}=q_{0}(q_{0}^{2}-1)(q_{0}^{3}+1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{3}(q)|}{q^{2}|{\rm GL}_{1}(q)|}=q(q^{2}-1)(q^{3}-1)$
and one can verify that the claim holds. Similarly, if $x$ has Jordan form $J_{2},J_{1}]$, then we get
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm GU}_{3}(q_{0})|}{q_{0}^{3}|{\rm GU}_{1}(q_{0})|^{2}}=(q_{0}-1)(q_{0}^{3}+1)$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{3}(q)|}{q^{3}|{\rm GL}_{1}(q)|^{2}}=(q+1)(q^{3}-1)$
and the required bound holds in this case as well.
Now suppose that $x$ is semisimple, and let $i$ be minimal such that $r|q^{i}-1$ and $i_{0}$ be minimal such that $r|q_{0}^{2i_{0}}-1$. Note that $i=i_{0}/2$ if $i_{0}$ is even and $i=i_{0}$ otherwise. There are two possible centraliser types for $x$, namely ${\rm GL}_{2}(q)\times{\rm GL}_{1}(q)$ and ${\rm GL}_{1}(q)^{3}$. If $C_{G}(x)$ is of type ${\rm GL}_{2}(q)\times{\rm GL}_{1}(q)$, then we compute
$|x^{G_{I}}\cap H|\leqslant q_{0}^{2}(q_{0}^{2}-q_{0}+1)\,\,\,\text{and}\,\,\,|x^{G_{0}}|\geqslant q^{2}(q^{2}+q+1)$
and if $C_{G}(x)$ is of type ${\rm GL}_{1}(q)^{3}$, then
$|x^{G_{I}}\cap H|\leqslant q_{0}^{3}(q_{0}-1)(q_{0}^{2}-q_{0}+1)+q_{0}^{3}(q_{0}^{3}+1)\,\,\,\text{and}\,\,\,|x^{G_{0}}|\geqslant q^{3}(q+1)(q^{2}+q+1)$
and one can check that those bounds are sufficient.
Proposition 4.13.
Suppose that $G_{0}={\rm PSL}^{\epsilon}_{3}(q)$ and $x\in G$ is a field, graph, or graph-field automorphism of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-3/10}$.
Proof.
The cases where $q\leqslant 19$ can be checked using Magma, so we may assume that $q\geqslant 23$. Lemma 3.18 implies that $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{4}$, and so unless $H$ is of type ${\rm GL}_{3}^{\epsilon^{\prime}}(q_{0})$, ${\rm O}_{3}(q)$ or ${\rm GL}_{1}^{\epsilon}(q)\wr S_{3}$, then the bound $|x^{G}\cap H|<|H|$ is sufficient.
If $H$ is of type ${\rm GL}_{3}^{\epsilon^{\prime}}(q_{0})$ or ${\rm O}_{3}(q)$ and $x$ is not an involutory graph-field automorphism then the bounds in the proofs of [15, 5.1] and [15, 8.2] are sufficient, so now assume that $x$ is an involutory graph-field automorphism. In both cases, Lemma 3.8 implies that $x^{G}\cap H\subseteq\widetilde{H}x$, where $\widetilde{H}=H\cap G_{I}$ and using [51, 1.3] we find
$|x^{G}\cap H|\leqslant i_{2}(\widetilde{H})+1<2(1+1q^{-1/2})q^{5/2}$
which is sufficient.
Finally assume that $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{3}$. If $r>2$, then the bound $|x^{G}\cap H|<|H|$ is sufficient, so now assume that $r=2$. In this case, again appealing to Lemma 3.8 gives
$|x^{G}\cap H|\subseteq\widetilde{H}x\leqslant i_{2}(\widetilde{H})+1\leqslant 16,$
and recall that $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{4}$, so we are done. ∎
Proposition 4.14.
Suppose that $G_{0}={\rm PSp}_{4}(q)$ for $q\geqslant 4$, and $x\in G_{I}$ is an element of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-9/20}$.
Proof.
Using [12, Tables 8.12-8.14], as well as Table 3, we can determine the possibilities of $H$. For most of those cases we use entirely similar methods as the ones we use to prove Propositions 4.9, 4.10, and 4.12, so we only discuss the cases where we use different methods here, to avoid unnecessary repetition. In particular, we discuss the cases where $H$ is of type $^{2}B_{2}(q)$, $O_{2}^{\epsilon}(q)\wr S_{2}$, or $O^{-}_{2}(q^{2}).2$ for $q$ even, ${\rm L}_{2}(q)$ for $q$ odd.
If $H$ is of type $^{2}B_{2}(q)$, then we obtain the required bounds by inspecting the proof of [18, 4.4], so it remains to discuss the other four cases.
Suppose that $H$ is of type $O_{2}^{\epsilon}(q)\wr S_{2}$. Note that if $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient. Now suppose that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$. Lemma 3.18 implies that this only occurs when $x$ is a $b_{1}$ involution. In this case we know that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$, where $H_{0}=H\cap G_{0}$. We can infer from [12, Table 8.14] that $H_{0}<K<G_{0}$, where $K={\rm Sp}_{2}(q)\wr S_{2}$, and so $|x^{G}\cap H|=|x^{G}\cap H_{0}|\leqslant|x^{G}\cap K|$. Now [16, 2.9] gives $|x^{G}\cap K|\leqslant 2q^{2}$ and we get $|x^{G_{0}}|=q^{4}-1$, and one can check that those bounds are sufficient for $q>32$.
If $q\leqslant 32$, then we can construct $H_{0}$ using Magma as the normaliser of an appropriate Sylow subgroup. We only give the details for the case where $H$ is of type $\Omega^{-}_{2}(q)\wr S_{2}$ when $q=32$, as the other cases are similar. In this case we can obtain $H_{0}$ as the normaliser of a Sylow $3$-subgroup. Note that $3$ divides $q+1$, and in particular, the $3$-factor of $|G_{0}|$ is the same as the $3$-factor of $(q+1)^{2}$, so a Sylow $3$-subgroup $P$ of $G_{0}$ is contained in some maximal torus $T\cong C_{q+1}\times C_{q+1}$. In particular, $P\cong C_{3}\times C_{3}$ is a characteristic subgroup of $T$, so $N_{G}(T)\leqslant N_{G}(P)$. One then checks using Magma that $|N_{G}(P)|=|H_{0}|$, so $N_{G}(P)=H_{0}$, and we can then check that the bound $|x^{G}\cap H|<i_{2}(H_{0})$ is sufficient.
The case where $H$ is of type $O^{-}_{2}(q^{2}).2$ is very similar. If $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient, so it remains to prove the claim for $b_{1}$-type involutions. Again we note that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$, and this time $H_{0}=H\cap G_{0}$ embeds inside a maximal subgroup $K$ of type ${\rm Sp}_{2}(q^{2})$, so $|x^{G}\cap H|\leqslant|x^{G}\cap K|$. However, $K$ does not contain any $b_{1}$-type involutions, and therefore $|x^{G}\cap H|=|x^{G}\cap K|=0$, and the claim trivially holds.
Finally, suppose that $H\in\mathscr{S}$ is of type ${\rm L}_{2}(q)$. Let $W=\langle v_{1},v_{2}\rangle$ be the natural module for $L={\rm SL}_{2}(q)$. Then the embedding of $L$ in ${\rm Sp}_{4}(q)$ is the image of the representation $\rho:L\rightarrow{\rm Sp}_{4}(q)$ corresponding to the $L$-module $S^{3}(W)$. Let $\hat{x}$ be the lift of $x$ in ${\rm GSp}_{4}(q)$ and first suppose that $x$ is unipotent. In this case $x$ has associated Jordan decomposition $J_{2}]$ one $W$ and we note that
$\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)=\begin{pmatrix}1&1&1&1\\ 0&1&2&3\\ 0&0&1&3\\ 0&0&0&1\end{pmatrix}=M$
with respect to the basis $\{v_{1}^{3},v_{1}^{2}v_{2},v_{1}v_{2}^{2},v_{2}^{3}\}$. Now, since $M-I_{4}$ has rank $3$, we deduce that $M$ is not a transvection, and therefore $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$. Moreover, we know that $|x^{G}\cap H|\leqslant q^{2}-1$, and so the claim holds.
Next suppose that $x\in H$ is semisimple and $r$ is odd, and let $i$ be minimal such that $r|q^{i}-1$. First suppose that $i=1$. Here $\hat{x}$ is $L$-conjugate to $\lambda,\lambda^{-1}]$ for some $\lambda\in\mathbb{F}_{q}^{\times}$ and we have
$\rho([\lambda,\lambda^{-1}])=[\lambda^{3},\lambda,\lambda^{-1},\lambda^{-3}]$
If $x$ has order $3$, then $C_{G}(x)$ is of type ${\rm GL}_{1}(q)\times{\rm Sp}_{2}(q)$. In this case we have
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{4}(q)|}{|{\rm GL}_{1}(q)||{\rm Sp}_{2}(q)|}=q^{3}(q+1)(q^{2}+1)$
and
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm GL}_{2}(q)|}{|{\rm GL}_{1}(q)|^{2}}=q(q+1)$
and so the bound holds.
If $r\geqslant 5$, then $x$ is a regular semisimple element, and we have
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{4}(q)|}{|{\rm GL}_{1}(q)|^{2}}=q^{4}(q+1)^{2}(q^{2}+1).$
The case where $i=2$ is similar, so we leave the details to the reader.
We finally consider involutions. Let $K$ denote the algebraic closure of $\mathbb{F}_{q}$ and let $\bar{G}={\rm PSp}_{4}(K)$ and $\bar{H}={\rm PSL}_{2}(K)$. Now let $\bar{W}$ be the natural module for ${\rm SL}_{2}(K)$ and $\bar{\rho}:{\rm SL}_{2}(K)\rightarrow{\rm Sp}_{4}(K)$ denote the representation afforded by the $K{\rm SL}_{2}(q)$-module $S^{3}(\bar{W})$. Now note that $\bar{H}$ has a unique class of involutions, $y^{\bar{H}}$, and $y$ lifts in ${\rm SL}_{2}(K)$ to a diagonal matrix $\hat{y}=[\lambda,\lambda^{-1}]$ for some $\lambda\in K^{\times}$ of order $4$. Then
$\bar{\rho}([\lambda,\lambda^{-1}])=[\lambda^{-1},\lambda,\lambda^{-1},\lambda]$
with respect to the basis $\{w_{1}^{3},w_{1}^{2}w_{2},w_{1}w_{2}^{2},w_{2}^{3}\}$. Now there are two classes of involutions in $\bar{G}$, one of type $A_{1}A_{1}$ and one of type $A_{1}T_{1}$, and the former lift to involutions in ${\rm Sp}_{4}(K)$, whilst the latter lift to elements of order $4$. We therefore deduce that $y$ embeds in $\bar{G}$ as an $A_{1}T_{1}$ involution.
Since both $\bar{H}$ and $\bar{G}$ are stable under the standard Frobenius morphism $\sigma$, we deduce that if $x\in\bar{H}_{\sigma}$ is an involution, then it embeds in $\bar{G}_{\sigma}={\rm Inndiag}(G_{0})$ as an involution with centraliser type ${\rm GL}_{2}^{\epsilon}(q)$. We therefore have
$|x^{\bar{G}_{\sigma}}\cap H|\leqslant i_{2}({\rm PSL}_{2}(q))=q^{2}-1$
and
$|x^{G_{0}}|=|x^{\bar{G}_{\sigma}}|\geqslant\frac{|{\rm Sp}_{4}(q)|}{2|{\rm GU}_{2}(q)|}=\frac{1}{2}q^{3}(q-1)(q^{2}+1)$
so the claim holds and the proof is complete. ∎
Proposition 4.15.
Suppose that $G_{0}={\rm PSp}_{4}(q)$ for $q\geqslant 4$, and $x\in G$ is a field or graph-field automorphism of prime order $r$. Then $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-3/10}$.
Proof.
The methods we use here are similar to the ones in the proofs of Propositions 4.11 and 4.13, so we leave the details for the reader. ∎
Proof of Theorem 4.1.
This follows by combining Propositions 4.3, 4.4, 4.5, 4.6, 4.7, 4.8, 4.9, 4.10, 4.11, 4.12, 4.13, 4.14, and 4.15. ∎
5. Eta function bounds
We will use Proposition 3.14 and variations as heavily as possible in the proof of Theorem 3.1. In most cases, the bounds in Proposition 3.15 suffice, but in some cases we will need stronger bounds on $T_{G}$, where $T_{G}$ is as defined in Definition 3.10, or even stronger bounds on $\eta_{G}(t)$ for some $t\geqslant T_{G}$. We establish those bounds in this section.
Proposition 5.1.
If $G_{0}={\rm PSp}_{6}(q)$ and $q$ is even, then $\eta_{G}(1/5)<0.9$.
Proof.
If $q\leqslant 16$, then we check the claim using Magma. In particular, we first call ClassicalClasses("Sp", 6, q), and we compute $\alpha=\sum_{i=1}^{k}|x_{i}^{G_{0}}|^{-1/5}$ directly, where $x_{1},\ldots,x_{k}$ denote representatives of the prime order classes of $G_{0}$. It now suffices to manually add the contribution $\beta$ from outer automorphisms to $\alpha$ to obtain an upper bound for $\eta_{G}(1/5)$. We note that
$\beta=\sum_{r\in\pi(f)}(r-1)\cdot\left(\frac{|{\rm PSp}_{6}(q)|}{|{\rm PSp}_{6}(q^{1/r})|}\right)^{-1/5}$
where recall that $f=\log_{p}(q)$ and we confirm that $\alpha+\beta<0.9$ for $q\leqslant 16$.
If $q\geqslant 32$, then we use the bound obtained by Burness in the proof of [18, 2.3]. In particular, we have
$\eta_{G}(t)$ $<2\left(\frac{1}{2}(q^{6}-1)\right)^{-t}+\left(2+\frac{q}{2}\log(q^{2}-1)\right)\cdot\left(\frac{1}{4}(q+1)^{-1}q^{11}\right)^{-t}$ (4)
$+((q^{2}+1)(q^{6}-1))^{-t}+\left(6+\frac{q}{2}\log(q^{2}-1)\right)\cdot\left(\frac{1}{4}(q+1)^{-1}q^{13}\right)^{-t}$
$+q^{2}\log(q^{2}+1)\left(\frac{1}{8}q^{14}\right)^{-t}+\frac{1}{3}q^{3}\log(q^{6}-1)\cdot\left(\frac{1}{4}q^{18}\right)^{-t}$
$+2(\log_{2}q-1)\cdot\log(\log_{2}q+2)\cdot\left(\frac{1}{4}q^{21/2}\right)^{-t}$
and one can check that this bound is sufficient if $q\geqslant 32$.
Finally, for $16<q<32$ we use a more precise version of (4). In particular, $\log(q^{2}-1),\log(q^{2}+1)$ and $\log(q^{6}-1)$ appear in (4) as approximations of the number of prime divisors of $q^{2}-1,q^{2}+1$, and $q^{6}-1$ respectively, and replacing those approximations with the exact number of prime divisors of those numbers for $16<q<32$ gives the desired bound. ∎
Proposition 5.2.
If $G_{0}={\rm\Omega}_{7}(q)$ and $q$ is odd, then $\eta_{G}(1/5)<0.9$.
Proof.
If $q\leqslant 13$, then again we compute directly using Magma exactly as in the proof of Proposition 5.1, so we may assume that $q\geqslant 17$ for the remainder of the proof.
Our aim here is to derive a similar bound to the one used in Proposition 5.1. To do this, we must bound the number of $G_{0}$-conjugacy classes of prime order elements of each type in $G$, as well as the size of each $G_{0}$-conjugacy class, and then group them accordingly. We record the relevant bounds for each type of prime order element in Table 11, but for the sake of brevity, we only discuss a few cases in detail. The cases not discussed are very similar.
Table 11 is organised as follows: The first row records the type of each prime order element in $G_{0}$. For unipotent elements we record the Jordan form on $V$, for odd order semisimple elements we record the centraliser type, and if $x$ is a semisimple involution or a field automorphism, then we use the standard notation from [22, Sections 3.5.2 and 3.5.5]. The second row records an upper bound $f(x,q)$ on the number of $G_{0}$-classes of each type and the third row records a lower bound $g(x,q)$ on the $G_{0}$-conjugacy class size of each element.
We first consider unipotent classes. Note that the possible Jordan forms on $V$ are as follows:
$J_{7}],[J_{5},J_{1}^{2}],[J_{3},J_{2}^{2}],[J_{3},J_{1}^{4}],[J_{2}^{2},J_{1}^{3}]$
We can estimate how many classes correspond to each Jordan decomposition by counting the number of labelled partitions corresponding to that form and consulting [22, 3.5.14] to determine which $G_{I}$-classes split in $G_{0}$. The $G_{0}$-class bounds in Table 11 are obtained via [15, 3.21].
We now look at semisimple classes of odd order. Let $i$ be minimal such that $r|q^{i}-1$. Note that since $r$ divides $|G|$ we must have $i\in\{1,2,3,4,6\}$. The $G_{0}$-class bounds are determined via [15, Tables 3.6 and 3.8], so it remains to determine $f(x,q)$. We must treat each possibility in turn. We indicatively discuss two cases in detail here.
Suppose first that $i=6$. In this case $C_{G}(x)$ is of type ${\rm O}_{1}(q)\times{\rm GU}_{1}(q^{3})$, and $x$ is of the form $\Lambda,1]$, for some $\sigma$-orbit $\Lambda$. We now know that if $i=6$, then $r|q^{3}+1$, and moreover that there are exactly $\frac{r-1}{6}\leqslant\frac{q^{3}}{6}$ $\sigma$-orbits, so the number of $G_{0}$-classes of elements of this type in $G$ is at most $\frac{q^{3}}{6}\pi(q^{3}+1)\leqslant\frac{q^{3}}{6}\log(q^{3}+1)$.
Now suppose that $i=1$ and $C_{G}(x)$ is of type ${\rm GL}_{1}(q)^{3}\times{\rm O}_{1}(q)$. In this case r|q1r|q-1, and since there are exactly $\frac{q-1}{2}$ distinct pairs $(\lambda,\lambda^{-1})\in(\mathbb{F}_{q}^{\times})^{2}$, we conclude that there are at most $\binom{(q-1)/2}{3}\pi(q-1)\leqslant\frac{q^{3}}{48}\log(q-1)$ $G_{0}$-classes of this type.
We next consider semisimple involutions. By inspecting [22, Table B.8] we find that there are at most six $G_{0}$-classes of semisimple involutions, namely $t_{i}$ for $i\in\{1,2,3\}$ and $t^{\prime}_{i}$ for $i\in\{1,2,3\}$.
Finally, we can infer from [22, 3.5.20] that for a fixed prime divisor $r$ of $\log_{p}q$, there are at most $2(r-1)\leqslant 2\log_{p}q$ $G_{0}$-classes of field automorphisms of order $r$, and $\log_{p}q$ has at most $\log(\log_{p}q+2)$ distinct prime divisors, so there are at most $2\log_{p}q\cdot\log(\log_{p}q+2)$ $G_{0}$-classes of field automorphisms.
Now if $x_{1},\ldots,x_{t}$ is a set of representatives for the $G_{0}$-classes of prime order elements in $G$, then we have
$\eta_{G}(1/5)\leqslant\sum_{i=1}^{t}f(x_{i},q)\cdot g(x_{i},q)^{-1/5}$
where $f(x_{i},q)$ and $g(x_{i},q)$ are recorded in Table 11 for every $i$ and one can check that this bound is sufficient for $q\geqslant 17$.
[…]
Table 11. Bounds on conjugacy classes and number of prime order elements in $\Omega_{7}(q)$.
Proposition 5.3.
If $G_{0}={\rm PSL}^{\epsilon}_{6}(q)$ and $q\geqslant 3$, then $\eta_{G}(1/3)<0.979$.
Proof.
We adopt essentially the same strategy as the one used in the proof of [18, 2.3]. For an element $x\in{\rm PGL}^{\epsilon}_{6}(q)$, we write $\nu(x)$ for the codimension of the largest eigenspace of $x$ on $V$, and for $1\leqslant s\leqslant 5$ we denote the number of distinct $G_{0}$-classes of unipotent (respectively semisimple) prime order elements $x\in G$ with $\nu(x)=s$ by $k_{s,u}$ (respectively $k_{s,s}$). Moreover, we will write $c_{u}(s)$ (respectively $c_{s}(s)$) for the size of the smallest unipotent (respectively semisimple) $G_{0}$-class $y^{G}$ with $\nu(y)=s$, $\pi$ for the distinct prime divisors of $\log_{p}q$ and for $r\in\pi$ we will write $\mathcal{F}_{r}$ for the set of $G$-classes of field automorphisms of order $r$. Finally, we will denote the set of $G$-classes of involutory graph-field automorphisms by $\mathcal{F}_{gf}$, and we will write $C_{1},C_{2},C_{3}$ for the $G_{I}$-classes of a graph automorphism of type ${\rm PGSp}_{6}(q),{\rm PGO}^{+}_{6}(q)$, and ${\rm PGO}^{-}_{6}(q)$ respectively. Noting that [22, 3.2.14] tells us that $C_{i}$ splits into at most 33 $G_{0}$-classes for all $i\in\{1,2,3\}$, we have
$\eta_{G}(t)\leqslant\sum_{C\in{\mathcal{F}_{gf}}}|C|^{-t}+\sum_{r\in\pi}\sum_{C\in\mathcal{F}_{r}}|C|^{-t}+\sum_{i=1}^{3}3\left(\frac{|C_{i}|}{3}\right)^{-t}+\sum_{s=1}^{5}(k_{s,u}+k_{s,s})\cdot c(s)^{-t}$
and therefore it suffices to find appropriate bounds for $|\mathcal{F}_{r}|$, $|\mathcal{F}_{gf}|$, $|C|$, $|C_{i}|$, $|\pi|$, $k_{s,u}$, $k_{s,s}$, $c_{u}(s)$, and $c_{s}(s)$ for $s\in\{1,\ldots,5\}$.
Firstly, we know that every $N\in\mathbb{N}$ has at most $\log(N+2)$ distinct prime divisors, and so $|\pi|\leqslant\log(\log_{p}q+2)$ if $q\neq p$ and $|\pi|=0$ if $p=q$. Moreover, [44, 4.9.1] gives $|\mathcal{F}_{r}|\leqslant(6,q-\epsilon)(r-1)\leqslant 6(\log_{p}q+1)$. Similarly, it follows straight from [22, 3.2.15] that the $G_{I}$-class of graph-field automorphisms splits into $(6,q^{1/2}-1)$ $G_{0}$-classes, and so $|\mathcal{F}_{gf}|\leqslant 6$.
We now consider unipotent elements. Note that $(8)$ in [15] tells us that there are at most $2^{s}$ unipotent ${\rm Inndiag}(G_{0})$-classes with $\nu(x)=s$ for $s\in\{1,\ldots,,5\}$ and each class can split to at most $6$ classes in $G_{0}$. Therefore, $k_{s,u}<6\cdot 2^{s}$. Moreover, we can infer from [15, 3.22] and [22, 3.2.7] that
[…]
Similarly, for semisimple elements [15, 3.40] tells us that there are at most $q^{s+1}$ prime order elements of nu-value $s$ of order $r$, so now it suffices to bound the possibilities for $r$ to bound $k_{s,s}$. Note that $r$ must divide $\prod_{i=1}^{3}(q^{i}-\epsilon)(q^{2i}-1)<q^{22}$, and since every integer $N>6$ has at most $\log N$ divisors, we conclude that $k_{s,s}<22\log q\cdot q^{s+1}$.
Finally, [15, 3.36-3.37] gives
[…]
and one can check that those bounds are sufficient for $q\geqslant 11$. If $3\leqslant q\leqslant 9$, then we can compute $k_{s,s}$, and $k_{s,u}$ precisely and obtain the result this way. For example, let us consider the case $(\epsilon,q,s)=(+,8,4)$. If $\nu(x)=4$ for a unipotent element $x\in G$, then the Jordan decomposition of $x$ on $V$ must comprise of two blocks. But one cannot have an involution in $G$ whose Jordan decomposition only has two blocks, so $k_{4,u}=0$.
Now suppose that $x$ is semisimple and let $i$ be minimal such that $r|q^{i}-1$, so $i\in\{1,\ldots,6\}$. Now it is easy to see that if $i\geqslant 4$, then the lift $\hat{x}$ of $x$ in ${\rm GL}(V)$ is of the form $\Lambda,I_{6-i}]$ for some $\sigma$-orbit $\Lambda$, so $\nu(x)=5$, a contradiction, and so $i\geqslant 3$. If $i=3$, then since $\nu(x)=4$ we must have $\hat{x}=[\Lambda^{2}]$ for some $\sigma$-orbit $\lambda$. Moreover, by inspecting the prime divisors of $|G_{I}|$, we find that $r=73$, and so, as explained in [22, Section 3.2.1], there are exactly $\frac{73-1}{3}=24$ distinct $\sigma$-orbits for $i=3$, so we get precisely $24$ $G_{0}$-classes with $\nu(x)=4$.
Now assume that $i=2$. Note that for all the eigenvalues of $\hat{x}$ to occur with multiplicity at least $2$, $\hat{x}$ must be of the form $\Lambda^{2},I_{2}]$. Moreover, the assumption $i=2$ implies that $r=3$, and there is a unique $\sigma$-orbit when $r=3$.
[…]
6. Proof of Theorem 3.1
Let $\mathcal{C}$ denote the set of simple classical groups. We will divide the possibilities for $G_{0}\in\mathcal{C}$ into the following six collections, according to the method we use to prove Theorem 3.1, and we will handle each one in turn:
$\mathcal{A}_{1}$ $=\{{\rm PSp}_{8}(q),{\rm PSL}_{8}^{\epsilon}(q),{\rm P\Omega}^{\epsilon}_{10}(q)\}$
$\mathcal{A}_{2}$ $=\{{\rm PSL}^{\epsilon}_{6}(q)\}$
$\mathcal{A}_{3}$ $=\{{\rm PSp}_{6}(q)\,:\,q\text{ even}\}\cup\{{\rm\Omega}_{7}(q)\,:\,q\text{ odd}\}$
$\mathcal{A}_{4}$ $=\{{\rm PSL}_{3}^{\epsilon}(q),{\rm PSL}_{4}^{\epsilon}(q),{\rm PSL}_{5}^{\epsilon}(q),{\rm PSp}_{4}(q)\}$
$\mathcal{A}_{5}$ $=\{{\rm PSL}_{2}(q)\}$
$\mathcal{A}_{6}$ $=\mathscr{C}\setminus(\mathcal{A}_{1}\cup\mathcal{A}_{2}\cup\mathcal{A}_{3}\cup\mathcal{A}_{4}\cup\mathcal{A}_{5})$
Proposition 6.1.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{6}$.
Proof.
In this case, it follows immediately from Theorem 3.9 and Propositions 3.14 and 3.15 that any non-standard $4$-tuple $\tau$ of $G$ is regular, and moreover Proposition 3.13 implies that $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$. ∎
Proposition 6.2.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{1}$.
Proof.
If $G_{0}={\rm PSp}_{8}(q)$ and $q=2$, then the claim can be verified using Magma, using the function RegTuplesPlus(), so we may assume for the remainder of the proof that $q\geqslant 3$ if $G_{0}={\rm PSp}_{8}(q)$.
Propositions 4.3,  4.4, and  4.5, together with Theorem 3.9, tell us that ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-1/3}$ for every core-free subgroup $H$ of $G$ and all $x\in G$ of prime order. Moreover, Proposition 3.15 gives $T_{G}<1/3$, and so if $\tau=(H_{1},\ldots,H_{4})$ is any non-standard tuple of $G$, then Proposition 3.14 implies that $\tau$ is regular. Finally, Proposition 3.13 implies that $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$. ∎
Proposition 6.3.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{2}$.
Proof.
If $q=2$, then as described in Section 2.2, we can check that all non-standard $4$-tuples (or $5$-tuples if $G=U_{6}(2).2$) are regular using the Magma function RegTuplesPlus(). If $G=U_{6}(2).2$, then we can also verify that the only non-regular $4$ tuple is $(H,H,H,H)$ for $H={\rm U}_{4}(3).2^{2}$ as the same function outputs all the non-regular $4$-tuples. We may therefore assume that $q\geqslant 3$ for the remainder. Now note that Proposition 4.6, together with Theorem 3.9 implies that
$\widehat{Q}(G,\tau)\leqslant\eta_{G}(1/3)+\alpha,$
where
$\alpha=\frac{q-1}{2}\log(q-1)\frac{(q^{6}(q+1)(q^{2}+1)(q^{3}+1))^{4}}{(q^{9}(q-\epsilon)(q^{3}-\epsilon)(q^{5}-\epsilon))^{3}}$
for any non-standard $4$-tuple $\tau$ of $G$. In view of Proposition 5.3, it suffices to show that $\alpha<0.021$ for all $q\geqslant 3$ to prove that $\tau$ is regular, and one can easily check that this is true. Moreover, we see that $\alpha\rightarrow 0$ as $q\rightarrow\infty$, and this combined with Proposition 3.13 implies that $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$. ∎
Proposition 6.4.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{3}$.
Proof.
We can check that all non-standard tuples of $G$ are regular if $q\leqslant 5$ using Magma, so we may assume that $q\geqslant 7$ for the remainder of the proof. Let $\tau=(H_{1},\ldots,H_{4})$ be a non-standard tuple of $G$. Note that if $H_{i}$ is of type $G_{2}(q)$ for all $i\in\{1,\ldots,4\}$, then it follows straight from [18, Theorem 1] that $\tau$ is regular, so we may assume that at least one of the $H_{i}$’s is not of type $G_{2}(q)$, say $H_{1}$ without loss of generality. Moreover, [17, 2.13] implies that if $H$ is of type $G_{2}(q)$ and $x\in H$ has prime order, then unless $x$ is unipotent with Jordan form $J_{2}^{2},J_{1}^{3}]$ or $J_{3}^{2},J_{1}]$, we have ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-0.268}$.
Now let $x_{1},\ldots,x_{t}$ be a set of representatives of the prime order conjugacy classes of $G$, and without loss of generality assume that $x_{t-1}$ and $x_{t}$ are the representatives of the unipotent $G$-classes with Jordan form $J_{2}^{2},J_{1}^{3}]$ and $J_{3}^{2},J_{1}]$ respectively. Propositions 4.7 and 4.8 tell us that ${\rm fpr}(x_{j},G/H_{1})<|x_{j}^{G}|^{-0.4}$ for all $j\in[t]$, whilst ${\rm fpr}(x_{j},G/H_{k})<|x_{j}|^{-0.268}$ for $k\neq 1$ and $j\in[t-2]$. Finally, for $j\in\{t-1,t-2\}$ ${\rm fpr}(x_{j},G/H_{k})$ can be recovered from [17, Tables 2.4 and 2.5]. Combining all of this we get
$\widehat{Q}(G,\tau)\leqslant\sum_{i=1}^{t}|x_{i}^{G_{0}}|^{-3\cdot 0.268-0.4+1}+\alpha+\beta$
where
$\alpha=|x_{t-1}^{G}|^{0.6}\prod_{i=2}^{4}\frac{|{x_{t-1}}^{G}\cap H_{i}|}{|{x_{t-1}}^{G}|}=\frac{(q^{6}-1)^{3}}{((q^{2}-1)(q^{6}-1))^{12/5}}$
and
$\beta=|x_{t}^{G}|^{0.6}\prod_{i=2}^{4}\frac{|{x_{t}}^{G}\cap H_{i}|}{|{x_{t}}^{G}|}=\frac{2^{12/5}(q^{2}(q^{2}-1)(q^{6}-1))^{3}}{(q^{6}(q-1)(q^{4}-1)(q^{6}-1))^{12/5}}.$
Therefore,
$\widehat{Q}(G,\tau)\leqslant\eta_{G}(1/5)+\alpha+\beta$
and in view of Propositions 5.1 and 5.2, it suffices to show that
$\alpha+\beta<0.1,$
to show that $\tau$ is regular. We can easily check that this is true.
Moreover, noting that $\alpha,\beta\rightarrow 0$ as $q\rightarrow\infty$ and combining this with Proposition 3.13, we find that $\mathbb{P}(G,\tau)\rightarrow 1$ as $|G|\rightarrow\infty$. Finally, if $\sigma$ is a $6$-tuple of $G$ all of whose entries are of type $G_{2}(q)$, then [25, Theorem 3] implies that $\mathbb{P}(G,\sigma)\rightarrow 1$ as $|G|\rightarrow\infty$, and so $\mathbb{P}(G,6)\rightarrow 1$ as $|G|\rightarrow\infty$, as required.
Proposition 6.5.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{4}$.
Proof.
Let $\tau=(H_{1},\ldots,H_{4})$ be a non-standard tuple. If $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$ or ${\rm PSL}^{\epsilon}_{4}(q)$, then in view of Propositions 4.9, 4.10, and 4.11, we may write
$\widehat{Q}(G,\tau)\leqslant\alpha_{1}\cdot\beta_{1}^{-4/5}+\alpha_{2}\cdot\beta_{2}^{-1/5}+\alpha_{3}\cdot\beta_{3}^{-1/5}$ (5)
where $\alpha_{1}$ denotes the number of prime order $G_{0}$-classes of unipotent and semisimple elements, $\alpha_{2}$ denotes the number of $G_{0}$-classes of field and graph-field automorphisms and $\alpha_{3}$ denotes the number of $G_{0}$-classes of graph automorphisms respectively, and
$\beta_{1}$ $=\min\{|x^{G_{0}}|\,:\,x\in{\rm Inndiag}(G_{0}),|x|\in\pi(|G|)\}$
$\beta_{2}$ $=\min\{|x^{G_{0}}|\,:\,x\in\mathcal{F}\}$
$\beta_{3}$ $=\min\{|x^{G_{0}}|\,:\,x\in\mathcal{G}\}$
where $\mathcal{F}$ denotes the set of prime order field and graph-field automorphisms in $G$ and $\mathcal{G}$ the set of involutory graph automorphisms in $G$. We will derive bounds for $\alpha_{i}$ and $\beta_{i}$ for $i\in[3]$ for each possibility for $G_{0}$ separately.
First suppose that $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$. If $q\leqslant 4$, then the regularity of $\tau$ can be verified using Magma, so we may assume that $q\geqslant 5$. In this case, we can infer from [41, Table 3] that the total number of conjugacy classes in ${\rm PGL}_{5}^{\epsilon}(q)$ is bounded above by $q^{4}+8q^{3}$. We now note that the only $G$-class of unipotent elements that potentially splits in $G_{0}$ is the one corresponding to the Jordan form $J_{5}]$, and [22, 3.2.7, 3.3.10] imply that it splits into at most 55 $G_{0}$-classes. Therefore, we may take $\alpha_{1}\leqslant q^{4}+8q^{3}+4$. Moreover, [44, 4.9.1] implies that there are at most
$\left(5,\frac{q-\epsilon}{q^{1/r}-\epsilon}\right)(r-1)|\pi(\log_{p}(q))|\leqslant 5\log_{p}q\cdot\log(\log_{p}q+2)$
$G_{0}$-classes of field automorphisms and similarly at most $5$ distinct $G_{0}$-classes of involutory graph-field automorphisms by [22, 3.2.15]. We also get at most 55 $G_{0}$-classes of involutory graph automorphisms by [22, 3.2.14, 3.3.17], so we have
$\alpha_{2}\leqslant 5+5\log_{p}q\cdot\log(\log_{p}q+2)\,\,\,\text{and}\,\,\,\alpha_{3}\leqslant 5.$
Finally, it follows straight from Lemma 3.18 that $\beta_{1}>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{8}$, $\beta_{2}>\frac{1}{10}\left(\frac{q}{q+1}\right)q^{12}$, and $\beta_{3}>\frac{1}{10}\left(\frac{q}{q+1}\right)q^{14}$, and by substituting the above bounds in (5) we find that $\widehat{Q}(G,\tau)<1$ for all $q\geqslant 5$, as required.
Next assume that $G_{0}={\rm PSL}^{\epsilon}_{4}(q)$. If $q\leqslant 9$, then the regularity of $\tau$ can again be verified using Magma, so we may assume that $q\geqslant 11$ for the remainder. Again, from [41, Table 3] we see that there are at most $\alpha_{1}\leqslant q^{3}+8q^{2}$ $G_{I}$-classes of unipotent and semisimple elements, and the only classes that possibly split in $G_{0}$ are the class of unipotent elements corresponding to the Jordan decompositions $J_{4}]$ and $J_{2}^{2}]$, so $\alpha_{1}\leqslant q^{3}+8q^{2}+4$. Moreover, [44, 4.9.1] again tells us that the contribution to $\alpha_{2}$ from field automorphisms is at most $4\log_{p}q\cdot\log(\log_{p}q+2)$, and [22, 3.2.15] tells us that there are at most 44 $G_{0}$-classes of involutory graph-field automorphisms. Finally, we can infer from [22, 3.2.14, 3.3.17] that each $G_{I}$-class of graph automorphisms splits into at most 22 $G_{0}$-classes, and so
$\alpha_{2}\leqslant 4\log_{p}q\cdot\log(\log_{p}q+2)+4\,\,\,\text{and}\,\,\,\alpha_{3}\leqslant 6$
Now Lemma 3.18 gives $\beta_{1}>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, $\beta_{2}>\frac{1}{8}\left(\frac{q}{q+1}\right)q^{15/2}$, and $\beta_{3}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$. Substituting into (5) we get $\widehat{Q}(G,\tau)<1$ for $q\geqslant 32$. If $q\leqslant 32$, then we replace the upper bound on $\alpha_{1}$, by the tighter bound $(4,q-1)\log_{p}q\cdot|\pi(\log_{p}q)|$ and we compute the exact contribution from involutory graph automorphisms, and we again verify that $\widehat{Q}(G,\tau)<1$.
Our strategy is similar for the cases where $G_{0}={\rm PSL}_{3}(q)$ or ${\rm PSp}_{4}(q)$. Let $\alpha_{1},\beta_{1},\alpha_{3},\beta_{3}$ be defined as above, and this time let $\alpha_{2}$ be the number of $G_{0}$-classes of field automorphisms of odd order and
$\beta_{2}\min\{|x^{G_{0}}|\,:\,x\in\mathcal{F},|x|\neq 2\}.$
Finally, let $\alpha_{4}$ denote the number of $G_{0}$-classes of involutory field and graph-field automorphisms and
$\beta_{4}=\min\{|x^{G_{0}}|\,:\,x\in\mathcal{F},|x|=2\}.$
Then Propositions 4.12 and 4.13 give us
$\widehat{Q}(G,\tau)\leqslant\alpha_{1}\beta_{1}^{-4/5}+\alpha_{2}\beta_{2}^{-1/5}+\alpha_{3}\beta_{3}^{-1/5}+\alpha_{4}\beta_{4}^{-1/5}$ (6)
so it suffices to find sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$. We can again infer from [41, Table 3] that there are at most $q^{2}+8q$ distinct prime order classes in $G_{I}$, none of which split in $G_{0}$, so $\alpha_{1}\leqslant q^{2}+8q$. Moreover, the same arguments we used when $G_{0}={\rm PSL}_{5}(q)$ gives $\alpha_{2}\leqslant 3\log_{p}q\cdot\log(\log_{p}q+2)$, $\alpha_{3}\leqslant 3$ and $\alpha_{4}\leqslant 6$. Next, as usual we obtain
$\beta_{1}$ $>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$
$\beta_{2}$ $>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{16/3}$
$\beta_{3}$ $>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{4}$
$\beta_{4}$ $>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{4}$
via Lemma 3.18 and one can check that these bounds are sufficient for $q\geqslant 67$. If $q\leqslant 25$, then we verify the claim using Magma, so it is left to verify that $\widehat{Q}(G,\tau)<1$ for $27\leqslant q\leqslant 64$. In this case, the tighter bounds $\alpha_{2}\leqslant(3,q-\epsilon)\log_{p}q|\pi(\log_{p}q)|$, $\alpha_{3}\leqslant(3,q-\epsilon)$ and $\alpha_{4}\leqslant 2(3,q-\epsilon)$ are sufficient.
Finally, suppose that $G_{0}={\rm PSp}_{4}(q)$. We verify the claim using Magma if $q\leqslant 11$ so we may assume that $q\geqslant 13$ for the remainder. Let $\alpha_{i},\beta_{i}$ be defined as in the case $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ for $i\in[4]$. Then Propositions 4.14 and 4.15 imply that (6) still applies, so we will again derive sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$.
By inspecting [41, Table 3] we deduce that there are at most $q^{2}+12q$ distinct $G_{I}$-classes of unipotent and semisimple elements and each unipotent class can split into at most 44 $G_{0}$-classes, so we get $\alpha_{1}\leqslant q^{2}+12q+9$. Moreover, Lemma 3.18 implies that $\beta_{1}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$. Next, using [22, 3.4.15] we find
$\alpha_{2}\leqslant(2,q-1)(r-1)\pi(\log_{p}q)\leqslant 2\log_{p}q\cdot\log(\log_{p}q+2).$
and once again via Lemma 3.18, we get $\beta_{2}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{20/3}$. Similarly, we deduce that there are at most 22 $G_{0}$-classes of involutory field automorphisms by inspecting [22, 3.4.15], and [22, 3.4.16] tells us that there is an additional $G_{0}$ class of involutory graph-field automorphisms when $p=2$, so $\alpha_{4}\leqslant 3$. Finally, we get $\beta_{4}>\frac{1}{4}q^{5}$ from Lemma 3.18, and trivially, $\alpha_{3}=\beta_{3}=0$, since $G$ contains no graph automorphisms. One can check that those bounds are sufficient for $q\geqslant 13$.
Finally, we observe that $\widehat{Q}(G,\tau)\rightarrow 0$ as $q\rightarrow\infty$ in all cases, so $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$. ∎
Proposition 6.6.
The conclusion to Theorem 3.1 holds for the groups with $G_{0}\in\mathcal{A}_{5}$.
Proof.
First we note that the claim can easily be verified using Magma for $q\leqslant 31$, so we may assume for the remainder of the proof that $q\geqslant 32$. If $\tau=(H_{1},\ldots,H_{4})$ is a non-standard tuple of $G$, then Proposition 3.17 yields
$\hat{Q}(G,\tau)\leqslant|G|\cdot\left(\frac{4}{3q}\right)^{4}+\alpha\cdot\left(\frac{2+q^{1/2}(q^{1/2}+1)}{q^{1/2}(q+1)}\right)^{4}$
where $\alpha$ denotes the number of involutory field automorphisms. We now note that
$\alpha\leqslant\frac{|{\rm PGL}_{2}(q)|}{|{\rm PGL}_{2}(q^{1/2})|}$
and we can now check that the above bound gives $\widehat{Q}(G,\tau)<1$ for all $q\geqslant 32$. Moreover, we can easily see that $\widehat{G}(G,\tau)\rightarrow 0$ as $q\rightarrow\infty$, and so $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$, as required. ∎
Proof of Theorem 3.1.
This follows by combining Propositions 6.1, 6.2, 6.3, 6.4, 6.5, $6.$ ∎
7. Preliminaries on exceptional groups
For the remainder of the paper, we may assume that $G$ is an almost simple exceptional group of Lie type with socle $G_{0}$ over $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$ and $\tau$ is a $6$-tuple of core-free subgroups of $G$. As in the previous section, our main aim is to prove Theorem 1. In particular, we will show that the stronger statement below holds, which in combination with Theorem 3.1 for classical groups, completes the proof of Theorem 1.
Theorem 7.1.
Let $G$ be a finite almost simple exceptional group of Lie type and let $\tau$ be a $6$-tuple of core-free subgroups of $G$. Then $\tau$ is regular, and moreover, $Q(G,\tau)\rightarrow 0$ as $q\rightarrow\infty$.
We start by fixing some initial notation. As before, we will write $\bar{G}$ for the ambient algebraic group (of the same type as $G$) over the algebraic closure of $\mathbb{F}_{q}$, so $G_{0}=(G_{\sigma})^{\prime}$ for some appropriate Steinberg endomorphism $\sigma$. Also, for a closed subgroup $\bar{H}$ of $\bar{G}$, we will write $\bar{H}^{0}$ to denote the connected component of $\bar{H}$ containing the identity. We will also be using the standard notation $E_{6}^{\epsilon}(q)$ for the finite exceptional groups of type $E_{6}$, where $E_{6}^{+}(q)=E_{6}(q)$ and $E_{6}^{-}(q)={}^{2}E_{6}(q)$.
We will denote the number of elements of order $r$ in some group $H$ by $i_{r}(H)$. We will often denote the set of prime numbers at most $r$ by $\pi_{r}$ and the number of odd primes at most $r$ by $\pi^{\prime}_{r}$.
For the remainder of the paper we will write $\mathcal{M}_{G}$ for the set of core-free maximal subgroups of $G$ and $\mathcal{P}_{G}$ for the set of parabolic maximal subgroups of $G$, and finally, following the notation in [51], we will often write $x=s,u,\phi,\tau,\phi\tau$ to mean that $x$ is semisimple, unipotent, a field, graph, or graph-field automorphism respectively.
7.1. Subgroup structure
We will repeatedly refer to standard sources in the literature for information regarding maximal subgroups of finite exceptional groups. We recall some important results in this section. We write $\mathcal{M}_{G}$ as a union $\mathcal{C}\cup\mathcal{S}$ of two collections, where $\mathcal{C}$ is the collection of maximal subgroups $H$ of $G$ of the following types:
  • (a)
    $H=N_{G}(\bar{H}_{\sigma})$ for some maximal $\sigma$-stable positive dimensional closed subgroup $\bar{H}$ of $\bar{G}$;
  • (b)
    $H$ is of the same type as $G$. That is, $H$ is either a subfield subgroup of a twisted version of $G$;
  • (c)
    $H$ is the normaliser of some exotic $r$-local subgroup for some prime $r\neq p$;
  • (d)
    $G_{0}=E_{8}(q)$, $p>5$, and ${\rm soc}(H)=A_{5}\times A_{6}$.
A description of the subgroups in $(a)$ and $(b)$ can be found in [60] and the groups in $(c)$ have been classified (up to conjugacy) in [31]. Finally, the group in case $(d)$ was first described by Borovik in [9] and is known as the Borovik subgroup. Note that the work of Borovik in [9] in combination with the work of Liebeck and Seitz in [58] implies that all maximal subgroups in $\mathcal{S}$ are almost simple.
At the time of writing, all maximal subgroups of $G$ have been determined up to conjugacy unless $G_{0}=E_{7}(q)$ or $E_{8}(q)$. We now give a few details for each case and the original sources where the information can be found:
  • $\mathbf{G_{0}={}^{2}B_{2}(q):}$ In this case, Suzuki determines the maximal subgroups of $G$ up to conjugacy in [74] and Bray, Hold and Roney-Dougal record them in [12, Table 8.16].
  • $\mathbf{G_{0}={}^{2}G_{2}(q)^{\prime}\textbf{ or }{}^{3}D_{4}(q):}$ Here the subgroups in $\mathcal{M}_{G}$ up to conjugacy are determined by Kleidman in [48] and [50] respectively, and unless $q=3$, they are listed in [12, Tables 8.43, 8.51].
  • $\mathbf{G_{0}={}^{2}F_{4}(q)^{\prime}:}$ In the case $q\geqslant 8$ the maximal subgroups of $G$ are determined by Malle in [66] apart from the omission of three classes of maximal subgroups isomorphic to ${\rm PGL}_{2}(13)$ when $q=8$. This is straightened out in [33, Remark 4.11].
    If $q=2$, then the subgroups in $\mathcal{M}_{G}$ are determined up to conjugacy by Wilson in [76] with the exception of the omission of a unique class of maximal subgroups isomorphic to ${\rm U}_{3}(2).2$.
  • $\mathbf{G_{0}=G_{2}(q)^{\prime}:}$ Here the members of $\mathcal{M}_{G}$ are determined up to conjugacy in [32] for even $q$ and [48] for odd $q$. For $q=2^{f}$ and $f>1$ one can find them listed [12, Table 8.30] and for odd $q$ they are listed in [12, Tables 8.41, 8.42].
  • $\mathbf{G_{0}=F_{4}(q):}\textbf{ or }\mathbf{{E}_{6}^{\epsilon}(q):}$ The definitive source for those groups is [33], where Craven determines their maximal subgroups up to conjugacy. The members of $\mathcal{C}$ are listed in [33, Table 7, 8] and the members of $\mathcal{S}$ are listed in [33, Tables 1-3].
  • $\mathbf{G_{0}=E_{7}(q):}$ The members of $\mathcal{M}_{G}$ are discussed in [34] in this case. In particular, the members of $\mathcal{C}$ are listed in [34, Table 4.1] up to conjugacy, and as stated in [34, Theorem ], the collection $\mathcal{S}$ consists of the groups listed in [34, Table 1.1] up to conjugacy possibly together with some additional subgroups with socle ${\rm L}_{2}(q^{\prime})$ for $q\in\{7,8,9,13\}$.
  • $\mathbf{G_{0}=E_{8}(q):}$ The subgroups in $\mathcal{C}$ have been completely determined up to conjugacy. In particular, the parabolic and maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$ are listed in [56, Tables 5.1 and 5.2], the subfield subgroups and exotic local subgroups are listed in [57, Table 1], and finally for the non-maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$, one can find the possibilities for ${\rm soc}(\bar{H}_{\sigma})$ listed in [31, Table 3]. If $p>5$ then $\mathcal{C}$ also contains the Borovik subgroup.
    It is currently an open problem to determine the members of $\mathcal{S}$ even up to isomorphism. However, significant progress has been made towards a classification. The current state of the art regarding the groups in class $\mathcal{S}$ will not be relevant to us, but we refer the reader to [24, p.19] for more information.
7.2. Conjugacy classes
We will repeatedly be using results from the literature about conjugacy classes of elements in finite exceptional groups. In this section we point the reader to the sources where the information is recorded.
  • $\mathbf{G_{0}={}^{2}B_{2}(q):}$ In this case the conjugacy classes of elements of $G$ are determined in [74].
  • $\mathbf{G_{0}={}^{2}G_{2}(q):}$ Detailed information on its conjugacy classes can be found in [75].
  • $\mathbf{G_{0}={}^{3}D_{4}(q):}$ The conjugacy classes of $G$ are recorded in [37] and [71].
  • $\mathbf{G_{0}={}^{2}F_{4}(q)^{\prime}:}$ In this case, the $G_{0}$-classes are described by Shinoda in [70].
  • $\mathbf{G_{0}=G_{2}(q)^{\prime}:}$ If $p\geqslant 5$ then the conjugacy classes of $G$ are determined in [30] and if $p<5$, then Enomoto determines the $G$-classes in [38].
  • $\mathbf{G_{0}=F_{4}(q):}$ Information on the semisimple classes of $G_{0}$ is recorded in [3, Table 23]. For unipotent classes the Foulkes functions for $G_{0}$ are given to us by Lübeck, and we discuss further in Section 9 how we use those to compute fixed point ratios.
  • $\mathbf{G_{0}=E_{6}^{\epsilon}(q),E_{7}(q)}\textbf{ or }\mathbf{E_{8}(q)}$: Our standard reference for semisimple classes is [39] when $G_{0}=E_{6}^{\epsilon}(q)$ and $E_{7}(q)$ and [40] when $G_{0}=E_{8}(q)$, where Fleischmann and Janiszczak describe the conjugacy classes of the simply connected groups of type $E_{6},E_{7}$, and $E_{8}$ respectively. For unipotent elements we again use the Foulkes functions given to us by Lübeck in [64].
Recall that $\tau$ is a $6$-tuple of core-free subgroups of $G$. As with classical groups, our central aim will be to show that $\widehat{Q}(G,\tau)<1$, and in particular bound $\widehat{Q}(G,\tau)$ by a function of $q$ that tends to zero as $q$ tends to infinity. We adopt a rather different approach on how we derive an upper bound for $\widehat{Q}(G,\tau)$, depending on the socle type of $G$ and so we will divide the various possibilities for $G_{0}$ into the following three collections, according to the proof strategy we use, and we will handle each one in turn:
$\mathcal{E}_{1}$ $=\{{}^{2}G_{2}(q)^{\prime},{}^{2}B_{2}(q),{}^{2}F_{4}(q)^{\prime},{}^{3}D_{4}(q),G_{2}(q)^{\prime}\}$
$\mathcal{E}_{2}$ $=\{E_{6}^{\epsilon}(q),E_{7}(q),F_{4}(q)\}$ (7)
$\mathcal{E}_{3}$ $=\{E_{8}(q)\}$
In all the above cases, our proof will rely on obtaining sufficient fixed point ratio estimates on all maximal subgroups of $G$, and we will be working with the bound
$\widehat{Q}(G,\tau)\leqslant\sum_{x\in\pi(|G|)}f(x,q),$
where
$f(x,q)=\max\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{M}_{G}\}$
for each choice of prime order element $x\in G$. Note that $f(x,q)$ is independent of the choice of conjugacy class representative. That is, $f(x,q)=f(x^{g},q)$ for all $g\in G$, so if $x_{1},\ldots,x_{t}$ is a list of representatives of the prime order conjugacy classes of $G$, then we will be working with the expression:
$\sum_{i=1}^{t}|x_{i}^{G}|\cdot f(x_{i},q).$
Let $H$ be a core-free maximal subgroup of $G$. The definitive source on bounds on ${\rm fpr}(x,G/H)$ is [51]. However, the bounds in [51] will not always be sufficient for our purposes.
When $H$ is non-parabolic, our standard approach will be to obtain sufficient bounds by inspecting the proofs of [25] and [19] when $H$. In certain rare cases, we will also need to improve existing bounds in the literature. For example, if $G_{0}=F_{4}(2)$ and $|H|>2^{22}$, then we will use the Character Table library in GAP to compute the precise value of ${\rm fpr}(x,G/H)$ for all $x\in G$ More details can be found in the proof of Proposition 8.4.
If $G$ has socle in $\mathcal{E}_{2}$, then obtaining fixed point ratio bounds that hold for all non-parabolic subgroups of $G$ is slightly more involved than in other cases, due to the larger number of possibilities for $H$, and so we carefully explain how to do this in Section 8. Surprisingly, if $G_{0}=E_{8}(q)$, then obtaining sufficient bounds is not as troublesome and can be done by careful inspection of [25]. These bounds may be of independent interest, so we record them in Tables 12, 13, 14, 15, and 16 for the reader’s convenience.
Our approach when $H$ is parabolic is rather different and we give detailed information on how we obtain bounds on ${\rm fpr}(x,G/H)$ in Section 9. If $G_{0}\in\mathcal{E}_{1}$, then the bounds in [51, Theorems 1–2] are always sufficient for unipotent and semisimple elements, and better bounds can be obtained by inspecting the proof of [51, 6.1] for field and graph-field automorphisms, if necessary. This is however not true if $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, where we will heavily use [51, 3.2] and [64] to derive better estimates for ${\rm fpr}(x,G/H)$ that are sensitive to ${\rm dim}\,x^{\bar{G}}$ for a given prime order element $x\in G$. For a unipotent element $x\in G$, it is always possible to compute precise values for ${\rm fpr}(x,G/H)$ using Lübeck’s values of the Foulkes functions for $G_{0}$ [64]. Now suppose that $x\in G$ is semisimple. If $G_{0}\in\mathcal{E}_{3}$, then an upper bound on the formula given by [51, 3.2] will be sufficient for our purposes but if $G_{0}\in\mathcal{E}_{2}$, then we will need to compute the precise value of ${\rm fpr}(x,G/H)$ to obtain our result. For field and graph-field automorphisms we will always appeal to a statement extracted from the proof of [51, 6.1] (see Proposition 9.4). Finally, for graph automorphisms when $G_{0}=E_{6}^{\epsilon}(q)$, we derive appropriate bounds in Proposition 9.5 when $q$ is odd using the approach described in  [51, 6.4], which we record in Table 18, and for even $q$, we consult [25, Table 6] and for the reader’s convenience we record this information in Table 17. We give more details in Section 9.
Organisation. The remainder of the paper is organised as follows: In Section 8 we give fixed point ratio bounds for non-parabolic actions for groups with socle in $\mathcal{E}_{2}$. In Section 9, we explain how we derive fixed point ratio estimates on parabolic subgroups for groups with socle in $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$. Finally, in Sections 10.1, 10.2, and 10.3, we prove Theorem 7.1 for groups with socle in $\mathcal{E}_{1},\mathcal{E}_{2}$ and $\mathcal{E}_{3}$, respectively.
8. Fixed point ratios for non-parabolic actions
In this section we establish fixed point ratio bounds for non-parabolic actions of groups with socle in $\mathcal{E}_{2}$. In particular, for every prime order element in $G$ we obtain a value $g(x,q)$ such that
${\rm fpr}(x,G/H)\leqslant g(x,q)$
for every $H\in\mathcal{M}_{G}\setminus\mathcal{P}_{G}$ and we record those values in Tables 12, 13, 14, 15, and 16. As mentioned above, $g(x,q)$ is readily obtained by inspecting the proofs of [25] in the majority of cases, and we carry out the necessary computations for the cases where more work is required.
For the remainder of the section, we fix $G$ with $G_{0}\in\mathcal{E}_{2}$ and $H\in\mathcal{M}_{G}\setminus\mathcal{P}_{G}$, and we proceed by establishing the bounds in question.
Proposition 8.1.
Suppose that $G_{0}=E^{\epsilon}_{6}(q)$ with $q\geqslant 3$, and $x\in G$ has prime order $r$. Then ${\rm fpr}(x,G/H)\leqslant g(x,q)$, where $g(x,q)$ is recorded in Table 12.
Proof.
Case 1. $x$ is semisimple: Here, [51, Theorem 2] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-12}$, and we claim that ${\rm fpr}(x,G/H)\leqslant q^{-14}$ if ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Let $L$ denote the simply connected group of type $E_{6}^{\epsilon}(q)$ and recall that $L=G_{0}.Z$, where $Z=Z(L)$. Also recall that $Z\cong C_{3}$ if $q\equiv 1\allowbreak\mkern 4.0mu({\operator@font mod}\,\,3)$ and $Z=1$ otherwise.
By close inspection of [39], we find that
$|C_{L}(x)|\leqslant(q^{2}-1)(q^{3}+1)(q^{2}-q+1)(q^{2}+q+1)$
when $\epsilon=+$ and ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Since $|x^{L}|$ can split into at most $G_{0}$-classes, we have
$|x^{G}|$ $\geqslant\frac{|E_{6}(q)|}{3q^{3}(q^{3}+1)(q^{2}+q+1)(q^{2}-q+1)}$
$\geqslant\frac{1}{3}q^{39}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{6}+1)(q^{8}-1)(q^{9}-1)$
$\geqslant\frac{q^{39}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{6}+1)(q^{8}-1)(q^{9}-1)}{3(q^{2}-1)}$
so by Lemma 2.1 we have $|x^{G}|>\frac{1}{6}q^{64}$.
Similarly, if $\epsilon=$-, then we find
$|C_{L}(x)|\leqslant q^{3}(q^{2}-1)(q^{3}+1)(q+1)^{4}$
and again using Lemma 2.1 we find $|x^{G}|>\frac{1}{6}q^{64}$.
If $|H|\leqslant q^{48}$, then we deduce that ${\rm fpr}(x,G/H)<2q^{-15}<q^{-14}$, as required, so now suppose $|H|>q^{48}$. Here, [33] tells us that $H$ must be of type $F_{4}$, in which case we verify that ${\rm fpr}(x,G/H)<q^{-14}$ by inspecting the proof of [25, 4.18].
Case 2. $x$ is unipotent: Here, $x\in G_{0}$. If $p>2$, then we check that the relevant bounds hold by inspecting [25, 4.14–4.20], so let us assume that $p=2$. If ${\rm dim}\,x^{\bar{G}}<40$, then the claim follows directly by [51, Theorem 2], so we may now assume that ${\rm dim}\,x^{\bar{G}}\geqslant 40$. That is, $x$ is in the $G$-class $A_{1}^{3}$. If $|H|\leqslant q^{32}$, or $H$ is of type $E_{6}^{\epsilon}(q^{1/2})$, then the claim is verified in the proofs of [25, 4.14, 4.20] respectively, so we may assume that $H=N_{G}(\bar{M}_{\sigma})$ where $\bar{M}$ is a $\sigma$-stable subgroup of $\bar{G}$.
First suppose that $\bar{M}=A_{1}A_{5}$. By Bala-Carter, the $\bar{M}$-class of $x$ corresponds to a pair $(L,P_{L^{\prime}})$, where $L$ is a Levi subgroup of $\bar{M}$ and $P_{L^{\prime}}$ is a distinguished parabolic subgroup of $L^{\prime}$. If $L$ is also a Levi subgroup of $\bar{G}$, then the $\bar{G}$-class has the same label as the $\bar{M}$-class, whilst if $L=A_{1}^{4}$, then we deduce that $x$ lies in the $G$-class $A_{1}^{3}$ via [54, Table 17]. Now it is shown in [25, 4.15] that ${\rm fpr}(x,G/H)<4q^{-\delta(x)}$, where $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}\,(x^{\bar{G}}\cap\bar{M})$, and so we compute that ${\rm fpr}(x,G/H)<4q^{-20}$.
Next consider the case $\bar{M}=D_{5}T_{1}$. Here [33, Tables 9 and 10] give $H_{0}=H\cap G_{0}={\rm P\Omega}_{10}^{\epsilon}(q)\times\frac{q-\epsilon}{e}$, where $e=(3,q-\epsilon)$. Therefore, [51, 1.3] implies that
$|x^{G}\cap H|\leqslant i_{2}(H_{0})=i_{2}({\rm Aut}({\rm P\Omega}_{10}^{\epsilon}(q))<2(q+1)q^{24}$
so the claim is satisfied.
Now suppose that $\bar{M}=T_{2}D_{4}.S_{3}$. Here we deduce via [33, Tables 9 and 10] that $x^{G}\cap H\subseteq\widetilde{H}$, where $\widetilde{H}={\rm O}_{8}^{+}(q)$ or $^{3}D_{4}(q)$. Hence, we have
$|x^{G}\cap H|\leqslant i_{2}(\widetilde{H})\leqslant 2q^{16}$
so certainly ${\rm fpr}(x,G/H)<2q^{-8}$, as required.
Finally, if $\bar{M}=F_{4}$, then we deduce via [53, Table A] that there are no involutions $x\in H$ with ${\rm dim}\,x^{\bar{G}}\geqslant 40$.
Case 3. $x$ is a graph or field or graph-field automorphism: If $r$ is odd and $|H|\leqslant q^{32}$ we observe that ${\rm fpr}(x,G/H)\leqslant q^{32-78(1-1/r)}$, as
$|x^{G}|\geqslant|E_{6}^{\epsilon}(q)\,:\,E_{6}^{\epsilon}(q^{1/r})|>\frac{1}{6}q^{78(1-1/r)}.$
In all other cases the bounds derived in the proofs of [25, 4.14–4.20] are again sufficient, so the proof is complete. ∎
Proposition 8.2.
Suppose that $G_{0}=E^{\epsilon}_{6}(2)$ and $x\in G$ has prime order $r$. Then ${\rm fpr}(x,G/H)\leqslant g(x,q)$, where $g(x,q)$ is recorded in Table 13.
Proof.
If $H$ is almost simple with socle ${\rm Fi}_{22}$, then the fusion of conjugacy classes of $H$ in $G$ is computable in GAP using the Character Table Library, so we can compute precise fixed point ratios for all $x\in G$, as noted in Section 2.2. The same is true if $G={}^{2}E_{6}(2).2$ and $H={\rm SO}_{7}(3)$. So, in view of [25, 4.13], we may assume from now on that $|H|\leqslant 2^{32}$ or $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
Case 1. $x$ semisimple: If ${\rm dim}\,x^{\bar{G}}$, then the claim follows directly by [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}\geqslant 42$. If $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=F_{4}$, then we note that $G\leqslant G_{0}.2$, and the character tables of $G_{0}$ and $H_{0}=H\cap G_{0}$ are both available in GAP. Since $x\in G_{0}$, we can compute the fusion of semisimple $H_{0}$-classes in $G_{0}$ and verify that the bounds in Table 13 are satisfied for all semisimple elements.
Next, if $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=D_{5}T_{1}$ then the bounds in the proof of [25, 4.16] suffice.
In all remaining cases, we can infer from [33, Tables 9 and 10] that $|x^{G}\cap H|\leqslant 2^{38}$ and we immediately get ${\rm fpr}(x,G/H)\leqslant 2^{38}/|x^{G}|$ when ${\rm dim}\,x^{\bar{G}}>42$, so finally let us assume that ${\rm dim}\,x^{\bar{G}}=42$, so $C_{\bar{G}}(x)^{0}=T_{1}A_{5}$.
First suppose that $|H|\leqslant 2^{32}$. We can infer from [25, Table 7] that $|x^{G}|>2^{41}$, so we have ${\rm fpr}(x,G/H)<2^{-9}$.
Now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$. If $\bar{M}=A_{1}A_{5}$, then ${\rm fpr}(x,G/H)<2^{-14}$ according to [25, 4.15].
Finally, if $\bar{M}=T_{2}D_{4}.S_{3}$, then inspection of [33, Tables 9 and 10] gives
$|x^{G}\cap H|<|H_{0}|\leqslant 18\cdot|{\rm P\Omega}_{8}^{+}(2)|<2^{33}$
where $H_{0}=H\cap G_{0}$, so the claim follows.
Case 2. $x$ is unipotent: If $x$ is in class $A_{1}$, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2^{-5}$, so next assume that $x$ is in class $A_{1}^{2}$ or $A_{1}^{3}$. If $|H|\leqslant 2^{32}$, then the claim is proved in [25, 4.14], so we may now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
If $\bar{M}=A_{1}A_{5}$, then the bound ${\rm fpr}(x,G/H)\leqslant 4\cdot 2^{-\delta(x)}$ is still valid, as highlighted in [25, 4.15], and we find that $\delta(x)=16$ if $x$ is in $A_{1}^{2}$ and $\delta(x)=20$ if $x$ is in $A_{1}^{3}$, so the claim holds.
If $\bar{M}=D_{5}T_{1}$, then we note as above that $x^{G}\cap H\subseteq{\rm\Omega}_{10}^{\epsilon}(2)$, and so using Magma we find
$|x^{G}\cap H|\leqslant i_{2}({\rm\Omega}_{10}^{\epsilon}(2))=21999615.$
One can check that this bound is sufficient in both cases. Now suppose that $\bar{M}=T_{2}D_{4}.S_{3}$. As above we have
$|x^{G}\cap H|\leqslant\max\{i_{2}({}^{3}D_{4}(2).3),i_{2}(\Omega_{8}^{+}(q).S_{3})\}\leqslant 2^{15}$
which is sufficent.
Case 3. x is an involutory graph automorphism: Here, the bounds in Table 13 can be verified by inspecting [25, 4.14–4.20], so the proof is complete.
[…]
Table 12. Bounds on ${\rm fpr}(x,G/H)$ for $G_{0}=E^{\epsilon}_{6}(q)$, $H$ non-parabolic and $q\geqslant 3$.
[…]
Table 13. Bounds on ${\rm fpr}(x,G/H)$ for $G_{0}=E^{\epsilon}_{6}(2)$ and $H$ non-parabolic.
Proposition 8.3.
Suppose that $G_{0}=E_{7}(q)$ and $x\in G$ has prime order $r$. Then ${\rm fpr}(x,G/H)\leqslant g(x,q)$, where $g(x,q)$ is recorded in Table 14.
Proof.
Appealing to [25, 4.6] we find that either $|H|\leqslant q^{46}$, or $H$ is of type $E_{7}(q^{1/2})$, or $H=N_{G}(\bar{M}_{\sigma})$, where $\bar{M}=T_{1}E_{6}.2,A_{1}D_{6},A_{7}.2$, or $A_{1}F_{4}$.
Case 1. $x$ is semisimple: We first consider the case $q\geqslant 3$, as $q=2$ needs special attention. Here, [51, Theorem 2] implies that ${\rm fpr}(x,G/H)<q^{-19}$, so the desired bound holds if ${\rm dim}\,x^{\bar{G}}\leqslant 70$. We may now thus assume that ${\rm dim}\,x^{\bar{G}}>70$. Recall that if $L$ denotes the simply connected group of type $E_{7}(q)$, then $L=G_{0}.Z$, where $Z=Z(L)$ and $Z\cong C_{2}$ if $q$ is odd and $Z=1$ if $q$ is even. By carefully inspecting [39], we find that
$C_{L}(x)\leqslant q^{21}(q+1)(q^{2}-1)(q^{3}+1)(q^{4}-1)(q^{5}+1)(q^{6}-1)(q^{7}+1)$
when ${\rm dim}\,x^{\bar{G}}>70$, so
$|x^{L}|$ $>q^{42}(q^{3}-1)(q^{4}+1)(q^{5}-1)(q^{6}+1)(q^{7}-1)(q^{17}+q^{16}+\cdots+1)$
$>q^{69}(q^{3}-1)(q^{5}-1)(q^{7}-1)$
and using Lemma 2.1, we find $|x^{L}|>\frac{1}{2}q^{82}$. Since $x^{L}$ can split into at most 22 $G_{0}$-classes, we get $|x^{G}|>\frac{1}{4}q^{82}$.
If $|H|\leqslant q^{46}$, then the trivial bound ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|$ is sufficient. Next, assume that $H$ is of type $E_{7}(q^{1/2})$ and let $H_{0}=G_{0}\cap H$. First suppose $q$ is odd. Inspecting [34, Table 4.1] tells us that $H_{0}=E_{7}(q^{1/2}).2$. Let $\xi:\alpha\mapsto\alpha^{q^{1/2}}$, so $H_{0}=\bar{G}_{\xi}$. If $C_{\bar{G}}(x)$ is connected, then a well-known corollary to the Lang-Steinberg theorem [72, I, 2.7] gives $x^{\bar{G}}\cap\bar{G}_{\xi}=x^{\bar{G}_{\xi}}$ and similarly $x^{\bar{G}}\cap\bar{G}_{\xi^{2}}=x^{\bar{G}_{\xi^{2}}}$. Hence, in view of Proposition 3.5 we have $x^{G_{0}}\cap H=x^{H_{0}}$. On the other hand, if $C_{\bar{G}}(x)$ is disconnected, then the fact that $|C_{\bar{G}}(x)\,:\,C_{\bar{G}}(x)^{0}|=2$, implies that $x^{\bar{G}}\cap H_{0}$ might be a union of two distinct $H_{0}$-classes and the same holds for $G_{0}$. We therefore get
${\rm fpr}(x,G/H)\leqslant\frac{2\max\{|y^{H_{0}}|\,:\,y\in x^{\bar{G}}\}}{\min\{|y^{G_{0}}|\,:\,y\in x^{\bar{G}}\}}$ (8)
We can now inspect Lübeck’s online data on semisimple classes [65] to inspect which elements have disconnected centraliser in $\bar{G}$ and compute ${\rm fpr}(x,G/H)$ using (8). For example, suppose that $C_{\bar{G}}(x)$ is of type $E_{6}T_{1}$. Then [65] tells us that $C_{\bar{G}}(x)$ is disconnected, and in particular we find that
$|x^{G_{0}}\cap H|\leqslant 2\frac{|E_{7}(q^{1/2})|}{2|{}^{2}E_{6}(q^{1/2})|(q^{1/2}+1)}<q^{27}$
and
$|x^{G_{0}}|>\frac{|E_{7}(q)|}{2|E_{6}(q)|(q-1)}=\frac{1}{2}q^{54}$
so the claim holds. The other cases are all similar, so we leave the details to the reader. If $q$ is even, then $G_{0}=\bar{G}_{\sigma}$ and again inspecting [34, Table 4.1] we find that $H_{0}=E_{7}(q^{1/2})$. Using [72, I, 2.7] once again, we deduce that $x^{G_{0}}\cap H_{0}$, so the bounds derived in the case where $q$ is odd are again sufficient.
Finally, if $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$, then one can check that the bounds in Table 14 hold by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4].
Now suppose that $q=2$. In this case the prime order $\bar{G}_{\sigma}$-classes and the possible centraliser types are recorded in [8, Table 2]. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the claim follows from [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>66$.
If $|H|\leqslant 2^{46}$, then one can can obtain much stronger bounds than the ones in Table 14 by arguing as in the case $q\geqslant 3$, so we may assume that $H=N_{G}(\bar{M})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $66<{\rm dim}\,x^{\bar{G}}\leqslant 84$, then we can check that ${\rm fpr}(x,G/H)<2^{-19}$ for all the remaining possibilities for $H$ by inspecting the proofs of  [25, 4.9, 4.10] and [19, 3.4]. Now if ${\rm dim}\,x^{\bar{G}}>84$, then by inspecting the same proofs, we again obtain sufficient fixed point ratio bounds that this time depend on the value of $z={\rm dim}\,Z(\bar{D}^{0})$, where $\bar{D}=C_{\bar{G}}(x)$. In particular, we find that ${\rm fpr}(x,G/H)<2^{-21}$ if $z\leqslant 3$ and ${\rm fpr}(x,G/H)<2^{-24}$ if $z>3$. The authors derive those bounds depending on $z$ by appealing to [51, 4.5], which implies
${\rm fpr}(x,G/H)<\frac{|W(\bar{G})\,:\,W(\bar{M})|\cdot|\bar{M}/\bar{M^{0}}|\cdot 2\cdot(q+1)^{z}\cdot 2}{q^{\delta(x)+z-l}(q-1)^{l}}$
where $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}(x^{\bar{G}}\cap\bar{M})$ and $l$ is the semisimple rank of $\bar{M}$.
Case 2. $x$ is unipotent: We first consider the case $p=2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 54$, then the claim follows directly by [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}>54$. In this case, $x$ does not lie in the class $A_{1}$ or $A_{1}^{2}$, and so inspecting [59, Table 22.2.2] we find that $|x^{G}|>q^{64}$.
If $|H|\leqslant q^{46}$, then we have ${\rm fpr}(x,G/H)<|H|/q^{64}<q^{-18}$. Next suppose that $H$ is of type $E_{7}(q^{1/2})$. Since $x$ is unipotent, we must have $x\in H_{0}=H\cap G_{0}$, and note that both the $H_{0}$-class and the $G_{0}$-class of $x$ are determined by the labelling of its class in $\bar{G}$. Therefore, $|x^{G_{0}}\cap H_{0}|=|x^{H_{0}}|$, and in particular, ${\rm fpr}(x,G/H)=|x^{H_{0}}|/|x^{G_{0}}|<q^{-18}$.
Finally, assume that $H=N_{G}(\bar{M}_{\sigma})$ for some $\sigma$-stable subgroup $\bar{M}$ of $\bar{G}$. If $\bar{M}=A_{1}D_{6}$, then the fusion of $H$-classes in $G$ is described in [25, Table 3] and from this we can compute precise fixed point ratios and verify that the claim holds. For example, suppose that $x$ is in class $A_{1}^{3^{\prime\prime}}$, and write $x=uy$, where $u\in A_{1}$ and $y\in D_{6}$. Then  [25, Table 3] tells us that either $u=1$ and $y$ embeds in $D_{6}$ as an $a^{\prime}_{6}$ involution (see [22, Section 3.5.4] for more information on unipotent involutions in orthogonal groups), or $u\neq 1$ (so $u$ has Jordan decomposition $J_{2}]$ on $\mathbb{F}_{q}^{2}$), and $y$ embeds in $D_{6}$ as a $c_{2}$ involution. We therefore get
$|x^{G_{0}}\cap H|\leqslant\frac{1}{2}q^{6}(q^{6}-1)(q^{8}-1)(q^{10}-1)+q^{4}(q^{2}-1)(q^{6}-1)(q^{10}-1)<q^{30}$
and so ${\rm fpr}(x,G/H)<q^{30-64}<q^{-18}$. The other cases are all similar, so we omit the details.
Finally, one can check that ${\rm fpr}(x,G/H)<q^{-18}$ for all other possibilities of $\bar{M}$ by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4].
Now assume that $p>2$. If ${\rm dim}\,x^{\bar{G}}\leqslant 66$, then the bound once again follows from [51, Theorem 2]. Now suppose that ${\rm dim}\,x^{\bar{G}}>66$. If $H$ is of type $A_{1}D_{6}$, then we verify the claim by inspecting [25, 4.10], and in every other case the claim is proved by arguing exactly as in the case $p=2$.
Case 3. $x$ is a field automorphism: If $r=2$, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant q^{-22}$, so now assume that $r$ is odd. Here Lemma 2.1 gives
$|x^{G}|=|E_{7}(q)\,:\,E_{7}(q^{1/r})|>\frac{1}{2}q^{-133(1-1/r)}\geqslant\frac{1}{2}q^{-88}.$
If $|H|\leqslant q^{46}$, then we get ${\rm fpr}(x,G/H)<\frac{1}{2}q^{-42}$, so the claim holds. If $H$ is of type $E_{7}(q^{1/2})$, then $x$ also acts as a field automorphism on $E_{7}(q^{1/2})$, and so
${\rm fpr}(x,G/H)\leqslant\frac{|E_{7}(q^{1/2})\,:\,E_{7}(q^{1/2r})|}{|E_{7}(q)\,:\,E_{7}(q^{1/r})|}<4q^{-44}.$
In all other cases, one can verify that the bounds in Table 14 are satisfied by inspecting the proofs of [25, 4.9, 4.10] and [19, 3.4], so the proof is complete. ∎
[…]
Table 14. Bounds on ${\rm fpr}(x,G/H)$ for $G_{0}=E_{7}(q)$ and $H$ non-parabolic.
Proposition 8.4.
Suppose that $G_{0}=F_{4}(q)$ and $x\in G$ has prime order $r$. Then ${\rm fpr}(x,G/H)\leqslant g(x,q)$, where $g(x,q)$ is recorded in Table 15 if $q\geqslant 3$ and Table 16 if $q=2$.
Proof.
First note that either $|H|\leqslant q^{22}$, or $H$ is one of the possibilities recorded in [25, 4.22]. The case $q=2$ will be dealt with using different methods, so for now we will assume $q\geqslant 3$.
Case 1. $x$ is semsimple: If $C_{\bar{G}_{\sigma}}(x)=B_{4}$, then the bound immediately follows from [51, Theorem 2], so it is left to deal with the remaining possibilities for $x$. First assume that $|H|\leqslant q^{22}$. If $C_{\bar{G}_{\sigma}}(x)$ has a $B_{3}$ or $C_{3}$ factor, then ${\rm dim}\,x^{\bar{G}}\geqslant 28$, and inspecting [65] gives
$|x^{G}|\geqslant\frac{|F_{4}(q)|}{|{\rm Sp}_{6}(q)||{\rm L}_{2}(q)|}=q^{14}(q^{8}+q^{4}+1)(q^{4}+1)(q^{2}+1)>q^{28}.$
It follows that ${\rm fpr}(x,G/H)<q^{-6}$, as required. In all remaining cases we find that $|x^{G}|>q^{31}$, so we get ${\rm fpr}(x,G/H)<q^{-9}$.
If $H$ is of type $F_{4}(q^{1/2})$, then we refer to [33, Table 8] and we find that $H_{0}=H\cap G_{0}=F_{4}(q^{1/2})$. Now we note that $\bar{G}$ is simply connected, and so $C_{\bar{G}}(x)$ is connected. Therefore, appealing to [72, I, 2.7] we find that $x^{G}\cap H_{0}=x^{H_{0}}$. We can therefore compute precise fixed point ratios for all semisimple elements and verify that the bounds in Table 15 are satisfied for this choice of $H$. In all other cases we appeal to the proofs of [25, 4.24–4.27] and we verify that the claim holds.
Case 2. $x$ is unipotent, field or graph-field automorphism: The bounds in Table 15 can be directly verified by inspecting [25, 4.23–4.27], unless $H$ is of type $F_{4}(q^{1/2})$. If $x$ is unipotent, then we claim that $x^{G}\cap H=x^{H_{0}}$. This follows from the fact that both the $G_{0}$ and $H_{0}$-class of $x$ are determined by the labelling of its class in $\bar{G}$, and so one can verify that the bounds in Table 15 are satisfied. Finally, if $x$ is a field (respectively, graph-field) automorphism, then $x$ also acts as a field (respectively, graph-field) automorphism on $H$, and so we get
${\rm fpr}(x,G/H)=\frac{|F_{4}(q^{1/2})\,:\,F_{4}^{\epsilon}(q^{1/2r})|}{|F_{4}(q)\,:\,F^{\epsilon}_{4}(q^{1/r})|}$
which is sufficient.
Now assume that $q=2$, so $G=G_{0}$, or $G_{0}.2$. If $x$ is an involutory graph-field automorphism, then [51, Theorem 2] implies that ${\rm fpr}(x,G/H)<2^{-6}$, so we may now assume that $x\in G_{0}$.
Recall that if $|H|>2^{22}$, then the possibilities for $H$ are determined in [25, 4.22], and so we can check that the character tables of both $G_{0}=F_{4}(2)$ and $H_{0}=H\cap G_{0}$ are computable in GAP. It follows that we can determine the fusion of $H_{0}$-classes in $G_{0}$ and we note that ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{0}}\cap H_{0}|}{|x^{G_{0}}|}$ by Lemma 2.4. Now assume that $|H|\leqslant 2^{22}$. If ${\rm dim}\,x^{\bar{G}}<28$, then as noted in [25, 4.23] $x$ is an involution belonging to one of the $\bar{G}$-classes labelled $A_{1},\tilde{A}_{1}$, or $\tilde{A}_{1}^{(2)}$, and we have ${\rm fpr}(x,G/H)\leqslant(2^{4}-2^{2}+1)^{-1}=13^{-1}$ by [51, Theorem 1] if $x\in A_{1}$ or $\tilde{A}_{1}$, and ${\rm fpr}(x,G/H)<2^{-4}$ if $x\in\tilde{A}_{1}^{(2)}$. If ${\rm dim}\,x^{\bar{G}}\geqslant 28$, then ${\rm fpr}(x,G/H)\leqslant 2^{22}/|x^{G}|$ We now set
$g(x,q)=\max\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{M}_{G}\setminus\mathcal{P}_{G}\}$
and we record this value in Table 16 for each prime order class. For semisimple elements we denote $x$ by its centraliser type, whereas for unipotent elements, we use the notation in [59]. ∎
[…]
Table 15. Bounds on ${\rm fpr}(x,G/H)$ for $G_{0}=F_{4}(q)$, $H$ non-parabolic and $q\geqslant 3$.
[…]
Table 16. bounds on ${\rm fpr}(x,G/H)$ for $G_{0}=F_{4}(2)$ and $H$ non-parabolic.
9. Fixed point ratios for parabolic actions
In this section, we describe how we obtain sufficient fixed point ratio estimates for groups with socle in $G_{0}$ over a given parabolic subgroup $H$ of $G$ when $G_{0}=\mathcal{E}_{2}\cup\mathcal{E}_{3}$. We will be using the same methods that are used and discussed in [51]. In particular, if $\Omega=G/H$, then we have
${\rm fpr}(x,G/H)=\frac{\chi(x)}{|\Omega|},$
where $\chi(x)=1_{\bar{H}_{\sigma}}^{\bar{G}_{\sigma}}$ is the corresponding permutation character. We divide the section into four subsections, where we discuss our approach for semisimple elements, unipotent elements, field and graph-field automorphisms, and graph automorphisms respectively.
9.1. Semisimple elements
Let $x\in\bar{G}_{\sigma}$ be a semisimple element and let $\Omega=G/H$. As described above, $|C_{\Omega}(x)|=\chi(x)$, where $C_{\Omega}(x)=\{\omega\in\Omega\,:\,\omega^{x}=\omega\}$ is the fixed point set of $x$ in $\Omega$.
Let $W$ denote the Weyl group of $\bar{G}$ and let $W_{\bar{H}}$ denote the Weyl group of $\bar{H}$, so $W_{\bar{H}}$ is a standard parabolic subgroup of $W$. We write $\Phi$ for the root system of $\bar{G}$ with respect to a fixed maximal torus, we let $\Pi$ be a simple system of roots for $\bar{G}$ and $\alpha_{0}$ to be the highest root in $\Phi$ with respect to $\Pi$. The possible centraliser types of semisimple elements in $\bar{G}_{\sigma}$ are in one-to-one correspondence with pairs $(J,[w])$, where $J\subsetneq\Pi\cup\{\alpha_{0}\}$ and is determined up to conjugacy in $W$ and $w]=W_{J}w$ is a conjugacy class representative of $N_{W}(W_{J})/W_{J}$, where $W_{J}$ denotes the subgroup of $W$ generated by reflections in the roots in $J$. This is all described in [36, 39, 40].
Recall that $W$ is $\sigma$-stable, and so $\sigma$ induces permutation of the elements of $W$. We say that two elements $w,w^{\prime}\in W$ are $\sigma$-conjugate if there exists $x\in W$ such that $x^{-1}w\sigma(x)$. Now let $C_{1},\ldots,C_{l}$ be the $\sigma$-conjugacy classes of $W$. As explained in [51, Section 3], the $\sigma$-stable maximal tori of $\bar{G}$ are parametrised by the $\sigma$-classes of $W$, so we will write $T_{i}$ for a representative of the $\bar{G}_{\sigma}$-class of maximal tori in $\bar{G}$ corresponding to $C_{i}$ for every $i\in\{1,\ldots,l\}$, so that $(T_{1})_{\sigma},\ldots,(T_{l})_{\sigma}$ is a list of representatives of the $\bar{G}_{\sigma}$-classes of maximal tori of $\bar{G}_{\sigma}$.
We will also write $\epsilon_{T_{i}}=(-1)^{r_{T_{i}}}$, where $r_{T_{i}}$ is the relative rank of $T_{i}$ (that is the multiplicity of $q-\epsilon$ as a divisor of $|(T_{i})_{\sigma}|$) and similarly $\epsilon_{C_{\bar{G}}(x)^{0}}=(-1)^{r_{C_{\bar{G}}(x)^{0}}}$. Finally, $w^{*}$ will denote the longest word of $W$ if $G_{0}={}^{2}E_{6}(q)$ and $w^{*}=1$ otherwise and $|(C_{\bar{G}}(x)^{0})_{\sigma}|_{p^{\prime}}$ will denote the largest factor of $|(C_{\bar{G}}(x)^{0})_{\sigma}|$ not divisible by $p$.
Let $x$ be in the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$. Using the above description of $\bar{G}_{\sigma}$-classes, Lawther, Liebeck, and Seitz provide an explicit formula for the number of fixed points of $x$ in $\Omega$ in [51, 3.2], which we record below:
Proposition 9.1.
Let $x\in G$ be semisimple. With the notation above, we have
$|C_{\Omega}(x)|=\chi(x)=\sum_{i=1}^{l}\frac{|W|}{|C_{i}|}\cdot\frac{|W_{\bar{H}}\cap C_{i}{w^{*}}^{-1}|}{|W_{\bar{H}}|}\cdot\frac{|W_{J}w\cap C_{i}|}{|W_{J}|}\cdot\frac{\epsilon_{C_{\bar{G}}(x)^{0}}\cdot\epsilon_{T_{i}}|(C_{\bar{G}}(x)^{0})_{\sigma}|_{p^{\prime}}}{|(T_{i})_{\sigma}|}$
Now using Proposition 9.1, we will derive an upper bound on $|C_{\Omega}(x)|$ that will be of use to us when $G_{0}=E_{8}(q)$. Let $\{[w_{1}],\ldots,[w_{t}]\}$ be a complete set of conjugacy class representatives in $N_{W}(W_{J})/W_{J}$ and order the $\sigma$-classes $C_{1},\ldots,C_{l}$ of $W$ so that $\epsilon_{T_{i}}=1$ for $i\leqslant s$ and $\epsilon_{T_{i}}=-1$ for $i>s$ for some $s\in\{0,\ldots,l\}$. Moreover, for every $i\in\{1,\ldots,l\}$, let $L_{i}=\max_{j\in[t]}|W_{J}w_{j}\cap C_{i}|/|W_{J}|$ and $R_{i}=\min_{j\in[t]}|W_{J}w_{j}\cap C_{i}|/|W_{J}|$. We then obtain the following estimate:
Corollary 9.2.
With the notation above, we have
$C_{\Omega}(x)\leqslant|A-B|,$
where
$A=\sum_{i=1}^{s}\frac{|W|}{|C_{i}|}\cdot\frac{|W_{\bar{H}}\cap C_{i}{w^{*}}^{-1}|}{|W_{\bar{H}}|}\cdot\frac{|(C_{\bar{G}}(x)^{0})_{\sigma}|_{p^{\prime}}}{|(T_{i})_{\sigma}|}\cdot L_{i}$
and
$B=\sum_{i=s+1}^{l}\frac{|W|}{|C_{i}|}\cdot\frac{|W_{\bar{H}}\cap C_{i}{w^{*}}^{-1}|}{|W_{\bar{H}}|}\cdot\frac{|(C_{\bar{G}}(x)^{0})_{\sigma}|_{p^{\prime}}}{|(T_{i})_{\sigma}|}\cdot R_{i}$
Proof.
This follows almost immediately from Proposition 9.1. In particular, we note that $\epsilon_{C_{\bar{G}(x)^{0}}}$ is independent of $i$, so we can omit it from all the summands, and take an absolute value of the summation at the end. Moreover, for every $i\in\{1,\ldots,l\}$, we check which conjugacy class of $N_{W}(W_{J})/W_{J}$ gives the largest intersection with $C_{i}$ and which class gives the smallest intersection in $C_{i}$, and substituting for $|W_{J}w\cap C_{i}|/|W_{J}|$ in the expression gives the required bound. ∎
Remark 9.3.
We can use the polynomial functionality in Magma to estimate $\chi(x)$ using Proposition 9.1 and Corollary 9.2.
The benefit of the upper bound in Corollary 9.2 is that it is not sensitive to the choice of $w$. This is particularly useful when working with groups with socle $E_{8}(q)$, as the number of possible pairs $(J,[w])$, which are listed in [40] is extremely large, and it would be an incredibly laborious task to consider them all separately. It turns out that the upper bound from Corollary 9.2 is sufficient for our purposes, apart from the case $q=2$. When $q=2$, the possibilities for the centraliser type for $x$ are very restricted and they are all given in [6, Section 3], so we can use Proposition 9.1 to compute the precise value of $|C_{\Omega}(x)|$.
However, the upper bound in Corollary 9.2 is not sufficient when $G_{0}\in\mathcal{E}_{2}$, in which case we use Proposition 9.1 to compute $|C_{\Omega}(x)|$ precisely in all cases.
We finish the discussion on semisimple elements by giving a description of how one can compute fixed point ratios of semisimple elements in Magma using Proposition 9.1 and Corollary 9.2. Suppose that we want to compute the fixed points for a semisimple element $x\in\bar{G}_{\sigma}$ corresponding to the pair $(J,[w])$. The key ingredient is the function fprSemisimple that takes as input a group of Lie type $G$, a set encoding a parabolic subgroup $P$ of $G$ in Magma (which in our case will be $H$), a set $J$, where $J$ is as defined above, a polynomial poly, which is the $p^{\prime}$-part of the centraliser of the element in question, and a word wrd in standard generators for $W$ that encodes $w$ in Magma. Below we provide the relevant version of fprSemisimple for the case where $\bar{G}_{\sigma}$ is untwisted. If $\bar{G}_{\sigma}$ is twisted, then one can obtain a similar function with very minor modifications. Similarly, if $G_{0}=E_{8}(q)$ and we want to use the approximation of $|C_{\Omega}(x)|$ given by Corollary 9.2, it is possible to do so only with slight modifications of the function provided.
fprSemisimple := function(G, P, J, poly, wrd)
W := WeylGroup(G);
id := Identity(W);
R := Roots(G);
Q := RationalField();
L<q> := PolynomialRing(Q, 1);
S := RationalFunctionField(L);
C := Classes(W);
A := TwistedToriOrders(G);
T := [];
for i in [1 .. #A] do
  t := A[i];
  b := 1;
  for l in t[1] do
    b := b*Evaluate(l, q);
  end for;
  Append(~T, b);
end for;
E := [];
for i in [1 .. #T] do
  f := Factorization(T[i]);
  e := 0;
  if f[1][1] eq q - 1 then
    e := f[1][2];
  end if;
  Append(~E, e);
end for;
WJ := ReflectionSubgroup(W, J);
w := wrd;
WP := StandardParabolicSubgroup(W, P);
CP := Classes(WP);
k := L!poly;
f := Factorization(k);
F := 0;
if f[1][1] eq q - 1 then
  F := f[1][2];
end if;
k := S!k;
c := Classes(WP);
A := [0: i in [1 .. #C]];
for i in [1 .. #c] do
  x := c[i][3];
  a := exists(j){j : j in [1 .. #C] | IsConjugate(W, x, C[j][3])};
  A[j] := A[j] + c[i][2];
end for;
B := [0: i in [1 .. #C]];
U := [x*w : x in WJ];
for i in [1 .. #U] do
  x := U[i];
  a := exists(j){j : j in [1 .. #C] | IsConjugate(W, x, C[j][3])};
  B[j] := B[j] + 1;
end for;
z := S!0;
for i in [1 .. #C] do
  z := z + (#W/C[i][2]) * (A[i]/#WP) * (B[i]/#WJ) * (-1)^(F+E[i]) * k/T[i];
end for;
return z;
end function;
On the right input, the function fprSemisimple will return $|C_{\Omega}(x)|$ and in this way one can compute ${\rm fpr}(x,G/H)$.
Let us see a concrete example when $G_{0}=E_{7}(q)$ and $H$ is a $P_{1}$ parabolic subgroup of $G$. Here we write $\Pi=\{\alpha_{1},\ldots,\alpha_{7}\}$ for the simple roots of $\bar{G}$ and the possible pairs $(J,[w])$ parametrising the $\bar{G}_{\sigma}$-classes are recorded in [39]. Note that Fleischmann and Janiszczak adopt a labelling of the simple roots of $W$ that is different to the standard Bourbaki labelling [11]. However, the labelling of simple roots in Magma agrees with the Bourbaki labelling, so one needs to translate the sets $J$ to Bourbaki labelling, when working with Magma. In this paper, we will always adopt the Bourbaki labelling.
Let $x\in G$ be a semisimple element whose $\bar{G}_{\sigma}$-class corresponds to the pair $(J,[w])$ for $J=\{\alpha_{0},\alpha_{1},\alpha_{2}\}$ and $w=a_{1}a_{3}$, as given in [39, p.123]. Translating the Fleischmann and Janiszczak to Bourbaki notation, we find that $a_{1}$ and $a_{3}$ are the reflections corresponding to $\alpha_{7}$ and $\alpha_{5}$ in $W$ respectively. We also see that $C_{\bar{G}_{\sigma}}(x)$ is of type $A_{2}A_{1}T_{4}$, and in particular $|C_{\bar{G}_{\sigma}}(x)|=q^{4}(q^{2}-1)^{4}(q^{3}-1)$. Finally, noting that the $P_{1}$ parabolic of $G$ is given by StandardParabolicSubgroup(W, {2, 3, 4, 5, 6, 7}) in Magma, we can run the following lines of code:
G := GroupOfLieType("E7", 3);
P := {2, 3, 4, 5, 6, 7};
W := WeylGroup(G);
id := Identity(W);
R := Roots(G);
Q := RationalField();
L<q> := PolynomialRing(Q, 1);
S := RationalFunctionField(L);
J := {1, 2, #R};
a1 := Reflection(W, 7);
a3 := Reflection(W, 5);
poly := (q^2-1)^4 * (q^3-1);
fprSemisimple(G, P, J, poly, a1*a3);
and we find that $|C_{\Omega}(x)|=q^{3}+8q^{2}+9q+8$. Then dividing by the index of $H$ in $G$ in Magma we can also obtain an expression for ${\rm fpr}(x,G/H)$ in $q$. We can perform a similar computation to obtain the fixed point ratio of any semisimple element in $\bar{G}_{\sigma}$ and any parabolic action.
9.2. Unipotent elements
Let $x\in\bar{G}_{\sigma}$ be a unipotent element and let $W$, $W_{\bar{H}}$, and $\Omega$ be as in Section 9.1. Note that we again have $|C_{\Omega}(x)|=\chi(x)$, where $\chi=1^{\bar{G}_{\sigma}}_{\bar{H}_{\sigma}}$. Now if $\widehat{W}$ denotes the set of (ordinary) irreducible characters of $W$, then [51, 2.4] gives
$\chi(x)=\sum_{\phi\in\widehat{W}}n_{\phi}R_{\phi}(x)$ (9)
where
$n_{\phi}=\langle 1^{W}_{W_{\bar{H}}},\phi\rangle=\langle 1_{W_{\bar{H}}},\phi|_{W_{\bar{H}}}\rangle_{W_{\bar{H}}}=\frac{1}{|W_{\bar{H}}|}\sum_{w\in W_{\bar{H}}}\phi(w)$ (10)
and $R_{\phi}(x)$ are the Foulkes functions of $\bar{G}_{\sigma}$. The values $n_{\phi}$ are recorded in [51, pp. 413–415] when $\bar{G}_{\sigma}$ is untwisted and $\bar{H}$ is a maximal parabolic subgroup of $\bar{G}$, and we can find them in [27, Table 8] if $G_{0}=E_{6}(q)$ and $H=P_{1,6}$, or $P_{3,5}$, or $G_{0}=F_{4}(q)$ and $H=P_{1,4}$ or $P_{2,3}$. Finally, if $G_{0}={}^{2}E_{6}(q)$, then the values $n_{\phi}$ are recorded in [25, p. 14].
The values $R_{\phi}$ of unipotent elements of order $p$ have been computed by Lübeck [64] and are given to us in GAP-readable form, so we can compute ${\rm fpr}(x,G/H)$ precisely for all groups with $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, all unipotent elements and all maximal parabolic subgroups of $G$. We take the opportunity to thank him for making this data available to us.
9.3. Field and graph-field automorphisms
Let $x\in G$ be a field automorphism of prime order $r$. We write $\bar{G}_{\sigma}=\bar{G}(q)$, $\bar{H}_{\sigma}=\bar{H}(q)$ and $C_{\bar{G}_{\sigma}}(x)=\bar{G}^{\epsilon}(q^{1/r})$. As explained in the proof of [51, 6.1], we have $x^{\bar{G}_{\sigma}}\cap\bar{H}_{\sigma}x=x^{\bar{H}_{\sigma}}$, and moreover $C_{\bar{H}_{\sigma}}=\bar{H}^{\epsilon}(q^{1/r})$, where $\bar{H}^{\epsilon}(q^{1/r})$ is the corresponding parabolic subgroup of $C_{\bar{G}_{\sigma}}(x)$ and $\epsilon=$- if $x$ is a graph-field automorphism and $\epsilon=+$ otherwise. In particular, by inspecting the proof of [51, 6.1], we get the following statement:
Proposition 9.4.
With the notation above we have:
${\rm fpr}(x,G/H)=\frac{|\bar{G}^{\epsilon}(q^{1/r})\,:\,\bar{H}^{\epsilon}(q^{1/r})|}{|\bar{G}(q)\,:\,\bar{H}(q)|}.$
We will be using this result repeatedly.
9.4. Graph automorphisms
Here $G_{0}=E_{6}^{\epsilon}(q)$ and there are two classes of graph automorphisms, namely those of type $F_{4}$, and those of type $C_{4}$. We record an upper bound $t(H,\tau,q)$ for every $H\in\mathcal{P}_{G}$ and every involutory graph automorphism $\tau\in G$. If $p=2$, then [5, Section 19] implies that either $C_{\bar{G}}(x)=F_{4}$ or $C_{F_{4}}(t)$ for a long root element $t\in F_{4}$. The value $t(H,\tau,q)$ can be obtained via [51, 2.6] when $C_{\bar{G}}(x)=F_{4}$, and if $G_{\bar{G}}(x)=C_{F_{4}}(t)$, then a bound $t(H,\tau,q)$ is recorded in [25, Table 6]. For the reader’s convenience, we record these values in Table 17.
[…]
Table 17. The value $t(H,\tau,q)$ when $p=2$.
If $p>2$, then $C_{\bar{G}}(\tau)=F_{4}$ or $C_{4}$. Sufficient fixed point ratio bounds can be obtained by working as in the proof of [51, 6.4]. In the following proposition we carry out these computations to obtain $t(H,\tau,q)$. Note that we will only need an upper bound over all parabolic subgroups for our purposes, but those bounds might be of independent interest, and so we record separate upper bounds for each choice of parabolic subgroup. The value $t(H,\tau,q)$ for each distinct pair $(H,\tau)$ is recorded in Table 18.
Proposition 9.5.
Let $G_{0}=E_{6}^{\epsilon}(q)$, where $q$ is odd, and $\tau\in G$ be an involutory graph automorphism. If $H\in\mathcal{P}_{G}$, then ${\rm fpr}(x,G/H)\leqslant t(H,\tau,q)$, where $t(H,\tau,q)$ is recorded in Table 18.
Proof.
First observe that $H=N_{G}(P)$, where $P$ is a $\tau$-stable parabolic subgroup of $\bar{G}=E_{6}$. Moreover, as shown in [51, 6.4], $C_{P}(\tau)$ is a parabolic subgroup of $C_{\bar{G}}(\tau)$. Write $P=QL$, where $Q$ is the unipotent radical of $P$ and $L$ is a Levi subgroup, and we treat each possibility for $P$ in turn.
First suppose that $P=P_{2}$, in which case $L=A_{5}T_{1}$. Here $\tau$ acts as a graph automorphism on $L^{\prime}=A_{5}$, which implies that $C_{L^{\prime}}(\tau)$ must either be of type $C_{3}$ or $A_{3}$.
If $C_{\bar{G}}(\tau)=F_{4}$, then we note that no proper parabolic subgroup of $F_{4}$ has a Levi factor containing $A_{3}$, so $C_{L^{\prime}}(\tau)=C_{3}$ and thus $C_{P}(\tau)$ is the $P_{1}$ parabolic subgroup of $F_{4}$. This implies that
${\rm fpr}(\tau,G/H)=\frac{|H\,:\,C_{H}(\tau)|}{|G\,:\,C_{G}(\tau)|}\leqslant\frac{1}{q^{6}-q^{3}+1}.$
If $C_{\bar{G}}(\tau)=C_{4}$, then either $C_{L^{\prime}}(\tau)$ is of type $C_{3}$, which corresponds to the $P_{1}$ parabolic of $C_{4}$, or $C_{P}(\tau)$ is of type $A_{3}$, in which case $C_{L^{\prime}}(\tau)$ is the $P_{4}$ parabolic of $C_{4}$, and so we find ${\rm fpr}(\tau,G/H)<2q^{-11}$.
Next assume that $P=P_{4}$, in which case $L=A_{2}^{2}A_{1}T_{1}$. In this case, $\tau$ must swaps the two $A_{2}$ factors of $L^{\prime}$ and it either centralises the $A_{1}$ factor, or acts on it as an involution with centraliser $T_{1}$. It follows that $C_{L^{\prime}}(\tau)$ is either $A_{2}A_{1}$ or $A_{2}T_{1}$.
First suppose that $C_{\bar{G}}(\tau)=F_{4}$. If $C_{L^{\prime}}(\tau)=A_{2}A_{1}$, then $C_{P}(\tau)$ is the $P_{2}$ or $P_{3}$-parabolic subgroup of $F_{4}$. We therefore get ${\rm fpr}(\tau,G/H)<2q^{-9}$. On the other hand, if $C_{L^{\prime}}(\tau)=A_{2}T_{1}$, then the Levi factor of $C_{P}(\tau)$ must contain $A_{2}$, and we compute ${\rm fpr}(\tau,G/H)<3q^{-8}$ for all choices of parabolic where this is possible.
Now let us assume that $C_{\bar{G}}(\tau)=C_{4}$. If $C_{L^{\prime}}(\tau)=A_{2}A_{1}$, then $C_{P}(\tau)$ must be the $P_{3}$ parabolic subgroup of $C_{4}$ and we compute ${\rm fpr}(\tau,G/H)<2q^{-17}$. On the other hand, if $C_{L^{\prime}}(\tau)=A_{2}T_{1}$, then the Levi factor of $C_{P}(\tau)$ must contain $A_{2}$, and by checking all possibilities for $C_{P}(\tau)$, we deduce that ${\rm fpr}(\tau,G/H)<2q^{-16}$.
We now consider the case $P=P_{1,6}$, in which case $L^{\prime}=D_{4}$ and $\tau$ acts as an involutory graph automorphism on $D_{4}$, and so $C_{L^{\prime}}(\tau)=B_{3}$ or $B_{2}A_{1}$.
Since $F_{4}$ has no parabolic subgroups with Levi factor containing $B_{2}A_{1}$, it follows that if $C_{\bar{G}}(\tau)=F_{4}$, we must have $C_{L^{\prime}}(\tau)=B_{3}$, and so $C_{P}(\tau)$ must be the $P_{4}$ parabolic of $F_{4}$. This gives ${\rm fpr}(\tau,G/H)<2q^{-9}$.
On the other hand, if $C_{\bar{G}}(\tau)=C_{4}$, then we must have $C_{L^{\prime}}(\tau)=B_{2}A_{1}$, and so $C_{P}(\tau)$ is the $P_{2}$ parabolic of $C_{4}$. From this we get ${\rm fpr}(\tau,G/H)<2q^{-12}$.
Finally let us assume that $P=P_{3,5}$, so $L^{\prime}=A_{2}A_{1}A_{1}$. Here $\tau$ swaps the two $A_{1}$ factors and it either centralises the $A_{2}$ factor, or acts on it as an involution with centraliser $A_{1}T_{1}$. Hence, we must have $C_{L^{\prime}}(\tau)=A_{2}A_{1}$, or $A_{1}^{2}T_{1}$.
If $C_{\bar{G}}(\tau)=F_{4}$ and $C_{L^{\prime}}(\tau)=A_{2}A_{1}$, then as above we get ${\rm fpr}(\tau,G/H)<2q^{-11}$. If $C_{L^{\prime}}(\tau)=A_{1}^{2}T_{1}$, then the semisimple part of the Levi factor of $C_{P}(\tau)$ contains $A_{1}^{2}$, and by checking all possibilities for which this containment is satisfied, we find ${\rm fpr}(\tau,G/H)<4q^{-9}$.
If $C_{\bar{G}}(\tau)=C_{4}$ and $C_{L^{\prime}}(\tau)=A_{2}A_{1}$, then as above we get ${\rm fpr}(\tau,G/H)<2q^{-19}$. On the other hand, if $C_{L^{\prime}}(\tau)=A_{1}^{2}T_{1}$, then the semisimple part of the Levi factor of $C_{P}(\tau)$ must contain $A_{1}^{2}$, and by inspecting all possibilities for $C_{P}(\tau)$ we find that ${\rm fpr}(\tau,G/H)<3q^{-17}$, as required. ∎
[…]
Table 18. The value $t(H,\tau,q)$ when $p$ is odd.
10. Proof of Theorem 7.1
10.1. Proof of Theorem 7.1 for groups with socle in $\mathcal{E}_{1}$
In this section we prove Theorem 7.1 for the groups with socle in $\mathcal{E}_{1}$. For any prime order element $x\in G$ our aim is to derive an upper bound $f(x,q)$, where we recall
$f(x,q)=\max\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{M}_{G}\}.$
Then noting that
$\widehat{Q}(G,\tau)\leqslant\sum_{x\in\pi}f(x,q)^{6}$ (11)
where by $\pi$ we denote the set of prime order elements in $G$, we can show that the probability in question is strictly less than some function of $q$ that tends to zero as $q$ tends to infinity and prove the required claim.
In most cases we refer to [51] to obtain an upper bound on $f(x,q)$. In the cases where the bounds in  [51] are not sufficient, sometimes trivial bounds on the orders of core-free maximal subgroups of $G$ are good enough, but in most cases we use bounds derived in [25].
Proposition 10.1.
The conclusion to Theorem 7.1 holds when $G_{0}={}^{2}G_{2}(q)^{\prime}$.
Proof.
Here $q=3^{2m+1}$ for some non-negative integer $m$. In view of the isomorphism $^{2}G_{2}(3)^{\prime}\cong{\rm L}_{2}(8)$, we may assume that $q\geqslant 27$. In this case we can infer from [51, Theorem 1] that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and therefore
$\widehat{Q}(G,\tau)<|G|\cdot(q^{2}-q+1)^{-6}<q^{-1},$
since $|G|<q^{7}\log_{3}(q)$. ∎
Proposition 10.2.
The conclusion to Theorem 7.1 holds when $G_{0}={}^{2}B_{2}(q)$.
Proof.
Here $q=2^{m+1}$ for some $m\geqslant 1$. Observe that [51, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2/3}+1)/(q^{2}+1)$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and thus
$\widehat{Q}(G,\tau)<|G|\cdot\left(\frac{q^{2/3}+1}{q^{2}+1}\right)^{6}<q^{-1}$
since $|G|<q^{5}\log_{2}(q)$. ∎
Proposition 10.3.
The conclusion to Theorem 7.1 holds when $G_{0}={}^{2}F_{4}(q)^{\prime}$.
Proof.
Here $q=2^{2m+1}$ for some non-negative integer $m$. The claim can easily be verified using Magma for $q=2$ using the function RegOrbitsPlus given in [1], so we may assume that $q\geqslant 8$ for the remainder. In particular, $G_{0}={}^{2}F_{4}(q)$.
As noted in [25, Section 4.6], we can infer from [70] that $G$ has two classes of involutions with respective representatives $t_{2}$ and $t^{\prime}_{2}$, where
$|{t_{2}}^{G}|<q^{14}\,\,\,\text{and}\,\,\,|{{t^{\prime}_{2}}^{G}}|<q^{11}.$
Moreover, [51, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant q^{-4}$ for all involutions $x\in G$ and all $H\in\mathcal{M}_{G}$, so the contribution to $\widehat{Q}(G,\tau)$ from unipotent elements is less than $2q^{14}\cdot(q^{-4})^{6}=2q^{-10}$.
On the other hand, if $x\in G$ is semisimple or a field automorphism of prime order, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2q^{-6}$, so we have
$\widehat{Q}(G,\tau)<2q^{-10}+|G|\cdot(2q^{-6})^{6}\leqslant(2+64\log_{2}q)\cdot q^{-10}<q^{-1}.$
Proposition 10.4.
The conclusion to Theorem 7.1 holds when $G_{0}={}^{3}D_{4}(q)$.
Proof.
Let $x\in G$ be an element of prime order $r$. We claim that if ${\rm dim}\,x^{\bar{G}}>20$ or $x$ is a field automorphism of prime order $r\geqslant 5$, then ${\rm fpr}(x,G/H)<(q^{2}(q^{3}-2))^{-1}$ for any $H\in\mathcal{M}_{G}$. If $H$ is parabolic, then the claim follows directly by [51, Theorem 2], so assume that $H$ is non-parabolic. If $H_{0}={}^{3}D_{4}(q^{1/2})$, then we verify that the bound holds by inspecting the proof of [25, 4.42], and in all other cases $|H_{0}|<q^{14}$, so the bound is again satisfied.
Now assume that ${\rm dim}\,x^{\bar{G}}\leqslant 20$, or $r\in\{2,3\}$ and $x$ is either a field or triality graph automorphism. In this case [51, Theorem 1] implies that ${\rm fpr}(x,G/H)<(q^{4}-q^{2}+1)^{-1}$ and we claim that there are at most $3q^{22}$ elements in $G$ with ${\rm dim}\,x^{\bar{G}}\leqslant 20$.
By inspecting [37, Table 4.4], we deduce that there are at most $q^{2}+q+1$ such semisimple classes, and we find that there are $4$ such unipotent classes by inspecting [71]. Moreover, if $x$ is a field automorphism, then either $r=2$ and we have $|x^{G}|<2q^{14}$, or $r=3$ and $|x^{G}|<2q^{56/3}<q^{20}$. Note that $G$ has at most two classes of order $3$ field automorphisms. Finally there are fewer than $4q^{20}+2q^{14}$ triality graph automorphisms, so in total we get
$(q^{2}+q+11)\cdot q^{20}+4q^{14}\leqslant 3q^{22}$
elements with ${\rm dim}\,x^{G}\leqslant 20$.
Hence, we obtain
$\widehat{Q}(G,\tau)<3q^{22}\cdot(q^{4}-q^{2}+1)^{-6}+|G|\cdot(q^{2}(q^{3}-2))^{-6}<3q^{-1}$
if $q\geqslant 5$, since $|G|<3q^{28}\cdot{\rm log}_{2}q$. If $q=3$, then we note that there are no field automorphisms and so we can replace $|G|$ with $|\bar{G}_{\sigma}|$ in the bound above and we obtain the result. ∎
Proposition 10.5.
The conclusion to Theorem 7.1 holds when $G_{0}=G_{2}(q)^{\prime}$.
Proof.
In view of the isomorphism $G_{2}(2)^{\prime}\cong{\rm U}_{3}(3)$, we may assume that $q\geqslant 3$. In particular, $G_{0}=G_{2}(q)$. Let $x\in G$ be a prime order element. If $x$ is an involutory field or graph field automorphism, or a long root element, then [51, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for any $H\in\mathcal{M}_{G}$ and we can check that there are at most $q^{8}$ such elements in $G$. For all other elements, we claim that ${\rm fpr}(x,G/H)\leqslant(q^{2}(q-1))^{-1}$.
First suppose that $x$ is a field automorphism of order $r\geqslant 3$. If $|H_{0}|\leqslant q^{6}$, then the claim trivially holds, since $|x^{G}|>\frac{1}{2}q^{28/3}$ as noted in [25, 4.29], and so
${\rm fpr}(x,G/H)<\frac{|H|}{|x^{G}|}<q^{-10/3}\log_{2}{q}.$
Next, if $H$ is of type ${\rm SL}_{3}^{\epsilon}(q)$ or $^{2}G_{2}(q)$, then we confirm the required bound holds by inspecting the proofs of [25, 4.30, 4.31] respectively, and if $H$ is of type $G_{2}(q^{1/2})$, then $x$ acts on $H_{0}$ as a field automorphism, so we obtain
${\rm fpr}(x,G/H)\leqslant\frac{|G_{2}(q^{1/r})\,:\,G_{2}(q^{1/2r})|}{|G_{2}(q)\,:\,G_{2}(q^{1/2})|}<q^{-6(1-1/r)}\leqslant q^{-4}$
and the claim holds.
It now remains to show that the bound holds if $H$ is parabolic. In this case Proposition 9.4 implies that
${\rm fpr}(x,G/H)\leqslant\frac{(q^{5}|{\rm GL_{2}}(q)|)/(q^{5/r}|{\rm GL}_{2}(q^{1/r})|)}{q^{14(1-1/r)}}<2q^{-5(1-1/r)}\leqslant 2q^{-10/3}$
which is sufficient. For all other possibilities for $x$, the claim follows immediately from [51, Theorem 2].
It now follows that
$\widehat{Q}(G,\tau)<q^{8}(q^{2}-q+1)^{-6}+\log_{2}q\cdot q^{14}\cdot(q^{2}(q-1))^{-6}<q^{-1}$
and the result follows. ∎
10.2. Proof of Theorem 7.1 for groups with socle in $\mathcal{E}_{2}$
In this section we prove Theorem 7.1 for the groups with $G_{0}\in\mathcal{E}_{2}$. In view of (11), it suffices to show that
$\sum_{x\in\mathcal{P}}f(x,q)^{6}<R(q)$
where $\mathcal{P}$ denotes the set of prime order elements in $G$ and $F$ is a function with the properties that $R(q)<1$ for all $q$, and $R(q)\rightarrow 0$ as $q\rightarrow\infty$.
Let $\mathcal{P}=\mathcal{S}\cup\mathcal{U}\cup\mathcal{F}$, where $\mathcal{S}$ denotes the set of semisimple elements of prime order in $G$, $\mathcal{U}$ denotes the class of prime order unipotent elements in $G$, and $\mathcal{F}$ denotes the prime order graph, field, and graph field automorphisms of $G$, then we may write
$\sum_{x\in\mathcal{P}}f(x,q)^{6}=\alpha+\beta+\gamma,$
where
$\alpha=\sum_{x\in\mathcal{S}}f(x,q)^{6},\,\,\,\beta=\sum_{x\in\mathcal{U}}f(x,q)^{6},\,\,\,\text{and}\,\,\,\gamma=\sum_{x\in\mathcal{F}}f(x,q)^{6}.$ (12)
Our aim is to derive suitable bounds on $\alpha$, $\beta$ and $\gamma$ to prove Theorem 7.1 for $G$, which is the strategy that we will be adopting for all groups with socle in $\mathcal{E}_{2}$.
Remark 10.6.
Let us now record an important remark about semisimple elements. Note that Fleischmann and Janiszczak work with simply connected groups of type $E_{6}$ and $E_{7}$ [39], whereas we are interested in adjoint type groups. However, for a semisimple element $x\in\bar{G}_{\sigma}$, we are only using their results on centraliser orders to compute ${\rm fpr}(x,G/H)$. As described in Section 9, we do this by appealing to Proposition 9.1, and so we are only interested in $C_{\bar{G}_{\sigma}}(x)^{0}$, which is the same both in the adjoint and in the simply connected case.
Proposition 10.7.
The conclusion to Theorem 7.1 holds when $G_{0}=E_{6}^{\epsilon}(q)$.
Proof.
For now we will assume that $q\geqslant 3$, as $q=2$ needs separate consideration. We will derive sufficient bounds for $\alpha,\beta$ and $\gamma$, as defined in (12), and to do this, we need to compute upper bounds on $m_{x}$ for each prime order element $x\in G$. First note that
$f(x,q)\leqslant\max\left(\{g(x,q)\}\cup\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{P}_{G}\}\right)=:F(x,q),$ (13)
where $g(x,q)$ is determined in Proposition 8.1. It therefore remains to consider fixed point ratio bounds for parabolic subgroups.
We first consider semisimple elements. Recall that the possible centralisers of semisimple elements in $\bar{G}_{\sigma}$ are parametrised by pairs $(J,[w])$, where $J$ and $w$ are defined in Section 9.1. The possibilities for $(J,[w])$ are listed in [39], and the number of $\bar{G}_{\sigma}$-classes with centraliser type corresponding to a fixed pair $(J,[w])$ is recorded in [65]. Therefore, if $\{x_{1},\ldots,x_{t}\}$ is a complete set of representatives of $\bar{G}_{\sigma}$-classes of each centraliser type and $c_{i}$ denotes the number of $\bar{G}_{\sigma}$-classes corresponding to a pair $(J,[w])$, then we can write
$\alpha\leqslant\sum_{i=1}^{t}c_{i}\cdot|x_{i}^{\bar{G}_{\sigma}}|\cdot f(x_{i},q)^{6}\leqslant\sum_{i=1}^{t}c_{i}\cdot|x_{i}^{\bar{G}_{\sigma}}|\cdot F(x_{i},q)^{6}.$ (14)
Now recall that using Magma, we can compute the exact value of ${\rm fpr}(x,G/H)$ for every parabolic subgroup $H\leqslant G$, which is given by the formula in Proposition 9.1, and so we can compute the value of $F(x_{i},q)$ for every $i\in[t]$. Moreover, we can read off $c_{i}$ for every $i\in[t]$ from [65], and using Magma we can compute the polynomial expression
$\sum_{i=1}^{t}c_{i}\cdot|x_{i}^{\bar{G}_{\sigma}}|\cdot F(x_{i},q)^{6}$ (15)
and check that it is less than $q^{-1}$ for all values of $q\geqslant 3$.
Let us see an example. Assume $\epsilon=+$ and consider the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$ with $J=\{\alpha_{0},\alpha_{2},\alpha_{5},\alpha_{6}\}$ and $w=a_{1}a_{2}$ as given in [39]. Let $y$ be a representative of a $\bar{G}_{\sigma}$-class corresponding to this pair. Inspecting [39] tells us that $C_{\bar{G}_{\sigma}}(y)$ is of type $A_{2}^{2}T_{2}$, and in particular, $|C_{\bar{G}_{\sigma}}(y)|=|{\rm SL}_{3}(q)|^{2}\Phi_{3}(q)$, where $\Phi_{3}(q)=q^{2}-q+1$ denotes the third cyclotomic polynomial. We now turn to [65] and we find that the entry corresponding to the class of $y$ is the one with label $(i,j,k)=(12,1,5)$, and from there we can read the number of classes.
We will now derive an upper bound for $\beta$. As described in Section 9.2, in this case we can compute precise fixed point ratios for parabolic subgroups of $G$ using the data in [64], and so using GAP, we can compute the expression
$\beta\leqslant\sum_{i=1}^{r}|x_{i}^{\bar{G}_{\sigma}}|\cdot F(x_{i},q)^{6}$
where $x_{1},\ldots,x_{r}$ represent the prime order unipotent classes of $G$. In this way, we find that
$\beta<\begin{cases}0.43\text{ if }q=3,4\\ q^{-1}\text{ if }q\geqslant 5\end{cases}$
We finally compute a bound on $\gamma$. Let $\gamma_{1}$ denote the contribution from involutory field and graph-field automorphisms, $\gamma_{2}$ denote the odd order field automorphisms and $\gamma_{3}$ denote the contribution from graph automorphisms.
If $\bar{H}\leqslant\bar{G}$ is parabolic and $x$ is a field or graph-field automorphism of prime order $r$, then Proposition 9.4 gives
${\rm fpr}(x,G/H)\leqslant\frac{|E_{6}^{\epsilon}(q^{1/r})\,:\,\bar{H}^{\epsilon}(q^{1/r})|}{|E_{6}(q)\,:\,\bar{H}(q)|}.$
where we recall that $\bar{H}(q)=H_{\sigma}$ and $\bar{H}(q^{1/r})=C_{\bar{H}_{\sigma}}(x)$. Note that this implies that ${\rm fpr}(x,G/H)$ is maximised when $|\bar{H}(q)|$ (and consequently $|\bar{H}(q^{1/r})|$) is maximised, and we certainly have $|\bar{H}(q)|\leqslant|P_{1}(q)|$, where $P_{1}$ is the $P_{1}$-parabolic subgroup of $E_{6}$. Moreover, if $r=2$, then we note that the fixed point ratio of a graph-field automorphism is greater than the fixed point ratio of a field automorphism. We therefore find
$\gamma_{1}\leqslant\frac{3+\epsilon}{2}\cdot\left(\frac{|E_{6}^{\epsilon}(q)|}{|{}^{2}E_{6}(q^{1/2})|}\right)^{-5}\left(\frac{|P_{1}(q)|}{|P_{1}(q^{1/r})|}\right)^{6}$
and
$\gamma_{2}\leqslant\sum_{r\in\pi^{\prime}_{13}}(r-1)\left(\frac{|E^{\epsilon}_{6}(q)|}{|E^{\epsilon}_{6}(q^{1/r})|}\right)^{-5}\left(\frac{|P_{1}(q)|}{|P_{1}(q^{1/r})|}\right)^{6}+q^{78}\log_{2}{q}\cdot\left(\frac{|E^{\epsilon}_{6}(q^{1/17})\,:\,P_{1}(q^{1/17})|}{|E_{6}^{\epsilon}(q)\,:\,P_{1}(q)|}\right)^{6}$
where recall that $\pi^{\prime}_{13}$ denotes the set of odd prime numbers at most $13$.
Finally, we infer from Propositions 8.1, 8.2, and 9.5 that if $x$ is an involutory graph automorhpism, then $f(x,q)\leqslant q^{-5}$ if $C_{\bar{G}}(x)=F_{4}$ and $f(x,q)\leqslant 12q^{-10}$ if $C_{\bar{G}}(x)=C_{4}$. If $q>4$, then we find
$\gamma_{3}\leqslant q^{26}(q^{-5})^{6}+q^{42}(12q^{-10})^{6}$
If $q\geqslant 5$, then one checks that
$\gamma=\gamma_{1}+\gamma_{2}+\gamma_{3}<q^{-1}.$
Finally, if $q=3$ or $4$, then by only considering the contribution by involutory graph automorphisms (and field and graph-field automorphisms if $q=4$), we easily compute $\gamma\leqslant 0.025$ using Magma.
It follows that $\widehat{Q}(G,\tau)<1$ for all $q$, and moreover, $\widehat{Q}(G,\tau)<3q^{-1}$ for all $q\geqslant 5$, so $\widehat{Q}(G,\tau)\rightarrow 0$ as $q\rightarrow\infty$.
Now assume that $q=2$. Here, the value $g(x,q)$ can be obtained by Proposition 8.2. For parabolic subgroups, the unipotent $\bar{G}_{\sigma}$-classes and the tables of Foulkes functions are given to us again by Lübeck in [64], so we can compute precise fixed point ratios for involutions in parabolic actions using GAP.
The odd order semisimple classes are listed in [25, Table 9]. From this table we can recover the pair $(J,[w])$ corresponding to each semisimple $\bar{G}_{\sigma}$-class via [39]. More specifically, the centraliser type for each class is listed in [25, Table 9], so we can immediately recover $J$ by consulting [39], and by inspecting the possible centraliser orders for this particular choice of $J$ we can then recover $w$. For example, assume that $\epsilon=+$ and let us look at the fourth column in [25, Table 9]. In this case our element $y$ has centraliser type $A_{3}T_{3}$. We then turn to [39] and we find that $J=\{\alpha_{0},\alpha_{2},\alpha_{4}\}$ and by computing all possible centraliser orders we deduce that $w=a_{1}a_{17}a_{18}$. In this way we can compute precise fixed point ratios for semisimple elements, as in the case $q\geqslant 3$.
Finally, sufficient fixed point ratio bounds for involutory graph automorphisms are given by Proposition 9.5.
If $\epsilon=$-, then we can show that $\widehat{Q}(G,\tau)<1$ for any $6$-tuple $\tau$ of $G$ using almost the same strategy as in the case $q\geqslant 3$. The only difference is that we only considering prime order elements when finding upper bounds for $\alpha$ and $\beta$. In particular, (13) still holds, where, as mentioned above, $g(x,q)$ is obtained via Proposition 8.2. Then using (15), we find that $\alpha\leqslant 0.1$ and by using Lübeck’s data on unipotent elements we find that $\beta\leqslant 0.055$. Finally, the contribution from involutory graph automorphisms is no more than $0.125$, and hence $\widehat{Q}(G,\tau)<1$.
The case $\epsilon=+$ requires special attention. As pointed out in [25], if $G_{0}=E_{6}(2)$ and $\tau$ is a conjugate $6$-tuple containing copies of $P_{1}$ or $P_{6}$, then $\widehat{Q}(G,\tau)>1$. Therefore, it is not possible to prove the claim using the same methods as in the case $q\geqslant 3$, so we adopt a slightly different approach. We find that it is possible to construct $P_{1}$ and $P_{6}$ in Magma, and using random search, one can show that every $6$-tuple of $G$ only containing copies of $P_{1}$ and $P_{6}$ is regular. We may therefore assume that $\tau$ contains at least one entry not isomorphic to $P_{1}$ or $P_{6}$.
For $x\in G$, we define
$l(x,q)=\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{M}_{G}\setminus\{P_{1},P_{6}\}\},$
and we note that
$l(x,q)\leqslant\max\left(\{g(x,q)\}\cup\{{\rm fpr}(x,G/H)\,:\,H\in\mathcal{P}_{G}\setminus\{P_{1},P_{6}\}\}\right)=:L(x,q)$
Further, observe that our assumption on $\tau$ gives
$\widehat{Q}(G,\tau)\leqslant\sum_{x\in\mathcal{P}}f(x,q)^{5}l(x,q)\leqslant\alpha^{\prime}+\beta^{\prime}+\gamma^{\prime},$
where
$\alpha^{\prime}=\sum_{x\in\mathcal{S}}f(x,q)^{5}l(x,q),\,\,\,\beta^{\prime}=\sum_{x\in\mathcal{U}}f(x,q)^{5}l(x,q),\,\,\,\text{and}\,\,\,\gamma^{\prime}=\sum_{x\in\mathcal{F}}f(x,q)^{5}l(x,q).$
We can now estimate $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ from above using the methods described above, and we find $\alpha^{\prime}\leqslant 0.053$, $\beta^{\prime}\leqslant 0.6$, and $\gamma^{\prime}\leqslant 0.125$, so $\widehat{Q}(G,\tau)<1$, and the proof is complete. ∎
Proposition 10.8.
The conclusion to Theorem 7.1 holds if $G_{0}=E_{7}(q)$.
Proof.
We will again estimate $\alpha,\beta,\gamma$, as defined in (12) to derive a bound on $\widehat{Q}(G,\tau)$. Our approach here is almost the same as in the case $G_{0}=E_{6}^{\epsilon}(q)$ for $q\geqslant 3$. In particular, (13) applies here as well, where $n_{x}$ is obtained this time via Proposition 8.3 for each choice of prime order element $x\in G$.
We first consider semisimple elements. The pairs $(J,[w])$ encoding the possible centraliser types of semisimple elements in $\bar{G}_{\sigma}$ are again determined in [39] and so we can compute precise fixed point ratios for each centraliser type using Magma, as described in Section 9.1. Now let $x_{1},\ldots,x_{t}$ be representatives of the semisimple $\bar{G}_{\sigma}$-classes of each centraliser type and $c_{i}$ be the number of $\bar{G}_{\sigma}$-classes of centraliser type $C_{\bar{G}_{\sigma}}(x_{i})$ for each $i\in[t]$. Then we get
$\alpha\leqslant\sum_{i=1}^{t}c_{i}\cdot|x_{i}^{G}|\cdot f(x_{i},q)^{6}\leqslant\sum_{i=1}^{t}C_{i}\cdot|x_{i}^{G}|\cdot f(x_{i},q)^{6}$
where $C_{i}$ is an upper bound on $c_{i}$ for all $i\in[t]$, so it is left to compute sufficient bounds $C_{i}$ for all $i\in[t]$ to estimate $\alpha$. Note that the exact number of classes corresponding to a given pair $(J,[w])$ is given to us by [65]. However, crude bounds are good enough for most cases, and due to the large number of possible pairs, we will be working with upper bounds instead.
We find in [65] that the total number of semisimple classes in $G_{\sigma}$ is no more than $q^{7}+q^{4}$ and unless $C_{G}(x)$ has an $E_{6},D_{6}$, or $E_{7}$ factor, it suffices to take $C_{i}=q^{7}+q^{4}$. In the final few cases we can estimate $c_{i}$ by inspecting [65] and we give a bound $C_{i}$ in Table 19. In particular, we record the type of $C_{\bar{G}}(x_{i})$, the set $J$ and $w$ and an upper bound on the number of classes of such type. Recall that we use the standard Bourbaki labelling [11] when recording the set $J$ in this paper, instead of the notation used in [39]. Now using Magma, we can compute the expression
$\alpha\leqslant\sum_{i=1}^{t}C_{i}\cdot|x_{i}^{G}|\cdot F(x_{i},q)^{6}$
and we find that $\alpha<q^{-1}$, and in particular $\alpha\leqslant 0.25$ if $q=2$.
For unipotent elements, we can again use the GAP-readable data provided by Lübeck in [64] to compute precise fixed point ratios. If $\{x_{1},\ldots,x_{r}\}$ is a complete set of representatives of the $\bar{G}_{\sigma}$-classes of elements of order $p$, then using GAP we compute
$\beta\leqslant\sum_{i=1}^{r}|x_{i}^{G}|\cdot F(x_{i},q)^{6}<\begin{cases}0.61\,\text{ if }\,q=2\\ 1/q\,\text{ if }\,q\geqslant 3\end{cases}$
Finally, if $x$ is a field automorphism of prime order $r$, then sufficient fixed point ratio bounds on non-parabolic subgroups are obtained in Proposition 8.3, and if $H\leqslant G$ is parabolic, then Proposition 9.4 gives
${\rm fpr}(x,G/H)\leqslant\frac{|E_{7}(q^{1/r})\,:\,\bar{H}(q^{1/r})|}{|E_{7}(q)\,:\,\bar{H}(q)|}.$
where $\bar{H}(q)$ and $\bar{H}(q^{1/r})$ are defined as in Section 9.3. By computing a bound for each parabolic subgroup and comparing with $g(x,q)$, we find that $f(x,q)$ is achieved when $\bar{H}$ is the $P_{7}$ parabolic of $\bar{G}$ for all $r$, and so we get
$\gamma\leqslant\sum_{r\in\pi_{11}}(r-1)\left(\frac{|E_{7}(q)|}{|E_{7}(q^{1/r})|}\right)^{-5}\left(\frac{|P_{7}(q)|}{|P_{7}(q^{1/r})|}\right)^{6}+q^{133}\log_{2}{q}\cdot\left(\frac{|E_{7}(q^{1/13})\,:\,P_{7}(q^{1/13})|}{|E_{7}(q)\,:\,P_{7}(q)|}\right)^{6}<q^{-1}$
for all $q\geqslant 4$.
It follows that $\widehat{Q}(G,\tau)<1$ for all $q$, with $\widehat{Q}(G,\tau)<3q^{-3}$ for all $q\geqslant 3$. ∎
[…]
Table 19. The value $C_{i}$ for $E_{7}(q)$-classes that contain an $E_{6}$ or $D_{6}$ factor, as defined in Proposition 10.8.
Proposition 10.9.
The conclusion to Theorem 7.1 holds when $G_{0}=F_{4}(q)$.
Proof.
We use the same approach we used in the proof of Proposition 10.7 when $q\geqslant 3$. In particular, (13) holds in this case as well, where $g(x,q)$ is obtained this time via Proposition 8.4 for each choice of prime order element $x\in G$, so we can estimate $\alpha,\beta$, and $\gamma$, as defined in (12) to show that $\widehat{Q}(G,\tau)<1$.
First let us consider semisimple elements. In this case the pairs $(J,[w])$ encoding the possible centraliser types of semisimple elements can be found in [3, Table 23] and the number of classes of each centraliser type can be found in [65] as usual, so we can compute a bound on $\alpha$ via (15) and we find that $\alpha<q^{-1}$.
The contribution to $\widehat{Q}(G,\tau)$ from unipotent elements can be estimated in the usual way. Sufficient fixed point ratio bounds for non-parabolic subgroups can be obtained via Proposition 8.4 and GAP-readable data that allow us to compute precise fixed point ratios for parabolic subgroups are again given to us by Lübeck [64]. Therefore, if $\{x_{1},\ldots,x_{r}\}$ is a complete list of representatives for the conjugacy classes of prime order unipotent elements in $G$, we can compute the expression
$\sum_{i=1}^{r}|x_{i}^{G}|\cdot F(x_{i},q)^{6}$
using GAP and we deduce that $\beta<q^{-1}$.
It is only left to obtain a sufficient bound for $\gamma$. Let $x$ be a field or graph-field automorphism of prime order $r$. We have already obtained a sufficient fixed point ratio bound for non-parabolic subgroups in Proposition 8.4, and for a parabolic subgroup $H$ of $G$ we can once again appeal to Proposition 9.4 to get
${\rm fpr}(x,G/H)\leqslant\frac{|F_{4}^{\epsilon}(q^{1/r})\,:\,\bar{H}^{\epsilon}(q^{1/r})|}{|F_{4}(q)\,:\,\bar{H}(q)|}.$
In this way, we find that $f(x,q)\leqslant q^{-6}$ if $r=2$, whilst $f(x,q)$ is achieved when $\bar{H}=P_{1}$ if $r\geqslant 3$. Noting that $|x^{G}|<2q^{26}$ if $x$ is an involutory field or graph-field automorphism, we find
$\gamma\leqslant 4q^{26}(q^{-6})^{6}+\sum_{r\in\pi^{\prime}_{17}}(r-1)\left(\frac{|F_{4}(q)|}{|F_{4}(q^{1/r})|}\right)^{-5}\left(\frac{|P_{1}(q)|}{|P_{1}(q^{1/r})|}\right)^{6}+2q^{52}\log_{2}{q}\cdot\left(\frac{|F_{4}(q^{1/19})\,:\,P_{1}(q^{1/19})|}{|F_{4}(q)\,:\,P_{1}(q)|}\right)^{-6}<q^{-1},$
where we recall that $\pi^{\prime}_{17}$ is the set of odd primes at most 17. Therefore $\widehat{Q}(G,\tau)<1$ and if $q\geqslant 3$, then $\widehat{Q}(G,\tau)<3q^{-1}$. ∎
10.3. Proof of Theorem 7.1 for groups with socle in $\mathcal{E}_{3}$
In this final section we prove Theorem 7.1 for groups with socle $G_{0}=E_{8}(q)$, which completes the proof of Theorem 1.
Proposition 10.10.
The conclusion to Theorem 7.1 holds if $G_{0}=E_{8}(q)$.
Proof.
We will divide the prime order elements of $G$ into five collections $\mathcal{L}_{1},\ldots,\mathcal{L}_{5}$. In particular, we have $x\in\mathcal{L}_{i}$ if and only if $x\in\bar{G}_{\sigma}$ and $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$, or $x$ is a field automorphism with $q^{a_{i}}<|x^{G_{0}}|<q^{b_{i}}$, where $a_{i}$ and $b_{i}$ are listed in Table 20 for all $i\in\{1,\ldots,5\}$. We claim that if $x\in\mathcal{L}_{i}$, then $f(x,q)\leqslant c_{i}$, where $c_{i}$ is listed in Table 20 for all $i$. Moreover, we claim there are no more than $d_{i}$ prime order elements in $\mathcal{L}_{i}$ for all $i\in\{1,\ldots,5\}$, where $d_{i}$ is again listed in Table 20 for all $i$. This would in turn imply that
$\widehat{Q}(G,\tau)<\sum_{i=1}^{5}d_{i}c_{i}^{6}<q^{-1}$
which proves the claim.
First let us record some lower bounds on $|x^{G_{0}}|$ for a prime order element $x\in G$, which we will need below. In particular, we claim that if $x\in\mathcal{L}_{i}$ for some $i\in\{1,\ldots,5\}$, then $q^{88}/c_{i}<|x^{G}|<q^{b_{i}}$. All cases are similar, so we look at the case $i=2$, and we leave the remaining cases for the reader to check. If $x\in\mathcal{L}_{2}$ is a field automorphism, then the claim follows straight from the definition, so we may assume that $x\in\bar{G}_{\sigma}$. First suppose that $x$ is semisimple. In this case, inspecting [40] tells us that the possible centraliser types for $x$ are
$D_{5},A_{4}^{2},A_{6},D_{5}A_{1},A_{6}A_{1},D_{5}A_{1}^{2},D_{5}A_{2},D_{5}A_{3},A^{\prime}_{7},A^{\prime\prime}_{7}\text{ and }A_{7}A_{1}$
and from this we deduce that
$|x^{G_{0}}|\geqslant\frac{|E_{8}(q)|}{|{\rm SU_{8}(q)}||{\rm SL}_{2}(q)|}>q^{139}(q^{28}+q^{27}+\dots+q^{2}+1)(q^{3}-1)(q^{7}-1)(q^{5}-1)>q^{125}.$
Similarly, we find that
$|x^{G_{0}}|\leqslant\frac{|E_{8}(q)|}{|{\rm SO}_{10}^{-}(q)|(q-1)^{3}}<2q^{200}$
by Lemma 2.1, so the claim follows.
Now suppose that $x$ is unipotent. In this case inspecting [59, Table 22.2.1] tells us that the smallest class of elements in $\mathcal{L}_{2}$ is $A_{3}A_{2}A_{1}$ and the largest is $D_{5}$. We therefore find
$|x^{G_{0}}|=\frac{|E_{8}(q)|}{q^{60}|{\rm SL}_{2}(q)|^{2}}>q^{125}$
Similarly, we get
$|x^{G_{0}}|\leqslant\frac{|E_{8}(q)|}{q^{27}|{\rm Sp}_{6}(q)|}<q^{202}$
We now show that $f(x,q)\leqslant c_{i}$ for all $i\in\{1,\ldots,5\}$. Fix an $i\in\{1,\ldots,5\}$, a prime order element $x\in G$ such that $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ and $H\in\mathcal{M}_{G}$. If $i=5$, then [51, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-24}=c_{5}$, so we may assume that $i\leqslant 4$. First assume that $H$ is non-parabolic. If $|H|\leqslant q^{88}$, then the claim holds from the claim above as ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|\leqslant c_{i}$.
Now assume that $|H|>q^{88}$. The possibilities for $H$ are described in [25, 4.1] and one can check that ${\rm fpr}(x,G/H)\leqslant c_{i}$ by inspecting the proofs of [25, 4.3–4.5].
Finally assume that $H$ is parabolic. If $x$ is unipotent, then using [64], we can compute precise fixed point ratios in GAP for each conjugacy class and verify the claim. Now suppose that $x$ is semisimple. If $q\geqslant 3$, then using Magma, we compute the upper bound for $\chi(x)$ derived in Corollary 9.2, we verify that ${\rm fpr}(x,G/H)\leqslant c_{i}$ in all cases. If $q=2$, then the prime order semisimple classes of $G$ are listed in [6, Tables 3–15], and by calculating precise fixed point ratios for each class, we again deduce that ${\rm fpr}(x,G/H)\leqslant c_{i}$ in all cases.
Finally, if $x$ is a field automorphism, then Proposition 9.4 gives us
${\rm fpr}(x,G/H)\leqslant\frac{|E_{8}(q^{1/r})\,:\,\bar{H}(q^{1/r})|}{|E_{8}(q)\,:\,\bar{H}(q)|}<c_{i}.$
It now remains to show that the number of prime order elements $x\in G$ with $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ is bounded above by $d_{i}$ for all $i\in\{1,\ldots,5\}$. We treat each possibility of $i$ in turn.
Case 1: $i=1$.
In this case $d_{1}\leqslant|G|$, so the claim about $d_{1}$ is trivially true.
Case 2: $i=2$.
Note that if $182\leqslant{\rm dim}\,x^{\bar{G}}<202$, then $x$ is either semisimple or unipotent, or a field automorphism of order $5$. It follows from [65] that $G$ has fewer than $2q^{8}$ semisimple classes and we verify using [64] that there are exactly 24 unipotent classes with $182\leqslant{\rm dim}\,x^{\bar{G}}<202$. Finally, there are no more than 4 classes of field automorphisms of order $5$, and $|x^{G_{0}}|<q^{202}$, as shown above, so we get
$q^{202}(2q^{8}+28)<3q^{210}=d_{2}$
as required.
Case 3: $i=3$.
Here $154\leqslant{\rm dim}\,x^{\bar{G}}<182$, and so $x$ is either semisimple or unipotent, or a field automorphism of order $3$. By inspecting [65] we deduce that there are at most $2q^{2}+3q+6$ semisimple classes with $154\leqslant{\rm dim}\,x^{\bar{G}}<182$ and it follows from [64] that there are exactly 11 unipotent classes whose dimension is in this range. Finally there are no more than $2$ classes of involutory field automorphisms. It is also clear that $|x^{G}|<q^{182}$, and so we get
$q^{182}(2q^{2}+3q+19)<q^{188}=d_{3}$
as claimed.
Case 4: $i=4$.
By inspecting [64] and [65] one deduces that there are only $4$ unipotent classes with dimension in this range. Moreover, one can verify that $|x^{G}|<q^{154}$ for all such classes, and so we get $4q^{154}<q^{156}=d_{4}$, as claimed.
Case 5: $i=5$.
Inspecting [65] tells us that there are at most $q+2$ semisimple classes such that ${\rm dim}\,x^{\bar{G}}<129$ and it follows from [64] that there are $6$ such unipotent classes. Finally, there is a unique class of involutory field automorphisms in $G$, and $|x^{G}|<q^{129}$ if ${\rm dim}\,x^{\bar{G}}<129$, so we find
$q^{129}(q+9)<q^{133}=d_{5}$
which completes the proof.
[…]
Table 20. The values $a_{i},b_{i},c_{i},d_{i}$ as defined in Proposition 10.10.